### v1

# An alternative theory of magnetic flux tubes in strong fields via axion origin photons

Abstract
In our alternative theory, built around the coincidence of experimental and theoretical data, three ”free” parameters – the magnetic field in the tachocline of the order of $\sim 10^{7}~G$ (see Fig. (A.1) and Eq. (A17) in V. D. Rusov et al. (2021)), the axion mass $m_{a}\sim 3.2\cdot 10^{-2}~eV$ (see Eq. (11) in V. D. Rusov et al. (2021)), and the asymmetric dark matter (ADM) in the Universe with $m_{ADM}\sim 5~GeV$ ((see V. D. Rusov et al. (2021); A. C. Vincent et al. (2016)) – give a complete solution to the problem of the theory of magnetic flux tubes in strong fields with 11-year variations of axion-origin photons, which are caused by and anticorrelated to the 11-year variations in density of ADM, gravitationally captured on the Sun.
[…]
show]siiis@te.net.ua
1 Introduction
It is known that the unsolved problem of energy transport by magnetic flux tubes at the same time represents another unsolved problem related to the sunspot darkness (see 2.2 in M. Rempel & R. Schlichenmaier (2011)). Of all the known concepts playing a noticeable role in understanding the connection between the energy transport and sunspot darkness, let us consider the most significant theory, in our view. It is based on the Parker-Biermann cooling effect (E. N. Parker, 1955a; L. Biermann, 1941; E. N. Parker, 1979a) and originates from the early works of L. Biermann (1941) and H. Alfvén (1942).
As you know, the Parker-Biermann cooling effect (E. N. Parker, 1955a; L. Biermann, 1941; E. N. Parker, 1979a), which plays a role in our current understanding, originates from L. Biermann (1941) and H. Alfvén (1942): in a highly ionized plasma, the electrical conductivity can be so large that the magnetic fields are frozen into the plasma. Biermann realized that the magnetic field in the spots themselves can be the cause of their coolness – it is colder because the magnetic field suppresses the convective heat transport. Hence, the darkness of the spot is due to a decrease in surface brightness.
Parker (E. N. Parker, 1955a, 1974a, 1974b, 1974, 1979a) has pointed out that the magnetic field can be compressed to the enormous intensity only by reducing the gas pressure within the flux tube relative to the pressure outside, so that the external pressure compresses the field. The only known mechanism for reducing the internal pressure sufficiently is a reduction of the internal temperature over several scale heights so that the gravitational field of the Sun pulls the gas down out of the tube (as described by the known barometric law $dp/dz=-\rho g$). Hence it appears that the intense magnetic field of the sunspot is a direct consequence of the observed reduced temperature (E. N. Parker, 1955a).
On the other hand, Parker (E. N. Parker, 1974, 1977) has also pointed out that the magnetic inhibition of convective heat transport beneath the sunspot, with the associated heat accumulation below, raises the temperature in the lower part of the field. The barometric equilibrium leads to enhanced gas pressure upward along the magnetic field, causing the field to disperse rather than intensify. Consequently, E. N. Parker (1974) argued that the temperature of the gas must be influenced by something more than the inhibition of heat transport!
Our unique alternative idea is that the explanation of sunspots is based not only on the suppression of convective heat transfer by a strong magnetic field (of the order of $\sim 10^{7}~G$ (V. D. Rusov et al., 2021)) through the enhanced cooling of the Parker-Biermann effect (E. N. Parker, 1974a), but also on the appearance of the axions of photonic origin (Fig 3, Fig. B.1 in V. D. Rusov et al. (2021)) from the tachocline to the photosphere, which is confirmed by the “disappearance” of the heat, and consequently, the temperature in the lower part of the magnetic tube (E. N. Parker, 1974, 1977) due to the axions of photonic origin from the photon-axion oscillations in the O-loop near the tachocline (see Fig. 3).
This means that the appearance of axions of photonic origin, which remove the problem of the temperature rise in the lower part of the magnetic tube, and the photons of axionic origin, which have the free path (Rosseland length; see Fig. B.3 in (V. D. Rusov et al., 2021)) from the tachocline to the photosphere, are the explanation of sunspots based not only on the suppression of convective heat transport by a strong magnetic field (V. D. Rusov et al., 2021), but also on the indispensable existence of the Parker-Biermann cooling effect. At the same time, we clearly understand that stronger fields would seriously suppress the action of the dynamo (see (E. E. Deluca & P. A. Gilman, 1986)).
On the other hand, we understand that the existence of the Parker-Biermann cooling effect is associated with the so-called thermomagnetic Ettingshausen-Nernst effect (see Apendix A in V. D. Rusov et al. (2021)). Due to the large temperature gradient in the tachocline, the thermomagnetic EN effect (A. V. Ettingshausen & W. Nernst, 1886; E. H. Sondheimer, 1948; L. J. Spitzer, 1956; Y. B. Kim & M. J. Stephen, 1969) creates electric currents that are inversely proportional to the strong magnetic field of the tachocline. As we showed earlier (V. D. Rusov et al., 2021), the toroidal magnetic field of tachocline by means of the thermomagnetic Ettingshausen-Nernst effect (Appendix A in V. D. Rusov et al. (2021)) “neutralizes” the magnetic field of the solar core $\sim 5\cdot 10^{7}~G$ (W. A. Fowler et al., 1955; S. Couvidat et al., 2003). It means that, using the thermomagnetic EN effect, a simple estimate of the magnetic pressure of an ideal gas in the tachocline of e.g. the Sun,
$\frac{B_{tacho}^{2}}{8\pi}=p_{ext}\approx 6.5\cdot 10^{13}\frac{erg}{cm^{3}}~~at~~0.7R_{Sun},$ (1)
can indirectly prove that by using the holographic principle of quantum gravity (see Apendix C in V. D. Rusov et al. (2021)), the repelling toroidal magnetic field of the tachocline exactly “neutralizes” the magnetic field in the Sun’s core (see Fig. 2, and also Fig. A1 in V. D. Rusov et al. (2021))
$B_{tacho}^{Sun}=4.1\cdot 10^{7}~G=-B_{core}^{Sun},$ (2)
where the projections of the magnetic fields of the tachocline and the core have equal values but opposite directions.
Surprisingly, our experimental data provide observational evidence in support of a two-dimensional holographic explanation of the Sun and a powerful key to understanding the physics of quantum gravity in the Universe! This is explained once again by the fact that the experimental data, in which the toroidal magnetic field of the tachocline exactly ”neutralizes” the magnetic field in the solar core, should also theoretically describe both a two-dimensional holographic explanation of the solar tachocline, in which it is the magnetic field that must be solved, either in conformal field theory (see e.g. analogs in (R. Boyack et al., 2024; T. Kimura & S. Ozaki, 2019; A. B. Zamolodchikov, 1989; C. Callan et al., 1995; L. Hu et al., 2024; C. Cartwright et al., 2022; H. Nastase & K. Skenderis, 2020; V. A. Miransky & I. A. Shovkovy, 2015)) or at least in quantum mechanics (see e.g. analogs in (U. Gürsoy, 2021; S. Heefer, 2024)), and a three-dimensional magnetic field via gravity, which, in contrast to pseudo-Riemannian geometry, applies the extended (beyond Einstein geometry) Finsler geometry (see (U. Gürsoy, 2021; S. Heefer, 2024)), for example, inside the solar tachocline.
In other words, the “encoded” two-dimensional information about the magnetic field of the tachocline in conformal field theory corresponds to three-dimensional information about the magnetic field in Finsler gravity, which is reduced to the curvature of space-time in a three-dimensional volume with the radius of the tachocline of the Sun.
Such strong magnetic fields in a tachocline of e.g. the Earth, Sun, magnetic white dwarfs, accreting neutron stars and black holes can predict the exact “neutralization” of the magnetic field in the core of these stars and in a black hole (see Appendix C in V. D. Rusov et al. (2021)).
From here we understand that when space-time curvature appears, gravitational waves arise, which ultimately uncover, as not surprisingly, the more powerful Finslerian nature of quantum gravity. So Finslerian quantum gravity can be applied in our case, as we know, to holograms of the Universe, black holes, stars and planets, for example, the Sun.
Based on the “experimental example” of the existence and understanding of the physics of the holographic principle of quantum gravity, we apply the famous result of Stephen Hawking, who at a conference in Stockholm on August 28, 2015 announced that he had solved the information paradox. According to Hawking (S. W. Hawking, 2015a, b; S. W. Hawking et al., 2016), the AdS/CFT correspondence (see (J. Maldacena, 1998)) showed no loss of information, where all information about matter and energy in the three-dimensional volume of the black hole is encoded in the form of a hologram on its two-dimensional surface of the event horizon (see (G. ’t Hooft, 1993; L. Susskind, 1995; J. Maldacena, 2005; M. Hanada et al., 2014)). It resembles an analogue of the famous photohologram (see (D. Gabor, 1948)), which is a certain type of photograph that produces a three-dimensional image under good lighting; all the information describing a three-dimensional scene is encoded in a pattern of light and dark areas applied to a two-dimensional film. This means that the possible description and solution of the ”encoded” holographic two-dimensional ”film” is predetermined by the fact that the problem of the information paradox, known as the apparent contradiction between general relativity and quantum mechanics, completely disappears. This approach ultimately leads to a holographic principle of quantum gravity, which can be generalized to any physical system occupying space-time. For example, in the holograms of the Universe (N. Afshordi et al., 2017), black holes, white dwarfs, neutron stars or, surprisingly, in the hologram of the Sun’s tachocline!
Let us note one more important mechanism of quantum gravity (see Fig. 2b). The most surprising fact is that the antidynamo model of a magnetic flux tube at strong fields of the order of $10^{7}~G$ is a consequence of both the thermomagnetic EN effect and the so-called solar holographic antidynamo, in which the poloidal field arises directly from the toroidal field, but not vice versa (see Fig. 2b; and also Fig. C.1a in V. D. Rusov et al. (2021)).
Using the quantum gravity mechanism, and hence, the strong MFT field, we are interested in the existence of dark matter axions identical to solar axions, which strongly affects the the conversion of axion origin photons in magnetic O-loops near the tachocline, and are thus related to the so-called thermomagnetic Ettingshausen-Nernst effect through the Parker-Biermann cooling effect (see Apendix A in V. D. Rusov et al. (2021)).
Since the variations of axion origin photons are the result of the changes in Sun luminosity, then unlike a self-excited dynamo, an unexpected yet simple question arises: is there a dark matter chronometer hidden deep in the Sun? And consequently, the next question arises: who controls the Sun or who measures the precise “clock” inside the Sun’s core?
Figure 1: An illustration of the main possible processes of a magnetically active star of the Sun type. (a) $\alpha$-effect, $\Omega$-effect and BL mechanism as components of the solar dynamo model. The $\Omega$-effect (blue) depicts the transformation of the primary poloidal field into a toroidal field by differential rotation. Regeneration of the poloidal field is then performed either by the $\alpha$-effect (top) or by the BL mechanism (yellow in the middle). In case of $\alpha$-effect, the toroidal field at the base of the convection zone is subject to cyclonic turbulence. In the BL mechanism, the main process of regeneration of the poloidal field (based on the $\Omega$-effect (blue)) is the formation of sunspots on the surface of the Sun from the rise of buoyant toroidal flux tubes from the base of the convection zone. The magnetic fields of these sunspots closest to the equator in each hemisphere diffuse and join, and the field due to the spots closer to the poles has a polarity opposite to the current that initiates rotation of the polarity. The newly formed polar magnetic flux is transported by the meridional flow to deeper layers of the convection zone, thereby creating a new large-scale poloidal field. Derived from S. Sanchez et al. (2014). (b) 11-year variations of sunspot magnetic fields as a component of the solar anti-dynamo. In contrast to the component of the solar dynamo model (a) [^2] If the magnetic field at these levels is maintained by a dynamo, and the magnetic fields are diffuse, then the field strength is unlikely to be more than $\sim 10^{5}~G$, because larger fields would severely inhibit dynamo action (see E. E. Deluca & P. A. Gilman (1986))., the mechanism of the fundamental holographic principle of quantum gravity (see H. Nastase & K. Skenderis (2020); N. Afshordi et al. (2017)), in which the hologram of the Sun is encoded in the magnetic field on its two-dimensional surface tachocline and is equivalent to the magnetic field (from the core to the tachocline) in the internal three-dimensional space of the Sun’s gravity, and consequently, the formation of the thermomagnetic EN effect (see L. J. Spitzer (1962, 2006); V. Rusov et al. (2015)), emphasizes that this process is associated with the continuous transformation of toroidal field into poloidal one ($T\rightarrow P$ transformation), but not vice versa ($P\rightarrow T$). For this purpose, we note (seeV. D. Rusov et al. (2021)) that the modulations of the ADM halo in the Galactic Center (GC), which is an indicator of the periods of S-stars, e.g. S-102 with a period of about 11 years, lead (via vertical density waves from the black hole to the solar neighborhood) to modulations of the ADM density, gravitationally trapped in the interior of the Sun. This means that anticorrelation modulations of the ADM density control the solar luminosity or sunspot cycles, in which the sunspot variations are identical to the 11-year variations of the magnetic flux tubes (see Section 2): maximum (red lines) and minimum (green-red lines).
In order to answer this question, let us first consider all the unexpected and intriguing implications of the 11-year modulations of the ADM density in the solar interior and around the BH (see Sect. 3 in V. D. Rusov et al. (2021)).
A unique result of our model is the fact (see Sect. 3 in V. D. Rusov et al. (2021)), that the periods, velocities and modulations of the S-stars are the essential indicator of the modulation of the ADM halo density in the fundamental plane of the Galaxy center, which closely correlates with the density modulation of the baryon matter near the supermassive BH. If the modulations of the ADM halo at the GC lead to modulations of the ADM density on the surface of the Sun (through vertical density waves from the disk to the solar neighborhood), then there is an ”experimental” anticorrelation identity between such indicators as the ADM density modulation in the solar interior and the number of sunspots. Or equivalently, between the modulation of solar axions (or photons of axion origin) and the sunspot cycles!
A hypothetical pseudoscalar particle called axion is predicted by the theory related to solving the CP-invariance violation problem in QCD. The most important parameter determining the axion properties is the energy scale $f_{a}$ of the so-called U(1) Peccei-Quinn symmetry violation. It determines both the axion mass and the strength of its coupling to fermions and gauge bosons including photons. However, in spite of the numerous direct experiments, axions have not been discovered so far. Meanwhile, these experiments together with the astrophysical and cosmological limitations leave a rather narrow band for the permissible parameters of invisible axion (e.g. $10^{-6}eV\leqslant m_{a}\leqslant 10^{-2}eV$ (G. G. Raffelt, 2004, 2008a)). The PQ mechanism, solving the strong CP problem in a very elegant way (R. D. Peccei & H. R. Quinn, 1977a, b; F. Wilczek, 1978; S. Weinberg, 1978), is especially attractive here, since the axion is also a candidate for dark matter (J. Preskill et al., 1983; L. Abbott & P. Sikivie, 1983; M. Dine & W. Fischler, 1983; M. Kawasaki & K. Nakayama, 2013; D. J. Marsh, 2016; L. Di Luzio et al., 2020; P. Sikivie, 2021).
At present, we have shown that axions born in the core of the Sun, and photons of axion origin (in the solar corona with a temperature of one million degrees) yield a tight constraint on the axion-photon coupling constant $g_{a\gamma}\sim 4.4\cdot 10^{-11}~GeV^{-1}$ and mass $m_{a}\sim 3.2\cdot 10^{-2}~eV$ (see Fig. 2a; also Fig.1 in V. D. Rusov et al. (2021)). Most importantly, hadronic axions produced in the solar core are controlled by anticorrelated 11-year variations in the density of asymmetric dark matter (ADM) gravitationally trapped in the solar interior (via vertical density waves from the black hole disk to the solar neighborhood), with the hard part of the solar photon spectrum being of axion origin, which is defined as the product of the fraction of axions in the solar core and the fraction of the sunspot area. The latter is the result of solar luminosity variations (see Eqs. (27)-(28) in V. D. Rusov et al. (2021)). This means that solar luminosity variations and other solar cycles are the result of 11-year variations in ADM density, which are anticorrelated with the density of hadronic axions in the solar core.
Figure 2: (a) Summary of astrophysical, cosmological and laboratory constraints on axions. Comprehensive axion parameter space, highlighting two main front lines of direct detection experiments: helioscopes (CAST (CAST Collaboration, 2017)) and haloscopes (ADMX (S. J. Asztalos et al., 2010), and RBF (W. U. Wuensch et al., 1989) and microwave resonators (B. M. Brubaker et al., 2017; L. Zhong et al., 2018; T. Braine et al., 2020; S. Lee et al., 2020; K. M. Backes et al., 2021)). The astrophysical bounds from horizontal branch and massive stars are labeled “HB” (G. G. Raffelt, 2008b) and “Cepheids” (G. Carosi et al., 2013), respectively, and there are also astrophysical hints (WD cooling hints, and HB hint). The QCD motivated models (KSVZ (J. E. Kim, 1979; M. Shifman et al., 1980)) for axions lay in the yellow diagonal band. A plot of $g_{a\gamma}$ versus $m_{a}$ with the most stringent results (solid lines) and sensitivity perspectives (dashed lines) of observations and experiments directly comparable to the different phases of IAXO are shown, BabyIAXO, IAXO, and an upgraded version of IAXO, IAXO+. The yellow band denotes the region of the parameter space favoured by QCD axion models. The red star marks the values of the axion mass $m_{a}\sim 3.2\cdot 10^{-2}~eV$ and the axion-photon coupling constant $g_{a\gamma}\sim 4.4\cdot 10^{-11}~GeV^{-1}$, which were first obtained experimentally (see V. D. Rusov et al. (2021)).
(b) $R$ parameter constraints, which compares the numbers of stars in the the horizontal branch (HB) and in the upper portion of the red giant branch (RGB), to helium mass fraction $Y$ and axion coupling $g_{a\gamma}$ (adopted from (A. Ayala et al., 2014)). The resulting bound on the axion ($g_{10}=g_{a\gamma}/(10^{-10}~GeV^{-1})$ is somewhere between rather conservative $0.5<g_{10}<0.8$ and most aggressive $0.35<g_{10}<0.5$ (A. Friedland et al., 2013). The red line marks the value of the axion–photon coupling constant $g_{a\gamma}\sim 4.5\cdot 10^{-11}~GeV^{-1}$ adopted from Eq. (2) in (A. Ayala et al., 2014). The blue shaded area represents the bounds from Cepheids observation. The yellow star corresponds to $Y=0.254$ and the bounds from HB lifetime (yellow dashed line).
Interestingly, in contrast to our Fig. 2a, in Fig. 2b, we can see a well-known strong constraint on the axion-photon coupling $g_{a\gamma}\sim 4.5\cdot 10^{-11}~GeV^{-1}$, but in a wide mass range (see Fig. 2b and Eq. (2) in (A. Ayala et al., 2014)) from an analysis of a sample of 39 Galactic Globular Clusters. However, in our case (Fig. 2a) the wide range of masses in (A. Ayala et al., 2014) iso, ddly enough, replaced by a single mass $m_{a}\sim 3.2\cdot 10^{-2}~eV$, which fits not only the axions of galactic globular clusters (A. Ayala et al., 2014), but also all axion stars in our galaxy, including neutron stars (M. Buschmann et al., 2021), supernovae (P. Carenza et al., 2019, 2021), and the Sun (V. D. Rusov et al., 2021).
Taking into account the above, we structure our paper as follows. In Section 2, we discuss the physics of nearly empty flux tubes and the connection with dark matter axions out of the solar core. In Section 2.1, we present convective heating and magnetic buoyancy of flux tubes by means of axion origin photons. In Section 2.2, we discuss the physics of primary and secondary reconnection of flux tubes in the lower layers and the features of the Joy’s law tilt angle. In Section 2.2.2, we discuss the physics of the observed double maxima of sunspot cycles, where the first maximum does not depend on the tilt angle of the Joy’s law, while the second one definitely depends on the tilt angle of the Joy’s law. We also showed the relationship between the ADM density in the solar interior (or, for example, the number of anticorrelated sunspots) and the Gnevyshev gap. Finally, in Section 3, we provide a summary and outlook for this paper.
2 Physics of the theory of practically empty magnetic flux tubes at strong fields, $\sim 10^{7}~G$, at which there is a free path for photons of axion origin from the tachocline to the photosphere.
This raises the fundamental question of the self-exciting action of the dynamo, the remarkable physics of which was discovered some 70 years ago (see E. N. Parker (1955b, 1957)), beginning with the advent of Babcock’s solar magnetograph (H. W. Babcock & H. D. Babcock, 1955), and still today, but which, as no surprisingly, has not received a complete and final solution. Avoiding an alternative theory, most physicists believe that it is necessary to study more deeply the known physical problems of various aspects of the theory of solar dynamo (see B. B. Karak et al. (2014)) before passing a verdict on it. This continues to this day (see B. B. Karak (2024)).
Conversely, if the question arises whether there is a satisfactory alternative theory against the dynamo action, our short answer is ”yes”. We understand that the strong toroidal magnetic field in the tachocline is generated not by the dynamo action, but by the holographic mechanism of the tachocline (see Fig. 2b), which is a consequence of the holographic principle of quantum gravity in the Universe, in the black hole of our galaxy, and therefore on the Sun.
In order for our alternative theory (for example, magnetic flux tubes in strong fields using photons of axion origin) to solve all its problems (see Sections 2.1-2.2), it is necessary to solve the most important problem: In what way come sunspots, generated by a strong magnetic field in a flux tube from the base of the convection zone to the surface of the Sun, have 11-year variations without any dynamo?
The answer is very simple! Almost all physicists, and in particular dynamo physicists, know that in the Universe, and, of course, in our galaxy, there are dark matters, and the highest density of dark matter is located exactly at the center of the Galaxy, i.e. near the black hole. At the same time, we know that between the black hole and the disk there are also so-called amazing S-stars [^3] Let us remind from where the fairly young S-stars (Fig. 8a in (V. D. Rusov et al., 2021)) appear in our galaxy between the black hole and the disk. Since we know that in our galaxy, in addition to the supermassive black hole (SMBH), there is the so-called intermediate black hole (IMBH; see evidence (S. Takekawa et al., 2019)). If we take into account the evolution of binary SMBH-IMBH, then when one body (SMBH) exchanges partners of binary stars orbiting the IMBH (Fig. 8b in V. D. Rusov et al. (2021)), and through the extreme gravitational-tidal field, one star is captured (SMBH) and loses energy, while the other escapes, receives all this energy and, due to the hypervelocity of B-type stars – the IMBH sling (Fig. 8b in V. D. Rusov et al. (2021)), simply flies out of the galaxy. From here it becomes clear how in other galaxies a young S-star can end up between a black hole and a disk., which have a little more than 30 stars. The most interesting thing is that the closest S-star to the black hole is called S-102, which has an 11-yer period of around the black hole (see Fig. 6 and Fig. 8 in V. D. Rusov et al. (2021)). From here, we understand that when the period of S-102 is very close to the black hole, then the high speed of this star is such that the dark matter is practically not captured by this star. But when the elliptical star S-102 rotates farther from the black hole, then the star gravitationally captures these dark matters.
This means that the 11-year variations in the density of dark matter via vertical density waves from the black hole disk to the solar neighborhood V. D. Rusov et al. (2021) are also captured in the solar interior! Hence, the main result is that the 11-year variations in e.g. sunspots are never associated with any dynamo action.
Another intriguing question arises here. Unlike our galaxy, such S-type stars (between the black hole and the disk) may simply not exist in many galaxies, and consequently, there may not be any solar dynamo at all – for example, there is no Sun-like dynamo on the active star $\zeta$ Andromedae, which is part of the Zeta Andromedae star system in the constellation Andromeda (R. M. Roettenbacher et al., 2016). (It is located approximately 189 light years from Earth.) When various stars and clusters (e.g. Galactic Globular Clusters) have strong toroidal magnetic fields of $\geqslant 10^{7}~G$ in the corresponding tachoclines, then, unlike the self-excited dynamo, the manifestation of the tachocline itself in various stars and clusters is a result of the funda-mental holographic principle of quantum gravity (see J. M. Maldacena (1999); J. Maldacena (2005); S. W. Hawking (2015a, b); S. W. Hawking et al. (2016); and also N. Afshordi et al. (2017) and Fig.A.1 and Fig.C.1 in V. D. Rusov et al. (2021)). This means that the self-excited dynamo in the tachocline may not exist at all!
We could raise a few more questions, but it is enough for us that “…not enough for a scientific understanding of the solar dynamo. For there is an unfortunate revelation: with four free parameters you can provide a satisfactory fit to the horizon of New York or Beijing” – as was said by the great E. N. Parker (2009).
In this regard, let us return to the introduction of our section. Of all the known concepts that play a significant role in understanding the connection between the energy transfer of a magnetic flux tube and the darkness of sunspots, we will consider what we believe to be the most significant. It is based on the Parker-Biermann cooling effect (E. N. Parker, 1955a; L. Biermann, 1941; E. N. Parker, 1979a) in strong magnetic fields, which explains how the result of the suppression of Parker’s convective heat transfer manifests itself in the lower part of the magnetic flux tube (see Fig. 3).
In order to understand the physics of the Parker’s suppression of the convective heat transport in strong magnetic fields, we need to tun to the dark matter axions, born in the core of the Sun. With solar axions and the existence of a magnetic O-loop inside the MFT near the tachocline, the answer becomes very simple. When a magnetic O-loop is formed inside the MFT near the tachocline through the Kolmogorov turbulent cascade (see Fig. 4), the high-energy photons from the radiation zone experience the axion-photon oscillations in this O-loop, and the so-called axions of photonic origin appear under the sunspot. This means that the cooling effect of Parker-Biermann exists due to the disappearance of barometric equilibrium (E. N. Parker, 1974a) and, as a consequence, the manifestation of the photon free path (Rosseland length; see Fig. B.3 in V. D. Rusov et al. (2021)) from the tachocline to the photosphere, which is confirmed by axions of photonic origin after photon-axion oscillations in the O-loop (see Fig. 3).
Figure 3: Topological effects of magnetic reconnection inside the magnetic tubes with the “magnetic steps” (Fig. B.1 in V. D. Rusov et al. (2021)). The left panel shows the temperature and pressure change along the radius of the Sun from the tachocline to the photosphere (J. Bahcall & M. Pinsonneault, 1992), $L_{MS}$ is the height of the magnetic shear steps. At $R\sim 0.72~R_{Sun}$ the vertical magnetic field, which is developed from the horizontal part of the magnetic field (step) with the participation of the O-loop through the well-known Kolmogorov turbulent cascade (see Fig. 4), reaches $B_{z}\sim 3600$ T, and the magnetic pressure $p_{ext}=B^{2}/8\pi\simeq 5.21\cdot 10^{13}~erg/cm^{3}$ (J. Bahcall & M. Pinsonneault, 1992). The very cool regions along the entire convective zone caused by the Parker-Biermann cooling effect have the virtually zero internal gas pressure, i.e. the maximum magnetic pressure in the magnetic tubes. The narrow “purple” rings between the O-loop and the tube walls ($\rho_{ext}=\rho_{int}^{0}\gg\rho_{int}$) with the Parker-Bierman cooling effect inside, are a very important result of the existence of convective heating $(dQ/dt)_{2}$ in Section 2.1.
On the other hand, high-energy photon fluxes coming from the radiation zone through the thin “ring” part of the magnetic tube, i.e. between the tachocline and the “ring” (or more precisely between the magnetic wall and the O-loop of the tube; see Figs. 3 and 4), and which is applied without the Parker-Biermann cooling effect, allows to determine the convective heating $(dQ/dt)_{2}$ (see Sect. 2.1).
Figure 4: Kolmogorov turbulent cascade (A. Kolmogorov, 1941; A. N. Kolmogorov, 1968, 1991) and primary magnetic reconnection (which differs sharply from the secondary (see Fig. 5b) in the lower layers inside the unipolar magnetic tube (Fig. B.2 in V. D. Rusov et al. (2021)). Common to these various turbulent systems is the presence of the inertial region of Kolmogorov, through which the energy is cascaded from large to small scales. In this case, dissipative mechanisms (as a consequence of the “primary” magnetic reconnection) overcome the turbulent energy during plasma heating.
This solution clearly depends on the lifetime of the magnetic tubes rising from the tachocline to the surface of the Sun. That’s why due to the primary magnetic reconnection in the lower layers of the flux tubes (see Fig. 4d) this is not the final stage of the simulation. The essence of a practically empty tube in strong fields of $\sim 10^{7}~G$, which is born for the first time without a dynamo of any type (see Fig. 5a), is associated with the physics of turbulent reconnection of magnetic bipolar structures (see Fig. 5b,c). It is very important that this is due to the topological effects of secondary magnetic reconnection in the lower layers of the magnetic tube. From here, the bipolar part of the $\Omega$-loop rearranges itself at its base, compressing the $\Omega$-loop (blue lines; see Fig. 5b,c) and, as a consequence, organize a free O-loop (blue lines; Fig. 5d).
Figure 5: Sketch of primary or secondary magnetic reconnection near the tachocline. (a) The $\Omega$-loop forms a sunspot shadow (with photons of axion origin from O-loops associated with the primary reconnection) due to the indirect thermomagnetic EN effect in the tachocline, but without secondary reconnection; (b) The $\Omega$-loop with a spot in the presence of secondary reconnection, but without photons of axion origin (see b,c); the pink arrows show the upward convective flow between the “legs” of the $\Omega$-loop as it rises from the tachocline to the visible surface of the Sun; (c) The $\Omega$-loop with secondary reconnection and without a sunspot; (d) The O-loop without spots and reconnection. Passing stages (a), (b), (c) (from left to right), the convection around the ascending $\Omega$-loop “closes” it at the base (d) and, thus, a free O-loop is formed, and the initial configuration of the azimuthal field is restored near the tachocline. The blue arrows show the motion of matter leading to the connection of the “legs” of the loop and their “flying away” from the surface of the Sun.
As described by E. N. Parker (2009, 1994); H. C. Spruit et al. (1987); P. R. Wilson et al. (1990), the upward convection flow around the rising $\Omega$-loop brings its “legs” together in such a way that the magnetic field reconnection occurs across this loop. This cuts off the magnetic loop from the azimuthal magnetic field, turning it into an O-loop (see Fig. 4 in (E. N. Parker, 1994) and Fig. 3 in (H. C. Spruit et al., 1987)). After that the azimuthal magnetic field restores its initial configuration and becomes ready for another process with $\Omega$-loop.
Let us now make some important remarks on the turbulent reconnection, the $\Omega$-loop transformation into the O-loop by rapid “legs” closure, the restoration of the initial azimuthal field and the preconditions for another $\Omega$-loop formation in the same place. It is also necessary to explain the physical interpretation of the overshoot process near the tachocline and estimate the velocity $v_{rise}$ and time $\tau_{d}$ of the magnetic tube rise from the overshoot boundary layer – starting with the azimuthal magnetic flux strength of $B_{tacho}\sim 4\cdot 10^{7}~G$ (see Eq. A.1 in (V. D. Rusov et al., 2021)).
One of the final goals of this section is to determine the general regularities of the theory of magnetic flux tubes in strong fields of the order of $\sim 10^{7}~G$ using photons of axion origin – the so-called universal Ballegooijen-Fan-Fisher model (Sec. 2.1), which are generated by the magnetic buoyancy of practically empty tubes, and as a consequence, their rise from the tachocline to the solar surface (Figs. 3-5). Another goal is the essence of the physics of the primary and secondary reconnection of magnetic flux tubes in the lower layers of the tachocline (Sect. 2.2), which is directly related to the observed features of the cycle of both double sunspot maxima (see Sect. 2.2.2) and the tilt angle of Joy’s law (see Sect. 2.2.1). Moreover, both effects (see Sect. 2.1-2.2) are caused by the existence (of dark matter?) – solar axions generated in the core of the Sun.
2.1 Radiative $(dQ/dt)_{1}$ sec-reconnectionand convective $(dQ/dt)_{2}$ heatings and the buoyant rise of magnetic flux tubes and the phenomenon of solar axions
The first problem is devoted to the study of the effect of virtually empty magnetic tubes and the phenomenon of solar axions.
The assumption that the virtually empty magnetic tubes (Fig. 3) are neutrally buoyant ($\rho_{int}=\rho_{ext}$ (E. N. Parker, 1994)) implies that the temperature inside these tubes is lower than that of the ambient medium (Fig. 3 and Fig. 7a). This leads to the heat inflow, and consequently, the flux tube rises up (see E. N. Parker (1975) or Sect. 8.8 in (E. N. Parker, 1979b)). For a horizontal tube with a cross-section of radius $a$ the rise velocity follows from the Parker’s analysis (E. N. Parker (1975, 1979b), also Eq. (60) in (A. A. van Ballegooijen, 1982)):
$v_{rise}=2\frac{H_{p}}{\tau_{d}}\frac{B^{2}}{8\pi p_{ext}}\left(-\delta+0.12\frac{B^{2}}{8\pi p_{ext}}\right)^{-1},$ (3)
where $H_{p}=\Re T_{ext}/g=p_{ext}/g\rho_{ext}=0.08R_{Sun}$ (E. Böhm-Vitense, 1958; H. C. Spruit, 1977; A. S. Brun et al., 2011) is the pressure scale height at the tachocline, $T_{ext}$ and $p_{ext}$ are the external gas temperature and pressure, $\delta\equiv Y=\nabla_{e}-\nabla_{ad}=-c_{p}^{-1}dS/d\xi=-c_{p}^{-1}H_{p}dS/dz$ is the dimensionless entropy gradient (see A. A. van Ballegooijen (1982); R. Smolec & P. Moskalik (2008, 2010)), $\nabla_{e}\equiv d\ln T_{e}/d\ln p_{e}$ and $\nabla_{ad}\equiv(\partial\ln T_{e}/\partial\ln p_{e})_{s}$ are the local and adiabatic temperature gradients in external and internal plasma (H. Spruit, 1974; A. A. van Ballegooijen, 1982; J. Christensen-Dalsgaard et al., 1995), $s$ is the specific entropy, $c_{p}$ is the heat capacity at constant pressure, and $\tau_{d}$ is the rise time of the radiative heating of the magnetic flux tube:
$\tau_{d}=\frac{c_{p}\rho a^{2}}{k_{e}}\simeq c_{p}\rho a^{2}\left[\frac{c_{p}F_{tot}}{g}\left(1+\frac{2\ell_{ov}}{5H_{p}}\right)^{\nu}\right]^{-1}.$ (4)
where for the fully ionized gas $c_{p}=2.5\Re=2.5p_{ext}/\rho_{ext}T_{ext}$ ($\Re$ is the the equation of state $p_{ext}=\rho_{ext}\Re T_{ext}=nk_{B}T_{ext}$ (E. Böhm-Vitense, 1958)), $T(z)$ and $\rho(z)$ are the mean temperature and density; $k_{e}$ is the radiative heat conductivity (see Eq. (35) in A. A. van Ballegooijen (1982)); $\ell_{ov}\approx 0.37H_{p}$ (A. A. van Ballegooijen, 1982; J. Christensen-Dalsgaard et al., 2011) is the thickness of the overshoot layer; the total radiative energy flux $F_{tot}=L/(4\pi r^{2})$ depends on the Sun luminosity $L$; $g$ is the gravitational acceleration.
Next we apply the condition of hydrostatic equilibrium, $dp/dz=\rho g$, when the adiabatic temperature gradient $(dT/dz)_{ad}=g/c_{p}$ may be used, and the neutral buoyancy of the flux tube in the overshoot zone $(|\delta T|/T_{ext})^{-1}\sim\beta\equiv 8\pi p_{ext}/B^{2}$. This way we are able to estimate the time of the radiative and/or convective diffusion $\tau_{d}$ (see Eq. (4)) of the flux tube:
$\tau_{d}=\frac{c_{p}\rho a^{2}}{k_{e}}\approx|\delta T|c_{p}\rho\frac{a^{2}}{(1.148)^{\nu}\delta z|F_{tot}|},$ (5)
where
$\delta z\sim(1.148)^{-\nu}\frac{1}{2}\left(\frac{a}{H_{p}}\right)^{2}H_{p}\frac{\nabla_{e}}{\nabla_{rad}},~~where~~\nu\geqslant 3.5,$ (6)
$F_{tot}=\frac{L}{4\pi R_{tacho}^{2}}=H_{p}\frac{\nabla_{rad}}{\nabla_{e}}\left(\frac{dQ}{dt}\right)_{1}.$ (7)
Here $\nabla_{rad}=(\partial\ln T_{ext}/\partial\ln p_{ext})_{rad}$ is the radiative equilibrium temperature gradient; $(dQ/dt)_{1}$ is the rate of radiative heating, which only depends on the thermodynamic parameters $k_{e}$ and $T_{ext}$ of the ambient plasma, depending only on the radial distance from the Sun center (A. A. van Ballegooijen, 1982; J. Christensen-Dalsgaard et al., 1995).
As a result, it is not difficult to show that the van Ballegooijen model combining equations (3)-(7) gives the final expression for the rise time by radiation heating from the boundary layer of the overshoot to the solar surface,
$\tau_{d}\approx\frac{2}{\beta}T_{ext}\left[\frac{1}{c_{p}\rho_{e}}\left(\frac{dQ}{dt}\right)_{1}\right]^{-1},$ (8)
at which we have a very important (see section 2.2.1) value of
$\tau_{d}\approx\left.5\frac{1}{\beta}p_{ext}\middle/\left(\frac{dQ}{dt}\right)_{1}\right.\approx 0.84\cdot 10^{8}~sec\approx 2.5~years,$ (9)
by means of $1/\beta=B^{2}/8\pi\rho_{ext}\approx 2.9\cdot 10^{-5}$, which is identical to the magnetic field $B\approx B_{eq}\sim 2\cdot 10^{5}~G$ (see $B^{2}/8\pi=p_{ext}=nk_{B}T_{ext}\approx 1.665\cdot 10^{13}~erg\cdot cm^{-3}$ (at $0.8R_{Sun}$ (see e.g. Fig. 3)) and $(dQ/dt)_{1}\approx 29.7~erg\cdot cm^{-3}\cdot s^{-1}$ (see Eq. (18) in Y. Fan & G. Fisher (1996)).
At the same time, the rate of the MFT rise from the tachocline to the solar surface
$v_{rise}=H_{p}\nabla_{ad}\frac{1}{p_{ext}}\left(\frac{dQ}{dt}\right)_{1}\left(|\delta|+0.12\frac{B^{2}}{8\pi p_{ext}}\right)^{-1},~~~\nabla_{ad}=\nabla_{e}\simeq 0.4,$ (10)
at which we have a very important (see Sections 2.1-2.2) value of
$v_{rise}\approx 1.7\cdot 10^{2}~cm/s,$ (11)
by means of $H_{p}\approx 5.6\cdot 10^{4}~km$ and $(|\delta|+0.12B^{2}/8\pi p_{ext})=(3/5)\times B^{2}/8\pi p_{ext}+0.12\times B^{2}/8\pi p_{ext}\approx 0.72\times 2.9\cdot 10^{-5}$ (see Eqs. (58-60) in (A. A. van Ballegooijen, 1982)), are almost identical to the equations (29)-(30) of Y. Fan & G. Fisher (1996).
At the same time, it is easy to show that these values
$v_{rise}=\frac{2.77H_{p}}{\tau_{d}},$ (12)
provide a simple relationship between the rise time and the rate of ascent of the magnetic flux tube from the overshoot boundary layer to the surface of the Sun.
Surprisingly, these equations (Eqs. (8) and (10)) are almost identical to equations (29) and (30) from Y. Fan & G. Fisher (1996), which arise from completely different equations (60)-(61) of J. Bahcall & M. Pinsonneault (1992).
Unlike the special van Ballegooijen model, we adopt the universal model of MFTs with
$v_{rise}=2\frac{H_{p}}{\tau_{d}}\frac{B^{2}}{8\pi p_{ext}}\left(-\delta+0.12\frac{B^{2}}{8\pi p_{ext}}\right)^{-1},$
$where$ (13)
$\tau_{d}=\frac{c_{p}\rho a^{2}}{k_{e}}\left\{\left(\frac{dQ}{dt}\right)_{1}\left[1+\left(\frac{dQ}{dt}\right)_{2}\middle/\left(\frac{dQ}{dt}\right)_{1}\right]\right\}^{-1},$
Here the second term $(dQ/dt)_{2}$ represents a convective diffusion across the flux tube that is due to the temperature difference ($\delta T\equiv T-T_{e}$) between the flux tube and the external plasma (see (Y. Fan & G. Fisher, 1996)).
Using simple calculations of equations (3) and (2.1) for MFTs, it is easy to show that with the help of the
$\frac{dQ}{dt}=\left(\frac{dQ}{dt}\right)_{1}+\left(\frac{dQ}{dt}\right)_{2}$ (14)
of the universal model
$\tau_{d}\approx\frac{2}{\beta}T_{ext}\left[\frac{1}{c_{p}\rho_{e}}\frac{dQ}{dt}\right]^{-1},$ (15)
and
$v_{rise}=H_{p}\nabla_{ad}\frac{1}{p_{ext}}\frac{dQ}{dt}\left(-\delta+0.12\frac{B^{2}}{8\pi p_{ext}}\right)^{-1},~~~\nabla_{ad}=\nabla_{e}\simeq 0.4,$ (16)
is the general case of the so-called universal model of the van Ballegooijen-Fan-Fisher, which is completely identical to equation (27) of the model by Y. Fan & G. Fisher (1996):
$\frac{d\Delta\rho}{dt}=\rho_{e}\frac{v_{rise}}{H_{p}}\left[\delta+\left(\frac{1}{\gamma}-\frac{2}{\gamma^{2}}\right)\frac{1}{\beta}-\frac{1}{\gamma}\frac{\Delta\rho}{\rho_{e}}\right]+\frac{2}{\gamma\beta}\rho_{e}\left[\frac{\partial(v\cdot I)}{\partial s}-\vec{v}\cdot\vec{k}\right]+\frac{\rho_{e}}{p_{e}}\nabla_{ad}\left[\left(\frac{dQ}{dt}\right)_{1}+\left(\frac{dQ}{dt}\right)_{2}\right],$ (17)
where in case of quasi-static rising, it is necessary apply ($d\Delta\rho/dt$) and $\Delta\rho$ to zero (see equation (27) of Y. Fan & G. Fisher (1996)). Furthermore, according to Y. Fan & G. Fisher (1996), for simplicity, the lifting of a uniform horizontal magnetic flux tube in the region of a plane-parallel overshoot region is considered, under which the conditions $\partial(v\cdot I)/\partial s=0$ and $\vec{v}\cdot\vec{k}=0$ can be used in equation (27) of Y. Fan & G. Fisher (1996): $I\equiv\partial r/\partial s$ is the unit vector tangential to the flux tube, and $k\equiv\partial^{2}r/\partial s^{2}$ is the tube’s curvature vector.
It should also be recalled that, taking into account the presence of subadiabaticity $\delta<0$ (see Eq. (58) in (A. A. van Ballegooijen, 1982), Eq. (27) in (Y. Fan & G. Fisher, 1996), and also Fig. 6) of the overshoot tachocline and using, according to A. A. van Ballegooijen (1982), non-local mixing length theory (H. C. Spruit & A. A. van Ballegooijen, 1982; E. Böhm-Vitense, 1958; S. Z. Ali & S. Dey, 2020; M. A. Weber & Y. Fan, 2015), it can be shown that thin, neutrally buoyant flow tubes, stable in a stratified medium, provided that its field strength $B$ is smaller than a critical value $B_{c}$ (A. A. van Ballegooijen, 1982), which is approximately given by
$\frac{B_{c}^{2}}{8\pi p_{ext}}=-\gamma\delta=-\frac{5}{3}\delta.$ (18)
Ultimately, we call these Eqs. (2.1)-(18), in honor of these remarkable “solar” physicists, the universal van Ballegooijen-Fan-Fisher model (vanBFF model).
On the other hand, let us remind that on the basis of the holographic mechanism, generating the toroidal magnetic field in the tachocline, the universal model of flux tubes is predetermined by the existence of strong magnetic fields of the order of $B_{tacho}\sim 10^{7}~G$. Since the physics of the holographic mechanism does not involve a magnetic dynamo, we often refer to it as the universal antidynamo vanBFF model. It is determined by the following total energy rate per unit volume:
$\frac{dQ}{dt}=\left(\frac{dQ}{dt}\right)_{1}+\left(\frac{dQ}{dt}\right)_{2}=\left(\frac{dQ}{dt}\right)_{1}\left[1+\frac{\alpha_{1}^{2}}{\nabla_{e}}\left(\frac{H_{p}}{a}\right)^{2}\frac{1}{\beta}\right],$ (19)
where
$\left(\frac{dQ}{dt}\right)_{1}=-\nabla\vec{F}_{rad}=F_{tot}\frac{\nabla_{e}}{\nabla_{rad}}\frac{1}{H_{p}}=k_{e}\nabla_{e}\frac{T_{e}}{H_{p}^{2}},$ (20)
$\left(\frac{dQ}{dt}\right)_{2}=-k_{e}\frac{\alpha_{1}^{2}}{a^{2}}(T-T_{e}),$ (21)
where we used the MFT heating rate which consists of the radiative heating rate $(dQ/dt)_{1}$ and convective heating rate $(dQ/dt)_{2}$ (see Eqs. (2.1)–(19), Eqs. (13) and (19) in (Y. Fan & G. Fisher, 1996) and Eqs. (8)–(9) in (M. A. Weber & Y. Fan, 2015)); the approximate ratio $|\delta T|/T_{ext}\sim 1/\beta$; the parameter $\alpha_{1}^{2}\approx 5.76$ (Y. Fan & G. Fisher, 1996; M. A. Weber & Y. Fan, 2015); $F_{rad}$ is the radiative energy flux (see Eq. (15) in (Y. Fan & G. Fisher, 1996), and also Eq. (6) and Fig. 1 in (M. A. Weber & M. K. Browning, 2016)); $\nabla_{e}\sim 1.287\nabla_{rad}$ (see Table 2 in (H. C. Spruit, 1977)); $H_{p}/a$ is the factor for the lower convection zone (see (R. M. Roettenbacher et al., 2016)), where it was previously believed that $a=(\Phi/\pi B_{tacho})^{1/2}\leqslant 0.1H_{p}$ (see 3.2.1 in (R. M. Roettenbacher et al., 2016)) with an average value of typical magnetic flux of $\Phi\sim 10^{20}-10^{22}~Mx$ (Y. Fan et al., 1993).
From this, we understand that it is with the strong fields that a sharp increase in convective heating $(dQ/dt)_{2}$ is necessary in the tube, which simultaneously leads to a sharp decrease in the outer width of the ring $a\equiv a_{conv}$ (between the O-loop and the tube walls in Fig. 6) in a practically empty magnetic flux tube (see Fig. 3 and Fig. 6). In this case, the area of the “ring” of the magnetic tube is approximately equal to $2\pi r\times a_{conv}$, where $r$ is the radius of the magnetic tube, and $a_{conv}$ is the width of the “ring”.
What is the physics behind the appearance of the “ring” between the O-loop and the walls of the magnetic tube (see Fig. 6; also Fig. 3)? In simple words, one can say the following. The appearance of the “ring” cross-section is, as a consequence, the result of the production of both axion-origin photons (by converting solar axions into photons in the tachocline) and photon-origin axions (through the conversion of high-energy photons from the radiation zone to axions in the tachocline). From here, on the one hand, the axions of photonic origin are the part of the manifestation of the mean free path of axion origin photons from the tachocline to the photosphere. On the other hand, they are the part of the manifestation of magnetic tube “ring”, where the convective heating $(dQ/dt)_{2}$ dominates over the radiation heating $(dQ/dt)_{1}$.
Figure 6: A sketch of the magnetic tube born anchored to the tachocline and risen to the solar surface by the neutral buoyancy ($\rho_{ext}=\rho_{int}^{0}$). The strong convection suppression inside the tube leads to the abrupt decrease of temperature and density ($\rho_{int}^{0}\gg\rho_{int}$), which in its turn leads to the significant decrease in gas pressure above the umbra (sunspot). At the top of the overshoot tachocline the first term $(dQ/dt)_{1}$ characterizes the radiative heating which depends on the thermodynamic quantities $k_{e}xt$ and $T_{ext}$ of the external plasma, changing with the distance from the center of the Sun only (see Eq. (20)). The second term $(dQ/dt)_{2}$ represents the diffuse radiation through the flux tube because of the temperature difference between the tube ring (see $a\equiv a_{conv}$ between the O-loop and the tube walls) and the surrounding plasma (see Eq. (21)). The keV photons (see Fig. 2 in J. E. Bailey et al. (2009)) coming from the radiation zone are turned into axions in the horizontal magnetic field of the O-loop (see Fig. 3). Therefore, the radiative heating almost vanishes in the virtually empty magnetic tube. The base of the convection zone is defined as a radius at which the stratification switches from almost adiabatic ($\delta=\nabla_{e}-\nabla_{ad}=0$) to sub-adiabatic ($\delta=\nabla_{e}-\nabla_{ad}<0$). Meanwhile, the external plasma turns from sub-adiabatic to super-adiabatic ($\delta=\nabla_{e}-\nabla_{ad}>0$).
At the same time, we remember that anticorrelated 11-year ADM density modulations in the Sun interior (see Fig. 2b) and the ADM around BH (see Sect. 3 in (V. D. Rusov et al., 2021)) are not only interconnected, but simultaneously, as consequence, are driving the 11-year modulations of the solar axion density (see Eqs. (27)-(28) in (V. D. Rusov et al., 2021)).
As a result, assuming 11-year variations in convective heating $(dQ/dt)_{2}$ in the form of a maximum and minimum of the “ring” width of the $a\equiv a_{conv}$ (between the O-loop and the tube walls in Fig. 6)
$(a_{conv})_{max}\sim 3.7\cdot 10^{-4}H_{p}\approx 22~km,$ (22)
$(a_{conv})_{min}\sim 3.7\cdot 10^{-3}H_{p}\approx 220~km,$ (23)
we apply a new analysis of the universal vanBFF model (see Eqs. (32.1)-(18) analogous to (A. A. van Ballegooijen, 1982) or (19)-(21) (analogous to the last part of Eq. (13) and (19) in (Y. Fan & G. Fisher, 1996))), where the calculated values such as the magnetic flux $\Phi$ and the rise speed $(v_{rise})_{conv}$ of the MFT to the surface of the Sun, do not contradict the known observational data:
$\bullet$ values of the magnetic tube (see Fig. 5a), based on the physics of convective heating $(dQ/dt)_{2}$, are related to the radius of the “ring” between the O-shaped contour and the walls of the magnetic tube (Fig. 6), and which are also the result of maximum magnetic flux
$\Phi_{max}=2\pi r(a_{conv})_{min}(B_{tacho})_{max}\approx 10^{22}~Mx,$ (24)
where $r\sim 220~km$, $(a_{conv})_{min}\sim 22~km$ and $(B_{tacho})_{max}\sim 3.6\cdot 10^{7}~G$, and minimum magnetic flux
$\Phi_{min}=\pi(a_{conv})_{max}^{2}(B_{tacho})_{min}\approx 5\cdot 10^{21}~Mx,$ (25)
where $\pi(a_{conv})_{max}^{2}\approx\pi(220^{2}-20^{2})~km^{2}=1.5\cdot 10^{15}~cm^{2}$ and $(B_{tacho})_{min}\sim 3.6\cdot 10^{6}~G$, and which is in good agreement with the observational data of C. Zwaan (1987) (see also Section 2.1 in (Y. Fan, 2021)) since the total magnetic flux ranges are $1-5\cdot 10^{22}~Mx$ (near the maximum) and $2-5\cdot 10^{21}~Mx$ (near the minimum) (see (K. L. Harvey et al., 1982)).
It is very important here that, on the one hand, the anchored magnetic field in the tachocline rises from the bottom of the convective zone to the solar photosphere in the form of thin isolated filaments known as flux tubes (see e.g. (E. N. Parker, 1979a)), and on the other hand, the well-known nonlinear dynamics of a thin magnetic tube was modeled in the form that has been used by H. C. Spruit (1981). In this case, the magnetic tube is assumed to be thin in the sense that its cross-sectional radius $a_{conv}$ is negligibly small with respect to both the atmospheric scale height (i.e. $a_{conv}\ll H_{p}$ (see Fig. 6)) and any variation scales along the tube. This means that each magnetic field line is roughly parallel to the axis of the tube, and the Spruit’s model is applicable only to untwisted magnetic tubes.
This raises a rather complicated but simple question: In what way, without twisting, can a real magnetic tube, which is not integral and will not behave as a single object for a very long time (E. N. Parker, 1979a; H. C. Spruit, 1981; D. W. Longcope & I. Klapper, 1997), be obtained without destruction by hydrodynamic forces?
The answer, oddly enough, is very simple! For an untwisted thin magnetic tube, the pressure balance is always maintained not only along the entire radius of the “ring” between the O-shaped contour and the walls of the magnetic tube (Fig. 6), which is solved without destruction by hydrodynamic forces, but at the same time leads to the suppression of convection along the inner radius of the O-shaped contour due to a sharp decrease in temperature and density (see $\rho_{int}^{0}\gg\rho_{int}$ in Fig. 6). The latter means that in very strong magnetic fields (e.g. $5\cdot 10^{5}~G<B<10^{8}~G$) an untwisted thin magnetic tube will, surprisingly, always be perpendicular to the azimuthal direction from the tachocline to the surface of the photosphere!
The physics of this process is briefly as follows. The high-energy photons from the radiation zone through axion-photon oscillations in the O-loop inside the magnetic tube near the tachocline create the so-called axions of photon origin under the sunspot. This means that at such strong magnetic fields, the Parker-Biermann effect of cooling inside the magnetic tube, where the Sun’s gravitational field pulls gas out of the inner tube (as described by hydrostatic pressure, known as the barometric law $dp/dz=-\rho g$), develops due to the “disappearance” of the convective heat transfer $(dQ/dt)_{2}$ and, as a consequence, the temperature in the lower part of the magnetic tube with the help of axions of photon origin from photon-axion oscillations in the O-loop near the tachocline. As a result, it opens a free path (in the radial direction!) for photons of axion origin (Rosseland length; see Fig. B.3 in (V. D. Rusov et al., 2021)) from the tachocline to the photosphere (see Fig. 5a)!
$\bullet\bullet$ the time of rise of convective heating of the flux tube from the tachocline to the photosphere (see Eq. (15)), which corresponds to the equations
$(\tau_{d})_{conv}$ $=\frac{2}{\beta}T_{ext}\left\{\frac{1}{c_{p}\rho_{ext}}\left(\frac{dQ}{dt}\right)_{1}\left[1+\frac{\alpha_{1}^{2}}{\nabla_{e}}\frac{1}{\beta}\left(\frac{H_{p}}{a_{conv}}\right)^{2}\right]\right\}^{-1}$ (26)
$=\frac{2}{\beta}T_{ext}2.5\Re\rho_{ext}\left\{\left(\frac{dQ}{dt}\right)_{1}\left[1+\frac{\alpha_{1}^{2}}{\nabla_{e}}\frac{1}{\beta}\left(\frac{H_{p}}{a_{conv}}\right)^{2}\right]\right\}^{-1}$
$=\frac{5}{\beta}T_{ext}(p_{ext}/\rho_{ext}T_{ext})\rho_{ext}\left\{\left(\frac{dQ}{dt}\right)_{1}\left[1+\frac{\alpha_{1}^{2}}{\nabla_{e}}\frac{1}{\beta}\left(\frac{H_{p}}{a_{conv}}\right)^{2}\right]\right\}^{-1}$
$=\frac{5}{\beta}p_{ext}\left\{\left(\frac{dQ}{dt}\right)_{1}\left[1+\frac{\alpha_{1}^{2}}{\nabla_{e}}\frac{1}{\beta}\left(\frac{H_{p}}{a_{conv}}\right)^{2}\right]\right\}^{-1}$ (27)
$=\frac{5}{\beta}nk_{B}T_{ext}\left\{\left(\frac{dQ}{dt}\right)_{1}\left[1+\frac{\alpha_{1}^{2}}{\nabla_{e}}\frac{1}{\beta}\left(\frac{H_{p}}{a_{conv}}\right)^{2}\right]\right\}^{-1}$ (28)
$\cong\frac{5}{\beta}\rho_{ext}\mu^{-1}m_{p}^{-1}k_{B}T_{ext}\left\{\left(\frac{dQ}{dt}\right)_{1}\left[1+\frac{\alpha_{1}^{2}}{\nabla_{e}}\frac{1}{\beta}\left(\frac{H_{p}}{a_{conv}}\right)^{2}\right]\right\}^{-1}$ (29)
at the maximum value of the 11-year variations of the magnetic flux tube in the photosphere for sunspots (see (27)), in the form
$(\tau_{d})_{conv}^{max}\approx 5\times(6.5\cdot 10^{13}~erg\cdot cm^{-3})/3\cdot 10^{9}~erg\cdot cm^{-3}\cdot s^{-1}$
$\approx 10^{15}~erg\cdot cm^{-3}/3\cdot 10^{9}~erg\cdot cm^{-3}\cdot s^{-1}\approx 1.1\cdot 10^{5}~s,$ (30)
or
$(\tau_{d})_{conv}^{max}\approx 1.1\cdot 10^{5}~s\sim 1.3~day,$ (31)
where $p_{ext}=6.53\cdot 10^{13}~erg/cm^{3}$ (see A.16 in (V. D. Rusov et al., 2021), and also (J. Bahcall & M. Pinsonneault, 1992)), $T_{ext}=2.3\cdot 10^{6}~K$ and $\rho_{ext}=0.2~g/cm^{3}$ (J. Bahcall & M. Pinsonneault, 1992); $c_{p}=2.5\Re=2.5p_{ext}/\rho_{ext}T_{ext}$ ($\Re$ is the gas constant in the equation of state $p_{ext}=\rho_{ext}\Re T_{ext}=nk_{B}T_{ext}$ (E. Böhm-Vitense, 1958)); $\mu=0.58$ is mean molecular weight of hydrogen in overshoot tachocline, $m_{p}\cong 1.67\cdot 10^{-24}~g$ is the proton mass: ”2018 CODATA Value: proton mass”, The NIST Reference on Constants, Units, and Uncertainty, 20 May 2019; $k_{B}=1.38\cdot 10^{-23}~J/K$; $(dQ/dt)_{1}\approx 29.7~erg\cdot cm^{-3}\cdot s^{-1}$ (see Eq. (18) in (Y. Fan & G. Fisher, 1996)) is the rate of radiative heating; $\alpha_{1}^{2}\approx 5.76$ (Y. Fan & G. Fisher, 1996; M. A. Weber & Y. Fan, 2015); $\nabla_{e}\approx 0.4$ (see Table 2 in (H. Spruit, 1974)); $H_{p}=\Re T_{ext}/g=p_{ext}/g\rho_{ext}=0.08R_{Sun}$ (E. Böhm-Vitense, 1958; H. C. Spruit, 1977; A. S. Brun et al., 2011); $a_{conv}^{max}\sim 3.7\cdot 10^{-4}H_{p}$ (see Eq. (31)).
In order to obtain the minimum values of the 11-year variations of the magnetic flux tube in the photosphere for sunspots, we have to calculate a new value for $p_{ext}=(nk_{B}T)_{min}$, which together with $a_{conv}\sim 3.7\cdot 10^{-3}~H_{p}$ (Eq. (23)) is related to Eqs. (32)-(33). Using the value of fully ionized hydrogen plasma with $Z=1$ in the tachocline, which is related to the equation $T^{1/4}n=const$ (see Eq. (A.9) from Appendix A in (V. D. Rusov et al., 2021)), we obtain the relation
$(T^{1/4}\rho_{ext})_{min}=(T^{1/4}\rho_{ext})_{max}=((2.3\cdot 10^{6}~K)^{1/4}\times 0.2~g/cm^{3})_{max}.$ (32)
When we apply the temperature “minimum” value $(T\sim 10^{6}~K)_{min}$ (see Eq. (28) from (V. D. Rusov et al., 2021)), then we get the following relation
$\mu\times m_{p}\times(T^{1/4}n)_{min}=((10^{6}~K)^{1/4}\rho_{ext})_{min}=\mu\times m_{p}\times(T^{1/4}n)=((2.3\cdot 10^{6}~K)^{1/4}\times 0.2~g/cm^{3})_{max},$ (33)
from which we obtain the complete solution of the “minimum” rise time of the convective heating of the flux tube from the tachocline to the photosphere (see Eq. (15), which corresponds to the equation
$(\tau_{d})_{conv}^{min}\approx 5.7\cdot 10^{5}~s\sim 7~days,$ (34)
where $p_{ext}=(nk_{B}T)_{min}=3.56\cdot 10^{13}~erg/cm^{3}$, $T_{ext}=10^{6}~K$, $\rho_{ext}\approx 0.25~g/cm^{3}$ (see Eq. 33).
And, as a consequence, the magnitude of the MFT rise speed (see Eq. (16)) will be equal to
$(v_{rise})_{conv}=\frac{2.77H_{p}}{(\tau_{d})_{conv}},$ (35)
both at the maximum of the solar cycle
$(v_{rise})_{conv}^{max}\sim 1.40~km/s,$ (36)
which are almost identical to the observational data of the known works (see about 1.4 m/s at 62-75 and 42-55 Mm for AR11726, April 2013 in A. G. Kosovichev et al. (2016, 2018); see about 1.2 km/s at 16-32 Mm for AR NOAA 6432, South Pole in 1991 January 4-5, in A. Kosovichev (1996)), and at the minimum of the solar cycle
$(v_{rise})_{conv}^{min}\sim 0.25~km/s,$ (37)
which are also almost identical to the observational data, e.g. of the known works by S. Ilonidis et al. (2011, 2012, 2013). This is due to the fact that these works, on the one hand, showed the measured velocities of the magnetic flux escape from the deep interior of the Sun into the atmosphere in the form of sunspots (see $\sim 1~km/s$ for 2003 October 25 (Fig. 5.14 in S. Ilonidis et al. (2012); Fig. 8 in S. Ilonidis et al. (2013))), and on the other hand, using deep-focus time-distance helioseismology T. L. Duvall et al. (1993, 1996); A. G. Kosovichev et al. (2016) found several strong emerging flux events (for AR 10488, October 26, 2003, + ARs 8164 and 8171, 23-27 February 1998 + AR 7978, 06 July 1996) at depths of 42-75 Mm up to 6 h (using only previous data) before entering the photosphere and 1-2 days before until the appearance of the detected magnetic structures on the surface with an average speed of 0.3 to 0.6 km/s (S. Ilonidis et al., 2011). This means that, unlike the works of S. Ilonidis et al. (2011, 2012, 2013), we use the velocities of emerging flows on the surface of sunspots between the maximum ((31) and (36)) and minimum ((34) and (37)) of the solar cycle (for AR 10488, October 26, 2003, + ARs 8164 and 8171, 23-27 February 1998 + AR 7978, 06 July 1996), which correspond close to the maximum $\sim 1~km/s$ (for 2003 October), lower (by assumption!) than middle of the maximum and minimum 0.3 km/s +0.3 km/s (for ARs 8164 and 8171, 23-27 February 1998) and low $\sim 0.25~km/s$ (for AR 7978, 06 July 1996). From here, we get the average value of the rate of rise of the magnetic tube from the maximum (for 2003 October) to the minimum (for July 1996)
$\frac{1.0(2003)+0.30(1998)+0.30(1998)+0.25(1996)}{4}km/s\sim 0.45~km/s$ (38)
which due to our mean velocity uncertainty of $45_{-0.15}^{+0.15}~km/s$ almost exactly corresponds to the known uncertainty in the average velocity from 0.3 to 0.6 km/s (S. Ilonidis et al., 2011)!
In other words, this means that our (maximum (31) and (36)) and minimum ((34) and (37)) are the remarkable existence of a maximum and minimum of 11-year variations of the photosphere magnetic flux tube for sunspots, which, unlike the solar dynamo, are predetermined as flux tube magnetic field sources (anchored in the overshoot tachocline), and are coupled by the holographic mechanism of quantum gravity (see Fig. 2b), and anticorrelated 11-year ADM density variations gravitationally trapped in the interior of the Sun (see Section 3 in (V. D. Rusov et al., 2021)).
In order to show signatures of appearing areas of sunspots inside the Sun before they appear on the surface, we used, according to A. G. Kosovichev et al. (2016), the measurements of plasma flows in the upper convection zone, provided by the Time-Distance Helioseismology pipeline developed for the analysis of solar oscillation data obtained by the Helioseismic and Magnetic Imager (HMI) aboard the Solar Dynamics Observatory (SDO), to investigate the subsurface dynamics of the emerging active region (AR) NOAA 11726.
Figure 7c shows the distribution of the effective phase shift (“helioseismic index”) for the depth range 62-75 Mm measured on April 19, 2013, 03:00 UT, when there was no significant magnetic flux on the surface. The figure 7d shows the variations of the helioseismic perturbation associated with the emerging AR at two different depths: 62-75 Mm and 42-55 Mm depending on time. One of the greatest advantages of the remote helioseismology is that the full-disk observations of solar variations allow to build a special scheme of measurement and selection of acoustic waves signals (Fig. 7b), passing though certain sub-surface areas. For example, S. Ilonidis et al. (2013) developed a special procedure of deep focusing which is able to detect large emerging active areas more than a day before their appearance on the surface (Fig. 7). The rise velocity determined by tracking the magnetic flux declination with depth, is about 1.4 km/s, which is very close to the rise velocity in the deep layers A. G. Kosovichev et al. (2016), e.g. near tachocline! Interestingly, the results of several attempts to detect the emerging magnetic flux before its appearance in the photosphere may be compared to our theoretical estimates (maximum (31) and (36)) and minimum ((34) and (37)) and demonstrate a surprising agreement in terms of the magnetic tube rising velocity and time.
Figure 7: Detection of emerging active region AR 11726: a) magnetograms of helioseismic and magnetic imager (HMI) on board of the Solar Dynamics Observatory (SDO) of April 19, 2013, 3:00 UT, prior the emergence; the rings show the area where the acoustic oscillation signal was measured to detect the subsurface signal at the central point; b) a vertical cut through the computational grid used in the time-distance inversions, and a sample of ray paths of acoustic waves, which were used for measuring travel times (A. G. Kosovichev & J. Duvall, 2008; T. L. Duvall et al., 1993, 1996); c) variations of a travel-time index showing the AR perturbation located the depth of 62-75 Mm on April 19, 2013, 3:00 UT; d) variations of the helioseismic perturbation associated with the emerging AR at two different depths: 42-55 Mm and 62-75 Mm, as a function of time. The characteristic rise time estimated using the delays in two layers at depths 42-55 Mm and 62-75 Mm is about 1.4 km/s. Derived from A. G. Kosovichev et al. (2016).
Below we briefly and simply show how our fundamental physics of the rise of strong magnetic flux tubes with $B\gg 10^{5}~G$, but without dynamo action, differs from the classical physics of an active region, in which the rise of magnetic tubes from a tachocline to the photosphere is determined by the result of the convective action of the dynamo in the main part of the solar convection zone (with 0.15-0.25 km/s; see e.g. Y. Fan (2021); P. Charbonneau (2020); P. Charbonneau & D. Sokoloff (2023) and Fig. 21 of H. Hotta & H. Iijima (2020).
Let us first draw some attention to what the remarkable physicist Yuhong Fan says about the essence of magnetic fields in the solar convection zone: “The question of whether, or to what extent, a strong toroidal magnetic field stored in the overshoot region at the base of the convection zone, generated by a deep seated solar dynamo process, is responsible for the formation of solar active regions remains to be more rigorously investigated.(Y. Fan, 2021)
Recalling the words of Yuhong Fan that “it remains to be more thoroughly investigated” to what extent is strong toroidal magnetic field, control by means of the self-generating dynamo in the overshoot tachocline region, one must also remember the existence of the amazing 11-year variation of dark matter density (see e.g. V. D. Rusov et al. (2021)), and, for example, the holographic mechanism of quantum gravity (see S. W. Hawking (2015b, a); S. W. Hawking et al. (2016); V. D. Rusov et al. (2021)), which strongly influence the Sun, the stars, the Universe?! This means that, unlike the holographic mechanism of quantum gravity on the Sun, where the magnetic flux tubes in strong fields ($\sim 4.1\cdot 10^{7}~G$ in the tachocline are identical to the magnetic core of the Sun), such physics of formation, emergence and “decay” of sunspots, which is considered an integral part of the convective dynamo (see e.g. Y. Fan & F. Fang (2014) and Fig. 21 in H. Hotta & H. Iijima (2020)), and, as a consequence, the formation of active regions from the uppermost convection zone to the corona (see e.g. F. Chen et al. (2017, 2022)), is still an unsolved problem of the 70-year evolution and the current state of solar and stellar dynamo theory (see e.g. P. Charbonneau & D. Sokoloff (2023) and P. Charbonneau (2020)), which, although good, cannot be on the Sun, and certainly not in our Universe!
In contrast to the classical physics of rise of magnetic flux tubes from the tachocline to the photosphere, which are driven by a self-generating dynamo in the main part of the solar convective zone, we already know that,
  • first, a strong toroidal magnetic field stored in the overshoot tachocline region is determined by the thermomagnetic Ettingshausen-Nernst effect (see Apendix A in V. D. Rusov et al. (2021)), which exactly (in contrast to dynamo!) coincides with $B_{tacho}\sim 10^{7}~G$. In this case, we indirectly showed that, using the fundamental holographic principle of quantum gravity (see Apendix C in V. D. Rusov et al. (2021)), the repelling toroidal magnetic field of the tachocline exactly “neutralizes” the magnetic field in the Sun’s core (see Eq. 1 and W. A. Fowler et al. (1955); S. Couvidat et al. (2003)), in which the projections of the magnetic fields of the tachocline and the core have the equal value but the opposite directions.
  • second, we underline that this process is associated with continuous transformation of toroidal magnetic field into poloidal magnetic field ($T\rightarrow P$ transformation), but not vice versa ($P\rightarrow T$). It means that the holographic mechanism is the main process of regeneration of the primary toroidal field in the tachocline, and the formation of the processes of these fields shears at the base of the convective zone leads to the emergence of buoyancy of toroidal magnetic tubes rising to the surface of the Sun.
  • third, our physics of the magnetic flux tubes rise is related to the fact that if the axions, which are born in the Sun’s core, are directly converted into X-rays near the tachocline, then the axion-photon oscillations predetermine the appearance of magnetic sunspot cycles. This is due to the fact that the formation of sunspots and their cycles is a consequence, according to V. D. Rusov et al. (2021), of the anticorrelation 11-year cycles of ADM density modulation inside the Sun, which, not surprisingly, are a consequence of the 11-year ADM halo density modulation in the fundamental plane of the galactic center, which closely correlates with the density modulation of baryonic matter near the supermassive black hole. From here, it is easy to show (see V. D. Rusov et al. (2021)) in what way the anticorrelation identity between indicators of the ADM density modulation inside the Sun and the number of sunspots (or the correlation identity between indicators of modulation of solar axions (or photons of axion origin) and sunspot cycles) is realized.
Finally, let us show that if the rise velocity of the buoyant magnetic tube, determined by the magnetic field values maximum ($B_{max}\sim 3.6\cdot 10^{7}~G$) and minimum ($B_{max}\sim 3.6\cdot 10^{6}~G$) of the magnetic field and the width of the “ring” $(a_{conv})_{max}\sim 3.7\cdot 10^{-4}~H_{p}$ or $(a_{conv})_{min}\sim 3.7\cdot 10^{-3}~H_{p}$, causes the MFT to appear on the surface in the form of sunspots, then the parameters of the universal vanBFF model will be associated with the magnetic cycles that are directly related to the observed features of the cycle of both double sunspot maxima (see Sect. 2.2.2) and the tilt angle of Joy’s law (see Sect. 2.2.1).
2.2 Physics of primary and secondary reconnection of magnetic flux tubes in the lower layers of the tachocline
The problem is devoted to the physics of primary and secondary magnetic reconnection in the lower layers of the tube, which is directly related to the observed features of the cycle of both double maxima of sunspots and the tilt angle of Joy’s law.
First, let us recall the essence of the primary reconnection inside magnetic tubes near the tachocline ($B\sim 10^{7}~G$), which generates the O-loop (see Fig. 4d), and thus participates in the formation of axion origin photons (see Fig. 3, left) and axions of photon origin (see Fig. 3, right) from the O-loop to the photosphere, is the source of sunspots, which (due to the cooling of the Parker-Biermann effect inside the tube) rise from the tachocline to the surface of the Sun.
On the other hand, we are interested in the essence of the very important secondary magnetic reconnection (see Fig. 5b), which is the source of the observed double maxima of one sunspot cycle – it shifts by 2.5 years between the primary and secondary solar maxima when the magnetic tube rises, and is associated with both the observational data on the tilt angle of Joy’s law and the “disappearance” of sunspots on the solar surface.
In Fig. 8 one can see the conditions for the secondary reconnection between the O-loop (green lines) and the unipolar part of the $\Omega$-loop (blue lines) that can organize them in the lower layer (Fig. 8, left) or in the upper layer (Fig. 8, right), thereby showing the appearance of bipolar magnetic tubes in various versions (Fig. 8c,d and g,h).
Figure 8: Topological effects of magnetic reconnection in the lower (left) or upper (right) layers of the magnetic tube. Here, the unipolar part of the $\Omega$-loop is rebuilt on its base, compressing the $\Omega$-loop (blue lines) to form a free O-loop (green lines) (Fig. 5a). The yellow lines show the movement of the substance leading to the connection of the “leg” loop (see analogous Fig. 4 in E. N. Parker (1994)). If the O-loop (primary magnetic reconnection, green lines) can randomly have different directions of magnetic fields, then the secondary magnetic reconnection can generate loop “legs” in different layers, for example, in lower layers (Fig. 8c and also Fig. 8d as an analog of Fig. 5b) and upper layers (Fig. 8e and also Fig. 8f as an analog of Fig. 4 in S. Jabbari et al. (2016)).
From here, since the sunspot cycle is identical to the maxima and minima of the magnetic tube number on the surface of the Sun, we are interested in the physics of the maxima and minima both during the primary reconnection for the time $(\tau_{d})_{conv}$ of the rise of the almost empty magnetic tube from the tachocline to the solar photosphere, which are associated with the features of the tilt angle of Joy’s law (see Fig. 10c), and during the secondary reconnection for the time $(\tau_{n})_{rec}$ of the departure of the same rise of the almost empty magnetic tube from the surface of the Sun and, as a consequence, their “disappearance” on the Sun.
2.2.1 On the ascending magnetic tube and the formation of sunspots both near the tacho-cline with $\sim 10^{7}~G$ in the absence of Joy’s law and in the middle region with $\sim 10^{5}~G$ in the presence of Joy’s law
The first part of the scenario consists in discussing the physics of magnetic flux tube buoyancy at strong fields and with the participation of axion origin photons (see Fig. 3), associating with the formation of an O-loop (see the existence of primary magnetic reconnection in Fig. 4) inside the magnetic flux tube near the tachocline.
For the neutral buoyancy condition ($\rho_{int}=\rho_{ext}\approx 0.2~g/cm^{3}$; see Figs. 3 and 6) and the strong toroidal field of the magnetic tube, $B_{Sun}=4.1\cdot 10^{7}~G$, as well as the average width of the “thin” ring $a_{conv}\sim 3.7\cdot 10^{-4}~H_{p}$ (see Eq. (22) and Fig. 6) between the O-loop and the walls of the magnetic tube (see Fig. 6), it is easy to show the estimate of the time $(\tau_{d})_{conv}$ (see Eqs. (27)-(34)) and the velocity $(v_{rise})_{conv}$ (see Eq. (16) and also Eqs. (36)-(37) of the convective rise from the tachocline to the photosphere, which for such large magnetic fields has a significant amount of rise of the magnetic tubes at all latitudes (see Fig. 5a and the red lines in Fig. 9a).
Figure 9: Scheme of turbulent reconstruction of the toroidal magnetic field in the convective zone (CZ): (a) meridional circulation (closed blue poloidal field lines), which is the source of the toroidal field in the tachocline (thin black lines) based on the holographic mechanism of quantum gravity (see Fig. 2b), and the magnetic buoyancy of the flux tubes (red arrows) with a magnetic field of $\sim 10^{7}~G$; (b) In the CZ, through secondary turbulent reconnection of the magnetic tube near the tachocline (see Fig. 10b2,c), effects of both macroscopic turbulent plasma diamagnetism are observed, facilitating the formation of two layers on the surface (see analogous (V. N. Krivodubskij, 2005, 2021; V. N. Krivodubskii, 1992)) with different vertical directions of horizontal field transfer (see “black” arrows), and the conditions of magnetic field transfer along $\nabla\rho$ due to small-scale magnetic pulsations (the so-called magnetic pumping, see “white” arrows), acting against buoyancy. However, in the deep layers in the near-equatorial region, the $\nabla\rho$ effect provides upward advection, which, on the contrary, directs magnetic buoyancy, and thus, can facilitate the penetration of strong fields to the surface.
This solution clearly depends on the rise time of the magnetic tubes rising from the tachocline to the solar surface. Therefore, the primary magnetic reconnection itself in the lower layers of the magnetic tubes (see Fig. 5a), is not the final stage of the modeling. The essence of a practically empty tube in strong fields of $\sim 10^{7}~G$ is related (through the primary magnetic reconnection) to the physics of the secondary turbulent reconnection of magnetic bipolar structures (see Fig. 5b, and also Fig. 4 in (A. S. Brun et al., 2011)). Here, on its basis, the bipolar part of the $\Omega$-loop is reconstructed, compressing the $\Omega$-loop (blue lines: see Fig. 5b,c) near the tachocline and simultaneously lifting them (from 2 to 3 years) to the surface of the Sun, and, as a consequence, after some time (from several days to several months (see S. K. Solanki (2003); K. Petrovay & L. van Driel-Gesztelyi (1997)) a free O-loop appears (blue lines in Fig. 5d), which eventually “disintegrates” or, more precisely, the O-loop (without spots) flies away from the surface of the Sun (see Fig. 5d).
This raises a rather unexpected but important question: How is the secondary magnetic reconnection of flux tubes with $B_{tacho}^{Sun}=B_{z}=4.1\cdot 10^{7}~G$, leading to a real sharp decrease in the magnetic field to $B_{eq}\leqslant 2\cdot 10^{5}~G$ at $\sim 0.80R_{Sun}$ (see e.g. Fig. 10b,c), related to the existence of downward turbulent diamagnetic pumping and rotational $\nabla\rho$-pumping at relatively low magnetic fields, and as a consequence, at low escape velocities of flux tubes from the solar surface?
In this regard, we obtain an answer to this important question. For this purpose, we consider the second part of the scenario, based on the remarkable idea of L. L. Kichatinov (1991), which includes the generation of the magnetic field of $\leqslant 2\cdot 10^{5}~G$ near the bottom of the convective zone (see Fig. C.1b in V. D. Rusov et al. (2021) for the holographic mechanism) and transfer of the toroidal field from the deep layers at different latitudes.
It is very important that the efficiency of the magnetic buoyancy transfer is associated with the participation of two processes: macroscopic turbulent diamagnetism (Y. B. Zel’dovich, 1957, 1956; L. J. Spitzer, 1957) and rotational $\nabla\rho$-pumping (L. L. Kichatinov, 1991).
Let us note some important properties of macroscopic turbulent diamagnetism. It is known that Y. B. Zel’dovich (1956, 1957) and L. J. Spitzer (1956) discovered the diamagnetism of inhomogeneously turbulent conducting liquids, in which the inhomogeneous magnetic field moves as a single whole. In this case the turbulent fluid, for example, with nonuniform effective diffusivity $\eta_{T}\approx(1/3)vl$ (see Fig. 1 in  L. L. Kitchatinov & G. Rüdiger (2008), and also (Y. B. Zel’dovich, 1957, 1956; L. J. Spitzer, 1957)) behaves like a diamagnetic one and carries the magnetic field with the effective velocity (H. Köhler, 1973)
$\vec{v}_{dia}=-\frac{1}{2}\nabla\eta_{T},$ (39)
where $l$ is the mixing length of turbulent pulsations, and $v=\sqrt{\langle v^{2}\rangle}$ is the root-mean-square velocity of turbulent motion. The minus sign on the right in Eq. (39) shows the meaning of turbulent magnetism: it is not paramagnetic magnetism, so magnetic fields repel from regions with relatively high turbulent intensity. In other words, macroscopic turbulent plasma diamagnetism and, as a consequence, the so-called macroscopic diamagnetic effect (see K.-H. Rädler (1968a, b)) in the physical sense is the displacement of the averaged magnetic field $B$ from regions with increased intensity of turbulent pulsations to regions with less developed turbulence (F. Krause & K.-H. Raedler, 1980; L. L. Kichatinov & G. Ruediger, 1992).
However, there is an interesting problem of the diamagnetic process caused by inhomogeneous turbulent intensity with allowance for the total nonlinearities in the magnetic field. This is due to the fact that up to the present time analytical estimates have been obtained only for the limiting cases of weak and strong magnetic fields. For example, at strong magnetic fields of flux tubes, at $B>10^{5}~G$, e.g. $\geqslant 10^{7}~G$, the diamagnetic effect becomes almost negligible, in particular, strong magnetic damping of diamagnetism $\sim B^{-3}$ is obtained for super-equipartitions of fields (F. Chen et al., 2022) when turbulence is close to two-dimensional (Y. B. Zel’dovich, 1957). On the other hand, for very weak fields the diamagnetic pumping, which is predetermined by the intensity of turbulence at $B<10^{5}~G$, is a very effective process (F. Chen et al., 2022).
Among the known limiting cases of weak and strong magnetic fields from the tachocline to the surface, we are interested in the addition to the portions of strong fields (via the secondary Lazaryan-Vishniac reconnection (see A. Lazarian & E. T. Vishniac (1999); A. Lazarian et al. (2004); G. Kowal et al. (2009), and also §III and §IX in (A. Lazarian et al., 2020)) near the tachocline, which have a very slow Sweet-Parker velocity $(V_{rec})_{2}$ (see E. N. Parker (1957); P. A. Sweet (1958)) for Fig. 10b2), the part of the weak magnetic field from $0.8~R_{Sun}$ to near $0.85~R_{Sun}$ (see Fig. 10b2,c), which constitutes the super-equipartitions of fields $B\geqslant B_{eq}$ and is the source of two processes: diamagnetic pumping, and rotational $\nabla\rho$-pumping. This is due to the fact that the secondary magnetic reconnection of the flux tubes (see Figs. 8b,d,f,g) leads to the real increase in the upward magnetic field to $B\approx 2\cdot 10^{5}~G$ at $\sim 0.8~R_{Sun}$, which is identical to the $1/\beta=B^{2}/8\pi p_{ext}\approx 2.9\cdot 10^{-5}$ (see e.g. Eq. (9), and Fig. 10b2,c). This means that the real toroidal magnetic flux tube is a consequence of the formation of the secondary reconnection, which generates both the “anti-buoyant” effects of downwardly directed turbulent diamagnetic transport and the rotational effect of the magnetic $\nabla\rho$-pumping at polar latitudes, and the “buoyant” effects of magnetic flux tubes at equatorial latitudes as a result of the upward rotational effect of magnetic $\nabla\rho$-pumping.
Figure 10: An alternative theory of magnetic flux tubes in strong fields by means of axion origin photons associated with primary and secondary reconnections.
Figure 11: (cont.) An alternative theory of magnetic flux tubes in strong fields by means of axion origin photons associated with primary and secondary reconnections.
In this regard, we consider turbulence with quasi-isotropic spectral tensor (L. L. Kichatinov, 1987, 1991), which is certainly the simplest representation for inhomogeneous turbulence (see Eq. (2.12) in L. L. Kichatinov & G. Ruediger (1992)). As a consequence, the information on the spectral properties of turbulence (given by Eqs. (2.12)-(2.16) from L. L. Kichatinov & G. Ruediger (1992)) is sufficient to reduce the expression for the average electromotive force $\varepsilon$ (see Eq. (2.1) L. L. Kichatinov & G. Ruediger (1992)) to its traditional form, where only integrations over the wave number $k$ and frequency $\omega$ remain. After such shortening it is possible to get (see Eq. 3.1 in L. L. Kichatinov & G. Ruediger (1992))
$\vec{F}=(\vec{v}_{dia}+\vec{v}_{dens})\times\vec{B}$ (40)
with the speed of turbulent diamagnetic transfer
$\vec{v}_{dia}=-\nabla\int\limits_{0}^{\infty}\Re_{dia}(k,\omega,B)\frac{\eta k^{2}q(k,\omega,x)}{\omega^{2}+\eta^{2}k^{4}}dkd\omega,$ (41)
where $q$ stands for the local velocity spectrum, and the rate of rotational magnetic advection caused by the vertical heterogeneity of the fluid density in the convective zone, i.e. the magnetic $\nabla\rho$-pumping effect,
$\vec{v}_{dens}=\frac{\nabla\rho}{\rho}\int\limits_{0}^{\infty}\Re_{dens}(k,\omega,B)\frac{\eta k^{2}q(k,\omega,x)}{\omega^{2}+\eta^{2}k^{4}}dkd\omega.$ (42)
The effective speeds $v_{dens}$ and $v_{dia}$ are consequences of the non-uniformity of density and of turbulence intensity, respectively, where the latter is attributed to the known diamagnetic pumping. The velocities (41) and (42) depend on the magnetic field through the kernels $\Re_{dens}$ and $\Re_{dia}$.
We are interested in the problem of reconstructing a strong toroidal field of flux tubes $\sim 10^{7}~G$ (see Figs. 9a,b with the magnetic tube profile, and Figs. 10a,b,c with the magnetic tube in full face), transform regions of the average magnetic field $B\sim 2\cdot 10^{5}~G$ in the convective zone (see Figs. 10b,c), and thereby allow the organization of the amazing balance between the magnetic buoyancy, turbulent diamagnetism, and the rotationally modified $\nabla\rho$-effect.
In this regard, we first consider the physics of linear magnetic diffusivity, which is known to govern two processes: macroscopic turbulent diamagnetism and rotational $\nabla\rho$-pumping. This is because the standard theory of the turbulent mixing length $l$ and the characteristic velocity of the dominant eddies $u$ yields a diffusion coefficient $\eta_{T}=(1/3)ul$, that depends directly on the diamagnetic pumping rate $v_{dia}=-\nabla\eta_{T}/2$ and, indirectly, on the rotational $\nabla\rho$-pumping $v_{dens}$.
Here we assumed that, according to Figs. 10b2,c, the radial profile of turbulent diffusion is a smooth-convex function with a maximum, $\eta_{T}\approx 10^{11}~cm/s$, approximately in the middle of the CZ at a depth of $z=1.30\cdot 10^{5}~km$ at $\sim 0.8R_{Sun}$, which is described by the distance of $2.3\cdot 10^{4}~km$ between $0.8R_{Sun}$ and near $0.85R_{Sun}$ (see Figs. 10b2,c) and with the depth of the bottom of the convective zone $z_{0}\approx 1.84\cdot 10^{5}~km$ (see M. Stix (2002); R. Kippenhahn et al. (2012); J. Schumacher & K. R. Sreenivasan (2020)).
In order to consider the balance of the three velocities $v_{B}$, $v_{dia}$, and $v_{dens}$, it is necessary to apply the widely used approximation of the mixing length (see e.g. E. Böhm-Vitense (1958); P. Bradshaw (1974); D. O. Gough (1977); D. Gough (1977); A. J. Barker et al. (2014); A. Brandenburg (2016)), which, according to L. L. Kichatinov (1991), fully satisfies this goal. This approximation will be understood as the replacement of nonlinear terms along with time derivatives in the equations for fluctuating fields by means of $\tau$-relaxation terms, i.e. instead of equations (3.1) and (3.9) from L. L. Kichatinov (1991), we now have the equation of the radial speed of the toroidal field in the convection zone
[…]
and firstly, it can be shown that at $\eta_{T}\approx 10^{11}~cm/s$, we have
$v_{dens}^{light}\sim 3\times\eta_{T}\times g_{0.8R_{Sun}}/(\gamma-1)c_{p}T_{0.8R_{Sun}}\approx 4.1\cdot 10^{1}~cm/s,$ (46)
where $\tau\approx l/(\langle u^{2}\rangle^{\circ})^{1/2}$ is a typical lifetime of a convective eddy; $l$ is the mixing length; $\langle u^{2}\rangle^{\circ}$ is the mean intensity of fluctuating velocities for original turbulence; $\theta$ is the latitude (scalar); $e_{r}\nabla\rho/\rho=-e_{r}g/[(\gamma-1)c_{p}T]$ , where $e_{r}$ is the radial unit vector; $T\cong 1.352\cdot 10^{6}~K$ is the temperature at $0.8R_{Sun}$ (J. Bahcall & M. Pinsonneault, 1992); $g=g_{0}(R_{Sun}/r)$ is the gravity, where $g_{0}=2.74\cdot 10^{4}~cm/s^{2}$ is the surface value of solar gravity, and $g_{0.8R_{Sun}}=4.22\cdot 10^{4}cm/s^{2}$ at $0.8R_{Sun}$; $c_{p}=3.4\cdot 10^{8}~cm^{2}s^{-2}K^{-1}$ (fully ionized hydrogen) is the specific heat at constant pressure; $\gamma=5/3$ is the ratio of specific heats $c_{p}/c_{V}$; $3\kappa=\tau\langle u^{2}\rangle^{\circ}$, where $\kappa\equiv\eta_{T}\approx 10^{11}~cm^{2}/s$ is turbulent diffusivity coefficient, and the mixing length relation $\langle u^{2}\rangle^{\circ}=-\nabla\Delta Tl^{2}g/(4T)$, where the $\nabla\Delta T$ is superadiabatic temperature gradient; $v_{p}=(1/6)\tau\langle u^{2}\rangle^{\circ}(\nabla\rho/\rho)$ is the velocity of the magnetic field transfer caused by the density gradient (see Eq. (36) in S. I. Vainshtein & L. L. Kichatinov (1983)); $\hat{\Omega}=Co=2\tau\Omega$ is the Coriolis number (reciprocal of the Rossby number), where $\Omega$ is the rotation speed, $\tau$ is the turnover time, and the functions
$\phi_{n}(\hat{\Omega})=(1/8)I_{n}(\Omega,k,\omega)$ (47)
(see $I_{1}$ and $I_{2}$ in Eqs. (3.12) and (3.21) in L. L. Kichatinov (1991)) are
$\phi_{1}(\hat{\Omega})=\frac{1}{4\hat{\Omega}^{2}}\left[-3+\frac{\hat{\Omega}^{2}+3}{\hat{\Omega}}\arctan\hat{\Omega}\right$], (48)
$\phi_{2}(\hat{\Omega})=\frac{1}{8\hat{\Omega}^{2}}\left[1+\frac{\hat{\Omega}^{2}-1}{\hat{\Omega}}\arctan\hat{\Omega}\right$] (49)
(see also analogous Eq. (19) and Fig. 2 in L. L. Kitchatinov & A. A. Nepomnyashchikh (2016)), which describe the rotational effect on turbulent convection.
Secondly, we get diamagnetic pumping rate
$v_{dia}=-\nabla\eta_{T}/2\approx 1.7\cdot 10^{2}~cm/s.$ (50)
Thirdly, in addition to the radial profile of turbulent diffusion $\eta_{T}\approx 10^{11}~cm^{2}/s$, which is related to the diamagnetic pumping velocity $v_{dia}\approx 1.7\cdot 10^{2}~cm/s$ and the distance $2.3\cdot 10^{4}~km$ between $0.8R_{Sun}$ and near $0.85R_{Sun}$ (see the “green” line of the turbulent-toroidal magnetic field in Fig. 10b,c), we need to calculate the magnetic buoyancy velocity $v_{B}$, which is associated with a super-uniformly distributed field strength $B\sim 2\cdot 10^{5}~G$ (see Fig. 9b and Fig. 10b,c). It should be remembered that for buoyant flux tubes with super-equipartition field strength $B\sim 2\cdot 10^{5}~G\geqslant(H_{p}/a)^{1/2}B_{eq}\approx 3.16B_{eq}$ with $a\approx 0.1H_{p}$, where, as Yuhong Fan wrote (see Section 7 in (Y. Fan, 2021)), $B_{eq}\approx 6.1\cdot 10^{4}~G$ is in equipartition with the kinetic energy density of the convective downflows, the magnetic buoyancy of the tubes dominates the hydrodynamic force from the convective downflows and the flux tubes can rise unaffected by convection”, but significantly affects the turbulent diamagnetic pumping!
For this, using our so-called universal van Ballegooijen-Fan-Fisher model with equations (8)-(12) with the values of the magnetic flux tube with $\sim 2\cdot 10^{5}~G$, the pressure of $1.665\cdot 10^{13}~erg/cm^{3}$ at $0.8R_{Sun}$ and the rate of radiative heating $(dQ/dt)_{1}\approx 29.7~erg\cdot cm^{-3}\cdot s^{-1}$, we obtained the magnetic buoyancy velocity $v_{B}\approx 1.7\cdot 10^{2}~cm/s$, which, surprisingly, exactly coincides with the diamagnetic pumping velocity $v_{dia}\approx 1.7\cdot 10^{2}~cm/s$.
$v_{B}\approx 1.7\cdot 10^{2}~cm/s,$ (51)
which, surprisingly, exactly coincides with the diamagnetic pumping rate $v_{dia}\approx 1.7\cdot 10^{2}~cm/s$.
From here, a certain balance of three velocities, $\vec{v}_{B}$, $\vec{v}_{dia}$ and $\vec{v}_{dens}$, arises, which at $v_{B}=v_{dia}\approx 1.7\cdot 10^{2}~cm/s$ provide both the process of blocking (via $\vec{v}_{dens}$ at $0.8R_{Sun}$) magnetic buoyancy at high latitudes
$\uparrow v_{B}+\downarrow v_{dia}^{black}+\downarrow v_{dens}^{light}=\downarrow v_{dens}^{light}\cong\downarrow 4.1\cdot 10^{1}~cm/s,$ (52)
and the process of raising magnetic buoyancy (via $\vec{v}_{dens}$ at $0.8R_{Sun}$) to the surface of the solar photosphere at low latitudes
$\uparrow v_{B}+\downarrow v_{dia}^{black}+\uparrow v_{dens}^{light}=\uparrow v_{dens}^{light}\cong\uparrow 4.1\cdot 10^{1}~cm/s.$ (53)
One of the main tasks was related to the solution of the maximum near-equatorial time $(\tau_{n})_{max}$ of the rise of the magnetic flux tube in the region between $0.8R_{Sun}$ and near $0.85R_{Sun}$ (see Fig. 9b and Fig. 10b,c), which is associated with the distance of turbulent diffusion $l_{diffus}=2.3\cdot 10^{4}~km$ and the speed of rotational $\nabla\rho$-pumping $v_{dens}\approx 4.1\cdot 10~cm/s$:
$(\tau_{n})_{max}$ $=\frac{l_{diffus}}{v_{dens}}\approx 0.56\cdot 10^{8}~s\sim 2~years,$
$(\tau_{n})_{min}$ $=\frac{l_{diffus}}{v_{dens}}\approx 0.56\cdot 10^{6}~s\leqslant 25~days.$ (54)
When the toroidal fields of magnetic tubes rising from the tachocline to the solar surface, but without the participation of the turbulent diffusion distance (see Eq. 39), have strong fields of the order of $2\cdot 10^{5}~G$ (see Fig. 9a and Fig. 10a), then they
$v_{B}(>2\cdot 10^{5}~G)\gg v_{dia}+v_{dens},$ (55)
suppress the existence of the velocities $\vec{v}_{dia}$ and $\vec{v}_{dens}$! This means that the near-equatorial time $(\tau_{n})_{max}$ of the magnetic flux tube rise, equal to 2 years, neglects the time of convective heating of the flux tube, which corresponds to only a few days (see analogue of $(\tau_{d})_{conv}^{max}$ in Eq. 31 and $(\tau_{d})_{conv}^{min}$ in Eq. 34).
The most interesting thing is that, in addition to the sunspots after the rise of the magnetic tube with $B\sim 10^{7}~G$ at all latitudes (see Fig. 9a and Fig. 10a) and the sunspots after the rise of the magnetic tube with $B\leqslant 2\cdot 10^{5}~G$ only at low latitudes (see Fig. 9b and Fig. 10c), which are the source of the double maxima of the 11-year solar cycles (see the full experimental evidence in Section 2.2.2), we are interested in the problem or, more precisely, the physics of the tilt angle of Joy’s law (V. G. Ivanov, 2012; M. Dasi-Espuig et al., 2010) and how it is related to the magnetic $\nabla\rho$-pumping?
First, about the Coriolis force. According to Käpylä, P. J. et al. (2014), the rotational effect on the flow can be measured with the local Coriolis number $\hat{\Omega}\equiv Co=2\tau\Omega$, especially if $\tau$ is estimated on the basis of the theory of mixing length, which predicts values of $\hat{\Omega}$ reaching more than 10 in the deep layers (see e.g. M. Ossendrijver (2003); A. Brandenburg & K. Subramanian (2005); P. Käpylä (2011); P. Käpylä et al. (2011)). However, on the other hand, according to Käpylä, P. J. et al. (2014), the question arises of whether or not there are differential rotation bistabilities in which antisolar (stellar-like) and/or solar-like rotation depends on the initial state or launch history, or, in addition, the question of whether there are such deep layers of solar and stellar convection zones or not is still open.
In contrast to the work of Käpylä, P. J. et al. (2014), already in 2021, astrophysicist Krivodubskii (V. N. Krivodubskij, 2005, 2021) recalled (see page 72 in (V. N. Krivodubskij, 2005)) that exactly “…Modern helioseismologocal inversions (J. Schou et al., 1998; R. Howe et al., 2000; H. M. Antia, 2002; S. Basu & H. M. Antia, 2003; S. Basu & H. Antia, 2008) show that the differential rotation rate, $\Omega(r,\theta)$, in the SCZ can be divided into domains of fast (equatorial) and slow (polar) rotation. And, what is more important, the radial angular-velocity gradient $\partial\Omega/\partial r$ has opposite signs in these domains. The angular velocity increases along the radius, $\partial\Omega/\partial r>0$, at the near-equator region, and, in contrast, it decreases along the radius, $\partial\Omega/\partial r<0$, at the polar regions.” The modernity of this considered helio-seismological inversion remains to this day!
At the same time, we will show a little later that this rotating magnetic $\nabla\rho$-pumping due to the second part (slow) secondary reconnection of the $\Omega$-loop (see $(V_{rec})_{2}$ in Figs. 10b2,c) (through the primary part (fast) secondary reconnection of the $\Omega$-loop (see $(V_{rec})_{1}$ in Fig. 10b1, also in Fig.5a-d, Fig. 8b,d, and also Fig. 8c)) both in the polar region of slow rotation, where the angular velocity increases inward of the Sun ($\partial\Omega/\partial r<0$), in the equatorial region of fast rotation, where the angular velocity mainly decreases with depth ($\partial\Omega/\partial r>0$). Moreover, the direction of the magnetic $\nabla\rho$-pumping (up or down, see Fig 10b2,c) is susceptible to the sign of the factor in Eqs. (52)-(53), which depends on the polar angle (colatitude) and the behavior of the Coriolis number functions in the convection zone.
Below we show that the estimate of the Coriolis number for solar convection in the deep layer should be $\hat{\Omega}\approx 20$. Due to Eqs. (47)-(49) and the function $\phi_{n}$ for the quantity $\hat{\Omega}\approx 20$, the following values are adopted:
$\phi_{1}\cong 0.0171,~~~\phi_{2}\cong 0.0098,$ (56)
at which the radial velocity $v_{dens}^{light}$ in Eq. (43) of toroidal field transport changes the sign at the latitude $\theta^{*}=\arccos\sqrt{\varphi_{2}/\varphi_{1}}\cong 41^{\circ}$, being negative (downward) for $\theta>\theta^{*}$ and positive (upward) for $\theta<\theta^{*}$. Using (45), we find that the value of the radial velocity in the convective zone $v_{dens}^{light}$ (see Eqs. (43)-(45)) near low latitudes (e.g. $\theta^{*}=\arccos(0.985)\cong 10^{\circ}$; see also Fig. 9b) almost completely coincides with the value of the speed (46), which was previously calculated using the toroidal field $B\approx 2\cdot 10^{5}~G$ at $0.8R_{Sun}$ (see Eq. (9) and Fig. 10b,c) in the $\Re_{dens}(\beta,\varphi)$ (see Eq. (3.13) in (L. L. Kichatinov & G. Ruediger, 1992)).
From here, this raises an unexpected question: In what way, in addition to the visible magnetic tube at strong fields at all latitudes (see Fig. 9a and Fig. 10a), the “invisible” magnetic tube at high latitudes (see Fig. 10b2) and the “visible” magnetic tube at low latitudes (see Fig. 10c), depending on the time $(\tau_{rec})_{2}\sim 2.5~years$ of reconnection of the horizontal field in the tachocline (which is equal to the sum of $(\tau_{n})_{max}\approx 2~years$ and the lifetime of sunspots ($\leqslant 0.5~years$) from several days to several months (see S. D’Silva & A. R. Choudhuri (1993); Y. Fan et al. (1993, 1994)) surprisingly, practically does not depend on the direction in the radial plane down or up?
The answer, surprisingly, is not very simple. We know that the vertical compression (see Fig. 9a and Fig. 10b1) of the descending nearly empty anchored magnetic tube in strong fields of $\sim 10^{7}~G$ (and via convective heating $(dQ/dt)_{2}$ is associated with a horizontal “stretching” or expansion of this magnetic tube near the tachocline, where the counter-movement of magnetic fields leads to the emergence of a collisionless interaction with fast (opposite) velocities $(V_{rec})_{1}$ (and times $(\tau_{rec})_{1}$ (<<10 days)) of the first part of the fast secondary (turbulent) reconnection (see Fig. 10b1). When the vertical compression leads to horizontal stronger “stretching” of the magnetic tube in weak fields of $\geqslant 2\cdot 10^{5}~G$ (and through radiative heating $(dQ/dt)_{1}$, then the counter motion of magnetic fields (see $(V_{rec})_{1}$ for Fig. 10b1) leads to the emergence of the second part of the secondary (turbulent) reconnection with a very slow Sweet-Parker speed $(V_{rec})_{2}$ (see Fig. 10b2; also A. Lazarian & E. T. Vishniac (1999); A. Lazarian et al. (2004); G. Kowal et al. (2009); A. Lazarian et al. (2020); E. N. Parker (1957); P. A. Sweet (1958)). This means that the emerging weak fields of $\geqslant 2\cdot 10^{5}~G$ at distances from $0.80R_{Sun}$ to near $0.85R_{Sun}$ are a remarkable source of rotating magnetic $\nabla\rho$-pumping with the rise time of the magnetic tube $(\tau_{n})_{max}\sim 2.0~years$ from the tachocline to the solar surface (see Fig. 10c), and the lifetime of sunspots ($\leqslant 0.5~years$), and consequently, the time of the magnetic tube evaporation $(\tau_{rec})_{2}\leqslant 2.5~years$ and their disappearance on the solar surface.
Since the study of the tendency of the tilt angle of Joy’s law is very important for understanding the evolution of the solar magnetic field, then unlike the heavy calculations of theoretical estimates of the averaged angle tilt of sunspots with increasing latitude, we can summarize our results (see Eqs. (5)-(9) in S. D’Silva & A. R. Choudhuri (1993)), which can simultaneously be expressed in the form of simple and understandable physics of Joy’s law. It is known that, according to S. D’Silva & A. R. Choudhuri (1993) and Y. Fan et al. (1993), the Coriolis force is proportional to the magnetic buoyancy velocity $v_{B}$, which can be estimated taking into account the balance between the buoyancy force (the term on the left) and the drag force (left term) and the drag force (right term):
$\frac{B^{2}}{8\pi H_{p}}=C_{D}\frac{\rho_{ext}v^{2}_{B}}{\pi a},$ (57)
where the buoyancy velocity $v_{B}$ is related to the radius $a$ of the cross-section of the magnetic tube.
Taking into account Eqs. (18) and (19) by Y. Fan et al. (1994) and using the theoretical values of the tilt angle $\alpha$, the latitude angle $\theta$ and the radial parameter $\xi=r/R_{Sun}\sim 0.80$, one can easily obtain a simplified equation in the following form:
$\sin{(tilt)}\propto\frac{\sqrt{a(\xi,\theta)}}{B}\sin{(latitude)}.$ (58)
This means that looking at equation (58), we understand that when magnetic tubes contain relatively large radii $a$ (due to radiative heat transfer $(dQ/dt)_{1}$ (see Fig. 10b2,c) and low magnetic fields ($\leqslant 10^{5}~G$), then the tilt angle of Joy’s law are born from here, and when magnetic tubes have practically zero radii $a$ (due to convective heat transfer $(dQ/dt)_{2}$ (see Fig. 10b2,c) and strong high magnetic fields ($10^{5}~G$), then the tilt angle of Joy’s law disappears!
It can be said in a completely different way. If radiative heat transfer $(dQ/dt)_{1}$ (see Fig. 10b2,c) or the presence of magnetic $\nabla\rho$-pumping is used in magnetic tubes, then, as usual, the tilt angles of Joy’s law appear in sunspots, when convective heat transfer $(dQ/dt)_{2}$ or the disappearance of magnetic $\nabla\rho$-pumping (see Fig. 10b2,c) is used in magnetic tubes, then, strangely enough, the tilt angles of Joy’s law disappear in sunspots!
In this regard, we are interested in an intriguing question: Where and how do sunspots appear, in which, according to M. N. Gnevyshev (1963), the maxima of the double peak in the sunspot cycle appear? And how does the first maximum – the maximum number of spots – physically differ from the second – the maximum of large spots?
Below we will show that, firstly, the first maximum differs from the second maximum by being shifted in time by 2 years (see text after Eq. (55)) later than the main (first!) maximum of the solar activity cycle. Secondly, according to equation (57), when the flux tube has a large radius and a low magnetic field, then this magnetic tube, surprisingly, is the source of the formation of the tilt angle (via the Coriolis force, and not necessarily on the surface of the photosphere (see Fig. 10c)) and a large area of the spot on the surface of the Sun. Conversely, when the flux tube has a small radius and a high magnetic field, then the magnetic tube, on the one hand, is not connected to the source of the formation of the tilt angle, and on the other hand, surprisingly, has a small spot area on the surface of the Sun.
2.2.2 New physics of observed double maxima of sunspot cycles and solutions of time interval or Gnevyshev gap
The phenomenon of a double peak in the sunspot cycle was discovered by M. N. Gnevyshev (1963) during a study of coronal emission and areas of active formations in the 11-year cycle 19 (M. N. Gnevyshev, 1963). First, Gnevyshev analyzed the evolution of the annual average intensity of the green coronal spectral line $\lambda 530.3$ (which serves as a direct indicator of magnetic activity and coronal heating) in five-degree heliolatitude intervals during the cycle 19 and discovered two maxima in the intensity of the corona. According to data in (A. Antalova & M. N. Gnevyshev, 1965; M. N. Gnevyshev, 1967, 1977) the first maximum of the area of sunspot groups coincides with the main maximum of the 11-year cycle for Wolf numbers, while the second maximum is associated with an increase in the number of large sunspots (the so-called maximum of sunspot intensities) (Y. I. Vitinskij et al., 1986; M. Kopecký & G. V. Kuklin, 1969). In other words, the first maximum is the maximum in the number of sunspots, and the second is the maximum of large sunspots (M. Kopecký & G. V. Kuklin, 1969). The more large spots in the cycle, the more distinct the two peaks. These features were later confirmed and for the cycle 20 (see Fig. 12), including for the solar northern and southern hemispheres separately (M. N. Gnevyshev, 1963).
Figure 12: Variation in the number of sunspots for six cycles from 19 to 24 associated with double maxima of 11-year solar cycles: (a) solid red lines of sunspot cycles – smoothed values (https://sidc.be/SILSO) Royal Observatory of Belgium 2023, August 23. (b) theoretical assessment of double maxima of 11-year solar cycles: solid line – “smoothed” blue values of sunspot cycles, in which the sum of the curves with the first maximum in the form of a “red” value and the second maximum in the form of a “green” value almost exactly coincides with a solid “blue” line of the corresponding cycle.
In contrast to the “new solar dynamo era” (see e.g. V. N. Krivodubskij (2005); Y. Fan (2021); Z. Zhang & J. Jiang (2022); Z. Zhang et al. (2023); P. Charbonneau & D. Sokoloff (2023); G. M. Vasil et al. (2024)), we used our alternative theory of magnetic flux tubes at strong fields via photons of axion origin. The essence of double maximum 11-year sunspot cycles, the source of which is the generation of a magnetic field near the bottom of the convective zone and the subsequent rise of the field from the deep layers to the surface in the photosphere, appear only at strong fields (see Fig. 9a), which generate a very “fast” (on the order of several days; see Eq. (31) $\sim 1-2~days$) first maximum (see Fig. 9a), and, as a consequence, the obligatory addition of a section at weak fields (see Fig. 10b2,c), at which it is the first maximum that is transformed into a very “slow” (on the order of two years! See $(\tau_{n})_{max}\sim 2.0~years$ second maximum (see Fig. 10b2,c)!
In other words, the very “slow” second maximum depends only on two interrelated causes: the very slow secondary reconnection (see Fig. 10b2,c) of the magnetic flux tube, which depends both at a strong magnetic field of $2\cdot 10^{5}~G$, having two sections between $0.70R_{Sun}$ and near $0.80R_{Sun}$ and between $0.85R_{Sun}$ and the photosphere (see Fig. 10b2,c), and, surprisingly, at a weak magnetic field of $2\cdot 10^{5}~G$, where the so-called effects of turbulent diamagnetic transport and rotating $\nabla\rho$-pumping appear for the first time, which provide both the process of blocking magnetic buoyancy at high latitudes (see Fig. 10b2), and at near-equatorial latitudes predetermines the process of dominance of the Coriolis force with rotating $\nabla\rho$-pumping with $\geqslant 2\cdot 10^{5}~G$ at a distance $2.3\cdot 10^{4}~km$ between $0.80R_{Sun}$ and near $0.85R_{Sun}$ (see Fig. 10c), which ultimately generates a “shift” of the arcuate ascending loop in the form of the tilt angle of Joy’s law on the solar surface (see Fig. 10c).
In this case, the key role in the proposed mechanism of double maxima is played by two “waves” of toroidal fields, in which the first wave, without the participation of secondary reconnection, first appears at all latitudes at the lower base of the CZ, and then, as a consequence, instead of the first wave without the participation of secondary reconnection, a second wave appears, which is the same primary wave, but with the participation of secondary reconnection, only at the latitudes of the equatorial surface of the Sun.
That is to say, the first maximum is the maximum of the number of spots that are sources of the magnetic flux tube (anchored in the overshoot tachocline) with $B\sim 10^{7}~G$ and with the participation of primary reconnection (see Fig. 4 and Fig. 5a; also Fig. 9a), and the second maximum is the maximum of large spots that are the result of the appearance of magnetic flux tubes with $B\sim 10^{7}~G$ and with the participation of secondary reconnection (see Fig. 5b and Fig. 10b2,c). From here, we are interested, in fact, in the very important secondary magnetic reconnection (see Fig. 9b and Fig. 10b2,c) during the rise of the magnetic flux tube in the solar photosphere, which is one of the two sources of the observed double maxima of one sunspot cycle – it shifts by $\sim 2.5~years$ between the primary (see the “red” curves in Fig. 12b) and secondary (see the “green” curves in Fig. 12b) sunspot maxima (see “smoothed” blue curves in Fig. 12b). Moreover, it is the secondary maximum of sunspots that is also associated with both observational data of the tilt angle of Joy’s law (see Fig. 10c from the right) and with the disappearance of sunspots on the surface of the Sun (see Fig. 5c,d).
And finally, we realize that the observed double maxima, for example in cycles 19-24 (see Fig. 12a), also have different ratios between the numbers of very “fast” primary maxima $(S_{n})_{max}^{primary}$ (see the red curves in Fig. 12b; also Fig. 9a and Fig. 10a) and very “slow” secondary maxima (see the “green” curves in Fig. 12b)
$(S_{n})_{max}^{primary}/(S_{n})_{max}^{secondary}\approx\begin{cases}2&at\sim 300~sunspots\\ 1.5&at\sim 150\div 200~sunspots\\ 1&at\sim 100~sunspots\end{cases}$ (59)
and at the same time have almost the same difference between the secondary reconnection time $(\tau_{rec})_{2}\leqslant 2.5~years$ of the horizontal field in the tachocline (which is equal to the sum of $(\tau_{n})_{max}\approx 2$ and the lifetime of sunspots ($\leqslant 0.5~years$) from several days to several months (see S. K. Solanki (2003); K. Petrovay & L. van Driel-Gesztelyi (1997)), which shift by $\sim 2.0~years$,
$(\tau_{n})_{max}=(\tau_{rec})_{max}^{secondary}-(\tau_{rec})_{max}^{primary}\sim 2.0~years$ (60)
between the “experimental” data of double sunspot maxima – very “slow” second maxima $(\tau_{rec})_{max}^{secondary}$ and very “fast” primary maxima $(\tau_{rec})_{max}^{primary}$ (see Fig. 12b).
This raises an unexpected but valid question: How does the unique source of the butterfly diagram is only in the form of 11-year fundamental sunspots or is only in the form of alternating 22-year fundamental solar activity cycles?
Here we show that, as expected, the source of the butterfly diagram is found only in the form of 11-year fundamental sunspots (see M. N. Gnevyshev (1967)). Below we show our solution with the so-called latitude-time interval (or gap) between the nearest butterflies of the remarkable solar physicist M. N. Gnevyshev (1967).
To understand the physics of the Gnevyshev gap, we need to briefly describe the solution of the following interrelated problems:
  • First, these are 11-year variations of sunspots (see V. D. Rusov et al. (2021)). A unique result of our model is the fact that the 11-year periods, velocities and modulations of the S102 star (see Fig. 6b in V. D. Rusov et al. (2021)) near the black hole are a significant indicator of the density modulation of the ADM halo in the fundamental plane of the Galactic center, which closely correlates with the density modulation of baryonic matter near the SMBH. If the ADM halo modulations in the GC black hole lead to ADM density modulations on the solar surface (via vertical density waves between the disk and the black hole to the solar neighborhood), then there is an “experimental” anticorrelation identity between such indicators as the 11-year ADM density modulation in the solar interior and the modulations of solar axions (or photons of axion origin), and equivalently between the modulations of solar axions (or photons of axion origin) and the sunspot cycles!
  • Secondly, it is an important holographic principle of quantum gravity. The most surprising fact is that the magnetic flux tube model at strong fields of the order of $10^{7}~G$ is a consequence of the thermomagnetic EN effect in the tachocline (see Fig. C.1a in V. D. Rusov et al. (2021)), which is the source of the so-called solar holographic mechanism of quantum gravity (see G. ’t Hooft (1993); L. Susskind (1995); J. Maldacena (2005); S. W. Hawking (2015b, a); S. W. Hawking et al. (2016). At the same time, we know that the hologram of the Sun, which is encoded in the magnetic field on its two-dimensional tachocline surface and is equivalent to the magnetic field (from the core to the tachocline) in the internal three-dimensional space of the Sun’s gravity, is in our case a remarkable source of the process with a continuous transformation of the toroidal field into a poloidal one (transformation $T\rightarrow P$), but not vice versa ($P\rightarrow T$) (see Fig. 2b and Eq. (2); also Fig. C.1a in V. D. Rusov et al. (2021)).
  • Thirdly, we know why we never see sunspots at the solar poles. The answer is quite simple. We know that both the tilt angle between the geometric and magnetic meridians, which are tilted by 7.25 degrees relative to the ecliptic, and the magnetic field of the northern (or southern) part of the meridional circulation, which radially corresponds to a length of about $\sim 22^{\circ}$, add up to about $\sim 30^{\circ}$. It follows clearly from this that the magnetic field of the meridional circulation part (see e.g. Fig.2b), will always slow down the magnetic field of the flux tube, which will almost never rise to the surface of the photosphere. This means that the magnetic fields of the flux tubes will only appear below about $90^{\circ}-30^{\circ}\approx 60^{\circ}$ at high latitudes. And that is precisely why no one observes sunspots at the poles of the Sun above $60^{\circ}$!
  • And fourthly, we can finally observe (see Fig. 6a of V. D. Rusov et al. (2021) for Japanese x-ray telescope Yohkoh (1991–2001)) and obtain the physics of axion origin photons only in those sunspots which are practically independent of the differential rotation of the Sun in the volume butterfly diagram. Clearly, in other sunspots which depend exactly on secondary reconnection of magnetic flux tubes and, as a consequence, on the tilt angle of Joy’s law on the surface of the Sun (see Fig. 10c), there can never be occurrences of axion origin photons in sunspots due to the differential rotation of the Sun!!!
From here we now return to the physics of the Gnevyshev gap. The answer is quite simple. The Gnevyshev intervals in the butterfly diagrams are related to the 11-year modulations of the ADM density, trapped on the Sun (via vertical density waves from the disk to the solar neighborhood) and, as a result, to anticorrelated indicators of variations in the sunspot cycles or photons of axion origin. When the ADM density in the solar interior is low (or the sunspot count is high), then the Gnevyshev intervals will be short in time. Conversely, when the ADM density in the solar interior is high (or the sunspot count is low), then the Gnevyshev intervals will be longer in time. As no surprisingly, it is the presence of ADM density modulations that shows that the appearance of “our indicated” Gnevychev intervals is the result of precisely 11-year variations of butterfly cycles with sunspots!
The most interesting thing is that it is the 11-year variations of sunspot cycles or photons of axion origin (see Figs. 5a and 6b in V. D. Rusov et al. (2021)) that are anticorrelated not only with the 11-year modulations of ADM density on the Sun, but also with the 11-year modulations of ADM density in the solar neighborhood, and, in particular, on Earth (see e.g. Fig. 13).
Figure 13: The time variations of total solar irradiance (TSI; orange curve; see Fig. 9 in (J. Pelt & O. Kärner, 2012)), simulated solar magnetic flux (black curve; see Fig. 7 in M. Dikpati et al. (2008); and also Fig. 16 in V. D. Rusov et al. (2013)) (both curves were smoothed with 29-year weighted moving average); the geomagnetic field secular variations (the Y-component; red curve) measured at the Eskdalemuir observatory (World Data Centre of Geomagnetism (2025)World Data Centre of Geomagnetism (Edinburgh), Edinburgh); variations in the Earth’s rotation speed (green curve; joined data from Fig. 3 in N. S. Sidorenkov (2016), Fig. 3.9 in N. S. Sidorenkov (2009), and the EOP C01 IAU1980 (1846-now) dataset by International Earth Rotation and Reference Systems Service (2025)); the variations of the average global ocean level (Pacific Decadal Oscillation (PDO), NOAA Physical Sciences Laboratory (2025) + Atlantic Multidecadal Oscillation (AMO), NOAA Physical Sciences Laboratory (2023); blue curve) with a trend in mm/years near 60-year oscillation (see Fig. 1 in D. P. Chambers et al. (2012)). The latter three curves were first smoothed using 5-year moving average, and then the 11-year moving average. All considered time variations between the Sun and the Earth always have a time interval of approximately $\sim 10\div 11$ years. The only serious disagreements between the reconstructions and observations occur during the Second World War, especially in the period 1944-1945 (see C. K. Folland et al. (2018)).
First, we recall that the physics of experimental data related to the anticorrelation between 11-year variations in the luminosity of the Sun and the Earth (see Fig. 13) currently cannot explain one of the most mystical phenomena of solar-terrestrial physics, which is contained in well-known experiments, for example, between the temporal variations of the magnetic flux of the solar tachocline (black curve; see Fig. 7 in M. Dikpati et al. (2008); and also Fig. 16 in V. D. Rusov et al. (2013)), and secular variations in the Earth’s magnetic field (Y-component from World Data Centre of Geomagnetism (2025)World Data Centre of Geomagnetism (Edinburgh) (Edinburgh)), or variations in the Earth’s rotation speed (see Fig. 3 in N. S. Sidorenkov (2016), Fig. 3.9 in N. S. Sidorenkov (2009), and EOP C01 IAU1980 (1846-now) dataset by International Earth Rotation and Reference Systems Service (2025)), as well as the variations in the mean global sea level (Pacific Decadal Oscillation (PDO) (NOAA Physical Sciences Laboratory, 2025) + Atlantic Multidecadal Oscillation (AMO) (NOAA Physical Sciences Laboratory, 2023)) with a trend in mm/years near 60-year oscillation (see Fig. 1 in D. P. Chambers et al. (2012)).
From here follows the main argument for solving the above problem (see Fig. 13; and also Fig. 6 in V. D. Rusov et al. (2021)). We understand that it is the presence of asymmetric dark matter (ADM) in the Earth’s core that can, on the one hand, be the reason for solving the density deficit in the inner core, and on the other hand, achieve the necessary consensus (see D. Antonangeli et al. (2018); E. Edmund et al. (2019)) not only regarding the absolute values of the solid iron velocities under the conditions of the inner core, but also regarding the iron melt velocities from pressure and temperature. If this is true, then we can prove that the inner core is not only a hidden reservoir of hydrogen on Earth (see e.g. T. Sakamaki et al. (2016), but also the cause of the existence of a heat flow source from the Earth’s core, which, based on ADM variability (see V. D. Rusov et al. (2021)), generates 11-year thermal, magnetic and temperature cycles in all parts of the Earth, including from the Earth’s surface to the troposphere.
This means that when the maximum of magnetic fluxes on the Earth’s surface exceeds the energy of solar irradiation in the troposphere, and the minimum of these fluxes on the Earth’s surface becomes equal to solar irradiation in the troposphere, then the 11-year luminosity of the Earth will anticorrelate with the 11-year luminosity of the Sun.
3 Summary and Outlook
The main result is the following: three “free” parameters – the magnetic field in the tachocline of the order of $\sim 10^{7}~G$ (see Fig. (A.1) and Eq. (A17) in V. D. Rusov et al. (2021)), axions with mass $m\sim 3.2\cdot 10^{-2}~eV$ and coupling constant $g_{a\gamma}\sim 4.4\cdot 10^{-11}~GeV^{-1}$ (see Eq. (11) in V. D. Rusov et al. (2021)) and the asymmetry dark matter (ADM) in the Universe with $m\sim 5~GeV$ (see V. D. Rusov et al. (2021); A. C. Vincent et al. (2016)) – are the source of a complete solution of the problem of an alternative theory of magnetic flux tubes in strong fields using 11-year variations of axion origin photons, which are caused by anticorrelations of 11-year variations of ADM density, gravitationally captured on the Sun.
From here, we understand that the fraction of dark matter in the Universe, which is about $\sim$27%, should be the reason for the appearance of a very large amount of dark matter with $m\sim 5~GeV$, for example, in our Milky Way galaxy. This means that the dark matter ADM in our galaxy should be not only in stars, but also on planets, for example, the Sun and the Earth.
As not surprisingly, a simple but fundamental conclusion emerges. The anticorrelation of the 11-year variations in the ADM density, gravitationally trapped both on the Sun and on the Earth, must necessarily participate in all the physical problems of dark matter, which are related, for example, not only to our solution of magnetic flux tubes in strong fields, describing the physics of sunspots, but also to the solution of the global climate of the Earth, describing the physics of global warming due to temperature and magnetic fields, but, crucially, without the existence of anthropogenic enhancement of the greenhouse effect.
The latter suggests that, in an intriguing perspective, the experimental observation of anticorrelations between 11-year variations in the luminosity of the Sun and the Earth (see Fig. 13) leads to the emergence of dark matter asymmetry in the Earth’s core, which, as not surprisingly, controls, among other things, the magnetic field from the core to the Earth’s surface. This means that when the maximum magnetic field at the Earth’s surface exceeds the solar irradiation energy in the troposphere, and the minimum of such energy flux at the Earth’s surface is equal to the solar irradiation energy in the troposphere, then the 11-year luminosity of the Earth will anticorrelate with the 11-year luminosity of the Sun.
This opens up a powerful prospect for observing both the axions, which in this mass range may be the main component of dark matter, and the anticorrelation of 11-year variations in the density of ADMs gravitationally trapped on both the Sun and Earth.

### v2

@@
 show]siiis@te.net.ua
-1 Introduction
+I Introduction
 It is known that the unsolved problem of energy transport by magnetic flux tubes at the same time represents another unsolved problem related to the sunspot darkness (see 2.2 in M.
@@
 D.
-Rusov et al., 2021), the toroidal magnetic field of tachocline by means of the thermomagnetic Ettingshausen-Nernst effect (Appendix A in V.
-D.
-Rusov et al.
-(2021)) “neutralizes” the magnetic field of the solar core $\sim 5\cdot 10^{7}~G$ (W.
-A.
-Fowler et al., 1955; S.
-Couvidat et al., 2003).
-It means that, using the thermomagnetic EN effect, a simple estimate of the magnetic pressure of an ideal gas in the tachocline of e.g.
-the Sun,
+Rusov et al., 2021), the toroidal magnetic field of tachocline by means of the thermomagnetic Ettingshausen-Nernst effect
 $\frac{B_{tacho}^{2}}{8\pi}=p_{ext}\approx 6.5\cdot 10^{13}\frac{erg}{cm^{3}}~~at~~0.7R_{Sun},$ (1)
-can indirectly prove that by using the holographic principle of quantum gravity (see Apendix C in V.
-D.
-Rusov et al.
-(2021)), the repelling toroidal magnetic field of the tachocline exactly “neutralizes” the magnetic field in the Sun’s core (see Fig.
-2, and also Fig.
-A1 in V.
-D.
-Rusov et al.
-(2021))
-$B_{tacho}^{Sun}=4.1\cdot 10^{7}~G=-B_{core}^{Sun},$ (2)
-where the projections of the magnetic fields of the tachocline and the core have equal values but opposite directions.
-Surprisingly, our experimental data provide observational evidence in support of a two-dimensional holographic explanation of the Sun and a powerful key to understanding the physics of quantum gravity in the Universe!
-This is explained once again by the fact that the experimental data, in which the toroidal magnetic field of the tachocline exactly ”neutralizes” the magnetic field in the solar core, should also theoretically describe both a two-dimensional holographic explanation of the solar tachocline, in which it is the magnetic field that must be solved, either in conformal field theory (see e.g.
-analogs in (R.
-Boyack et al., 2024; T.
-Kimura & S.
-Ozaki, 2019; A.
-B.
-Zamolodchikov, 1989; C.
-Callan et al., 1995; L.
-Hu et al., 2024; C.
-Cartwright et al., 2022; H.
-Nastase & K.
-Skenderis, 2020; V.
-A.
-Miransky & I.
-A.
-Shovkovy, 2015)) or at least in quantum mechanics (see e.g.
-analogs in (U.
-Gürsoy, 2021; S.
-Heefer, 2024)), and a three-dimensional magnetic field via gravity, which, in contrast to pseudo-Riemannian geometry, applies the extended (beyond Einstein geometry) Finsler geometry (see (U.
-Gürsoy, 2021; S.
-Heefer, 2024)), for example, inside the solar tachocline.
-In other words, the “encoded” two-dimensional information about the magnetic field of the tachocline in conformal field theory corresponds to three-dimensional information about the magnetic field in Finsler gravity, which is reduced to the curvature of space-time in a three-dimensional volume with the radius of the tachocline of the Sun.
-Such strong magnetic fields in a tachocline of e.g.
-the Earth, Sun, magnetic white dwarfs, accreting neutron stars and black holes can predict the exact “neutralization” of the magnetic field in the core of these stars and in a black hole (see Appendix C in V.
-D.
-Rusov et al.
-(2021)).
-From here we understand that when space-time curvature appears, gravitational waves arise, which ultimately uncover, as not surprisingly, the more powerful Finslerian nature of quantum gravity.
-So Finslerian quantum gravity can be applied in our case, as we know, to holograms of the Universe, black holes, stars and planets, for example, the Sun.
-Based on the “experimental example” of the existence and understanding of the physics of the holographic principle of quantum gravity, we apply the famous result of Stephen Hawking, who at a conference in Stockholm on August 28, 2015 announced that he had solved the information paradox.
-According to Hawking (S.
-W.
-Hawking, 2015a, b; S.
-W.
-Hawking et al., 2016), the AdS/CFT correspondence (see (J.
-Maldacena, 1998)) showed no loss of information, where all information about matter and energy in the three-dimensional volume of the black hole is encoded in the form of a hologram on its two-dimensional surface of the event horizon (see (G.
-’t Hooft, 1993; L.
-Susskind, 1995; J.
-Maldacena, 2005; M.
-Hanada et al., 2014)).
-It resembles an analogue of the famous photohologram (see (D.
-Gabor, 1948)), which is a certain type of photograph that produces a three-dimensional image under good lighting; all the information describing a three-dimensional scene is encoded in a pattern of light and dark areas applied to a two-dimensional film.
-This means that the possible description and solution of the ”encoded” holographic two-dimensional ”film” is predetermined by the fact that the problem of the information paradox, known as the apparent contradiction between general relativity and quantum mechanics, completely disappears.
-This approach ultimately leads to a holographic principle of quantum gravity, which can be generalized to any physical system occupying space-time.
-For example, in the holograms of the Universe (N.
-Afshordi et al., 2017), black holes, white dwarfs, neutron stars or, surprisingly, in the hologram of the Sun’s tachocline!
-Let us note one more important mechanism of quantum gravity (see Fig.
-2b).
-The most surprising fact is that the antidynamo model of a magnetic flux tube at strong fields of the order of $10^{7}~G$ is a consequence of both the thermomagnetic EN effect and the so-called solar holographic antidynamo, in which the poloidal field arises directly from the toroidal field, but not vice versa (see Fig.
-2b; and also Fig.
-C.1a in V.
-D.
-Rusov et al.
-(2021)).
-Using the quantum gravity mechanism, and hence, the strong MFT field, we are interested in the existence of dark matter axions identical to solar axions, which strongly affects the the conversion of axion origin photons in magnetic O-loops near the tachocline, and are thus related to the so-called thermomagnetic Ettingshausen-Nernst effect through the Parker-Biermann cooling effect (see Apendix A in V.
+can indirectly prove that toroidal magnetic field of the tachocline is exactly equal to
+$B_{tacho}^{Sun}=4.1\cdot 10^{7}~G,$ (2)
+Using the strong MFT field, we are interested in the existence of dark matter axions identical to solar axions, which strongly affects the the conversion of axion origin photons in magnetic O-loops near the tachocline, and are thus related to the so-called thermomagnetic Ettingshausen-Nernst effect through the Parker-Biermann cooling effect (see Apendix A in V.
 D.
@@
 (b) 11-year variations of sunspot magnetic fields as a component of the solar anti-dynamo.
-In contrast to the component of the solar dynamo model (a) [^2] If the magnetic field at these levels is maintained by a dynamo, and the magnetic fields are diffuse, then the field strength is unlikely to be more than $\sim 10^{5}~G$, because larger fields would severely inhibit dynamo action (see E.
-E.
-Deluca & P.
-A.
-Gilman (1986))., the mechanism of the fundamental holographic principle of quantum gravity (see H.
-Nastase & K.
-Skenderis (2020); N.
-Afshordi et al.
-(2017)), in which the hologram of the Sun is encoded in the magnetic field on its two-dimensional surface tachocline and is equivalent to the magnetic field (from the core to the tachocline) in the internal three-dimensional space of the Sun’s gravity, and consequently, the formation of the thermomagnetic EN effect (see L.
+The formation of the thermomagnetic EN effect (see L.
 J.
@@
 S-102 with a period of about 11 years, lead (via vertical density waves from the black hole to the solar neighborhood) to modulations of the ADM density, gravitationally trapped in the interior of the Sun.
-This means that anticorrelation modulations of the ADM density control the solar luminosity or sunspot cycles, in which the sunspot variations are identical to the 11-year variations of the magnetic flux tubes (see Section 2): maximum (red lines) and minimum (green-red lines).
+This means that anticorrelation modulations of the ADM density control the solar luminosity or sunspot cycles, in which the sunspot variations are identical to the 11-year variations of the magnetic flux tubes (see Section II): maximum (red lines) and minimum (green-red lines).
 In order to answer this question, let us first consider all the unexpected and intriguing implications of the 11-year modulations of the ADM density in the solar interior and around the BH (see Sect.
@@
 Taking into account the above, we structure our paper as follows.
-In Section 2, we discuss the physics of nearly empty flux tubes and the connection with dark matter axions out of the solar core.
-In Section 2.1, we present convective heating and magnetic buoyancy of flux tubes by means of axion origin photons.
-In Section 2.2, we discuss the physics of primary and secondary reconnection of flux tubes in the lower layers and the features of the Joy’s law tilt angle.
-In Section 2.2.2, we discuss the physics of the observed double maxima of sunspot cycles, where the first maximum does not depend on the tilt angle of the Joy’s law, while the second one definitely depends on the tilt angle of the Joy’s law.
+In Section II, we discuss the physics of nearly empty flux tubes and the connection with dark matter axions out of the solar core.
+In Section II.1, we present convective heating and magnetic buoyancy of flux tubes by means of axion origin photons.
+In Section II.2, we discuss the physics of primary and secondary reconnection of flux tubes in the lower layers and the features of the Joy’s law tilt angle.
+In Section II.2.2, we discuss the physics of the observed double maxima of sunspot cycles, where the first maximum does not depend on the tilt angle of the Joy’s law, while the second one definitely depends on the tilt angle of the Joy’s law.
 We also showed the relationship between the ADM density in the solar interior (or, for example, the number of anticorrelated sunspots) and the Gnevyshev gap.
-Finally, in Section 3, we provide a summary and outlook for this paper.
-2 Physics of the theory of practically empty magnetic flux tubes at strong fields, $\sim 10^{7}~G$, at which there is a free path for photons of axion origin from the tachocline to the photosphere.
+Finally, in Section III, we provide a summary and outlook for this paper.
+II Physics of the theory of practically empty magnetic flux tubes at strong fields, $\sim 10^{7}~G$, at which there is a free path for photons of axion origin from the tachocline to the photosphere.
 This raises the fundamental question of the self-exciting action of the dynamo, the remarkable physics of which was discovered some 70 years ago (see E.
@@
 Karak (2024)).
-Conversely, if the question arises whether there is a satisfactory alternative theory against the dynamo action, our short answer is ”yes”.
-We understand that the strong toroidal magnetic field in the tachocline is generated not by the dynamo action, but by the holographic mechanism of the tachocline (see Fig.
-2b), which is a consequence of the holographic principle of quantum gravity in the Universe, in the black hole of our galaxy, and therefore on the Sun.
-In order for our alternative theory (for example, magnetic flux tubes in strong fields using photons of axion origin) to solve all its problems (see Sections 2.1-2.2), it is necessary to solve the most important problem: In what way come sunspots, generated by a strong magnetic field in a flux tube from the base of the convection zone to the surface of the Sun, have 11-year variations without any dynamo?
+In order for our alternative theory (for example, magnetic flux tubes in strong fields using photons of axion origin) to solve all its problems (see Sections II.1-II.2), it is necessary to solve the most important problem: In what way come sunspots, generated by a strong magnetic field in a flux tube from the base of the convection zone to the surface of the Sun, have 11-year variations without any dynamo?
 The answer is very simple!
-Almost all physicists, and in particular dynamo physicists, know that in the Universe, and, of course, in our galaxy, there are dark matters, and the highest density of dark matter is located exactly at the center of the Galaxy, i.e.
+Almost all physicists, and in particular dynamo physicists, know that in the Universe, and, of course, in our galaxy, there are ADM, and the highest density of ADM is located exactly at the center of the Galaxy, i.e.
 near the black hole.
@@
 From here, we understand that when the period of S-102 is very close to the black hole, then the high speed of this star is such that the dark matter is practically not captured by this star.
-But when the elliptical star S-102 rotates farther from the black hole, then the star gravitationally captures these dark matters.
+But when the elliptical star S-102 rotates farther from the black hole, then the star gravitationally captures these ADMs.
 This means that the 11-year variations in the density of dark matter via vertical density waves from the black hole disk to the solar neighborhood V.
@@
 sunspots are never associated with any dynamo action.
-Another intriguing question arises here.
-Unlike our galaxy, such S-type stars (between the black hole and the disk) may simply not exist in many galaxies, and consequently, there may not be any solar dynamo at all – for example, there is no Sun-like dynamo on the active star $\zeta$ Andromedae, which is part of the Zeta Andromedae star system in the constellation Andromeda (R.
-M.
-Roettenbacher et al., 2016).
-(It is located approximately 189 light years from Earth.) When various stars and clusters (e.g.
-Galactic Globular Clusters) have strong toroidal magnetic fields of $\geqslant 10^{7}~G$ in the corresponding tachoclines, then, unlike the self-excited dynamo, the manifestation of the tachocline itself in various stars and clusters is a result of the funda-mental holographic principle of quantum gravity (see J.
-M.
-Maldacena (1999); J.
-Maldacena (2005); S.
-W.
-Hawking (2015a, b); S.
-W.
-Hawking et al.
-(2016); and also N.
-Afshordi et al.
-(2017) and Fig.A.1 and Fig.C.1 in V.
-D.
-Rusov et al.
-(2021)).
-This means that the self-excited dynamo in the tachocline may not exist at all!
 We could raise a few more questions, but it is enough for us that “…not enough for a scientific understanding of the solar dynamo.
@@
 the maximum magnetic pressure in the magnetic tubes.
-The narrow “purple” rings between the O-loop and the tube walls ($\rho_{ext}=\rho_{int}^{0}\gg\rho_{int}$) with the Parker-Bierman cooling effect inside, are a very important result of the existence of convective heating $(dQ/dt)_{2}$ in Section 2.1.
+The narrow “purple” rings between the O-loop and the tube walls ($\rho_{ext}=\rho_{int}^{0}\gg\rho_{int}$) with the Parker-Bierman cooling effect inside, are a very important result of the existence of convective heating $(dQ/dt)_{2}$ in Section II.1.
 On the other hand, high-energy photon fluxes coming from the radiation zone through the thin “ring” part of the magnetic tube, i.e.
@@
 3 and 4), and which is applied without the Parker-Biermann cooling effect, allows to determine the convective heating $(dQ/dt)_{2}$ (see Sect.
-2.1).
+II.1).
 Figure 4: Kolmogorov turbulent cascade (A.
@@
 One of the final goals of this section is to determine the general regularities of the theory of magnetic flux tubes in strong fields of the order of $\sim 10^{7}~G$ using photons of axion origin – the so-called universal Ballegooijen-Fan-Fisher model (Sec.
-2.1), which are generated by the magnetic buoyancy of practically empty tubes, and as a consequence, their rise from the tachocline to the solar surface (Figs.
+II.1), which are generated by the magnetic buoyancy of practically empty tubes, and as a consequence, their rise from the tachocline to the solar surface (Figs.
 3-5).
 Another goal is the essence of the physics of the primary and secondary reconnection of magnetic flux tubes in the lower layers of the tachocline (Sect.
-2.2), which is directly related to the observed features of the cycle of both double sunspot maxima (see Sect.
-2.2.2) and the tilt angle of Joy’s law (see Sect.
-2.2.1).
+II.2), which is directly related to the observed features of the cycle of both double sunspot maxima (see Sect.
+II.2.2) and the tilt angle of Joy’s law (see Sect.
+II.2.1).
 Moreover, both effects (see Sect.
-2.1-2.2) are caused by the existence (of dark matter?) – solar axions generated in the core of the Sun.
-2.1 Radiative $(dQ/dt)_{1}$ sec-reconnectionand convective $(dQ/dt)_{2}$ heatings and the buoyant rise of magnetic flux tubes and the phenomenon of solar axions
+II.1-II.2) are caused by the existence (of dark matter?) – solar axions generated in the core of the Sun.
+II.1 Radiative $(dQ/dt)_{1}$ sec-reconnectionand convective $(dQ/dt)_{2}$ heatings and the buoyant rise of magnetic flux tubes and the phenomenon of solar axions
 The first problem is devoted to the study of the effect of virtually empty magnetic tubes and the phenomenon of solar axions.
@@
 $\tau_{d}\approx\frac{2}{\beta}T_{ext}\left[\frac{1}{c_{p}\rho_{e}}\left(\frac{dQ}{dt}\right)_{1}\right]^{-1},$ (8)
-at which we have a very important (see section 2.2.1) value of
+at which we have a very important (see section II.2.1) value of
 $\tau_{d}\approx\left.5\frac{1}{\beta}p_{ext}\middle/\left(\frac{dQ}{dt}\right)_{1}\right.\approx 0.84\cdot 10^{8}~sec\approx 2.5~years,$ (9)
@@
 $v_{rise}=H_{p}\nabla_{ad}\frac{1}{p_{ext}}\left(\frac{dQ}{dt}\right)_{1}\left(|\delta|+0.12\frac{B^{2}}{8\pi p_{ext}}\right)^{-1},~~~\nabla_{ad}=\nabla_{e}\simeq 0.4,$ (10)
-at which we have a very important (see Sections 2.1-2.2) value of
+at which we have a very important (see Sections II.1-II.2) value of
 $v_{rise}\approx 1.7\cdot 10^{2}~cm/s,$ (11)
@@
 Fisher, 1996)).
-Using simple calculations of equations (3) and (2.1) for MFTs, it is easy to show that with the help of the
+Using simple calculations of equations (3) and (II.1) for MFTs, it is easy to show that with the help of the
 $\frac{dQ}{dt}=\left(\frac{dQ}{dt}\right)_{1}+\left(\frac{dQ}{dt}\right)_{2}$ (14)
@@
 $\frac{B_{c}^{2}}{8\pi p_{ext}}=-\gamma\delta=-\frac{5}{3}\delta.$ (18)
-Ultimately, we call these Eqs. (2.1)-(18), in honor of these remarkable “solar” physicists, the universal van Ballegooijen-Fan-Fisher model (vanBFF model).
+Ultimately, we call these Eqs. (II.1)-(18), in honor of these remarkable “solar” physicists, the universal van Ballegooijen-Fan-Fisher model (vanBFF model).
 On the other hand, let us remind that on the basis of the holographic mechanism, generating the toroidal magnetic field in the tachocline, the universal model of flux tubes is predetermined by the existence of strong magnetic fields of the order of $B_{tacho}\sim 10^{7}~G$.
@@
 where we used the MFT heating rate which consists of the radiative heating rate $(dQ/dt)_{1}$ and convective heating rate $(dQ/dt)_{2}$ (see Eqs.
-(2.1)–(19), Eqs.
+(II.1)–(19), Eqs.
 (13) and (19) in (Y.
@@
 At the same time, we remember that anticorrelated 11-year ADM density modulations in the Sun interior (see Fig.
-2b) and the ADM around BH (see Sect.
+1b) and the ADM around BH (see Sect.
 3 in (V.
@@
 In other words, this means that our (maximum (31) and (36)) and minimum ((34) and (37)) are the remarkable existence of a maximum and minimum of 11-year variations of the photosphere magnetic flux tube for sunspots, which, unlike the solar dynamo, are predetermined as flux tube magnetic field sources (anchored in the overshoot tachocline), and are coupled by the holographic mechanism of quantum gravity (see Fig.
-2b), and anticorrelated 11-year ADM density variations gravitationally trapped in the interior of the Sun (see Section 3 in (V.
+1b), and anticorrelated 11-year ADM density variations gravitationally trapped in the interior of the Sun (see Section 3 in (V.
 D.
@@
 W.
-Hawking (2015b, a); S.
+Hawking (2015a, b); S.
 W.
@@
 Finally, let us show that if the rise velocity of the buoyant magnetic tube, determined by the magnetic field values maximum ($B_{max}\sim 3.6\cdot 10^{7}~G$) and minimum ($B_{max}\sim 3.6\cdot 10^{6}~G$) of the magnetic field and the width of the “ring” $(a_{conv})_{max}\sim 3.7\cdot 10^{-4}~H_{p}$ or $(a_{conv})_{min}\sim 3.7\cdot 10^{-3}~H_{p}$, causes the MFT to appear on the surface in the form of sunspots, then the parameters of the universal vanBFF model will be associated with the magnetic cycles that are directly related to the observed features of the cycle of both double sunspot maxima (see Sect.
-2.2.2) and the tilt angle of Joy’s law (see Sect.
-2.2.1).
-2.2 Physics of primary and secondary reconnection of magnetic flux tubes in the lower layers of the tachocline
+II.2.2) and the tilt angle of Joy’s law (see Sect.
+II.2.1).
+II.2 Physics of primary and secondary reconnection of magnetic flux tubes in the lower layers of the tachocline
 The problem is devoted to the physics of primary and secondary magnetic reconnection in the lower layers of the tube, which is directly related to the observed features of the cycle of both double maxima of sunspots and the tilt angle of Joy’s law.
@@
 10c), and during the secondary reconnection for the time $(\tau_{n})_{rec}$ of the departure of the same rise of the almost empty magnetic tube from the surface of the Sun and, as a consequence, their “disappearance” on the Sun.
-2.2.1 On the ascending magnetic tube and the formation of sunspots both near the tacho-cline with $\sim 10^{7}~G$ in the absence of Joy’s law and in the middle region with $\sim 10^{5}~G$ in the presence of Joy’s law
+II.2.1 On the ascending magnetic tube and the formation of sunspots both near the tacho-cline with $\sim 10^{7}~G$ in the absence of Joy’s law and in the middle region with $\sim 10^{5}~G$ in the presence of Joy’s law
 The first part of the scenario consists in discussing the physics of magnetic flux tube buoyancy at strong fields and with the participation of axion origin photons (see Fig.
@@
 Figure 9: Scheme of turbulent reconstruction of the toroidal magnetic field in the convective zone (CZ): (a) meridional circulation (closed blue poloidal field lines), which is the source of the toroidal field in the tachocline (thin black lines) based on the holographic mechanism of quantum gravity (see Fig.
-2b), and the magnetic buoyancy of the flux tubes (red arrows) with a magnetic field of $\sim 10^{7}~G$; (b) In the CZ, through secondary turbulent reconnection of the magnetic tube near the tachocline (see Fig.
+1b), and the magnetic buoyancy of the flux tubes (red arrows) with a magnetic field of $\sim 10^{7}~G$; (b) In the CZ, through secondary turbulent reconnection of the magnetic tube near the tachocline (see Fig.
 10b2,c), effects of both macroscopic turbulent plasma diamagnetism are observed, facilitating the formation of two layers on the surface (see analogous (V.
@@
 9b and Fig.
-10c), which are the source of the double maxima of the 11-year solar cycles (see the full experimental evidence in Section 2.2.2), we are interested in the problem or, more precisely, the physics of the tilt angle of Joy’s law (V.
+10c), which are the source of the double maxima of the 11-year solar cycles (see the full experimental evidence in Section II.2.2), we are interested in the problem or, more precisely, the physics of the tilt angle of Joy’s law (V.
 G.
@@
 Conversely, when the flux tube has a small radius and a high magnetic field, then the magnetic tube, on the one hand, is not connected to the source of the formation of the tilt angle, and on the other hand, surprisingly, has a small spot area on the surface of the Sun.
-2.2.2 New physics of observed double maxima of sunspot cycles and solutions of time interval or Gnevyshev gap
+II.2.2 New physics of observed double maxima of sunspot cycles and solutions of time interval or Gnevyshev gap
 The phenomenon of a double peak in the sunspot cycle was discovered by M.
@@
 If the ADM halo modulations in the GC black hole lead to ADM density modulations on the solar surface (via vertical density waves between the disk and the black hole to the solar neighborhood), then there is an “experimental” anticorrelation identity between such indicators as the 11-year ADM density modulation in the solar interior and the modulations of solar axions (or photons of axion origin), and equivalently between the modulations of solar axions (or photons of axion origin) and the sunspot cycles!
-  • Secondly, it is an important holographic principle of quantum gravity.
-The most surprising fact is that the magnetic flux tube model at strong fields of the order of $10^{7}~G$ is a consequence of the thermomagnetic EN effect in the tachocline (see Fig.
+  • Secondly, the most surprising fact is that the magnetic flux tube model at strong fields of the order of $10^{7}~G$ is a consequence of the thermomagnetic EN effect in the tachocline (see Fig.
 C.1a in V.
@@
 Rusov et al.
-(2021)), which is the source of the so-called solar holographic mechanism of quantum gravity (see G.
-’t Hooft (1993); L.
-Susskind (1995); J.
-Maldacena (2005); S.
-W.
-Hawking (2015b, a); S.
-W.
-Hawking et al.
-(2016).
-At the same time, we know that the hologram of the Sun, which is encoded in the magnetic field on its two-dimensional tachocline surface and is equivalent to the magnetic field (from the core to the tachocline) in the internal three-dimensional space of the Sun’s gravity, is in our case a remarkable source of the process with a continuous transformation of the toroidal field into a poloidal one (transformation $T\rightarrow P$), but not vice versa ($P\rightarrow T$) (see Fig.
-2b and Eq.
+(2021)), which is in our case a remarkable source of the process with a continuous transformation of the toroidal field into a poloidal one (transformation $T\rightarrow P$), but not vice versa ($P\rightarrow T$) (see Fig.
+1b and Eq.
 (2); also Fig.
@@
 This means that when the maximum of magnetic fluxes on the Earth’s surface exceeds the energy of solar irradiation in the troposphere, and the minimum of these fluxes on the Earth’s surface becomes equal to solar irradiation in the troposphere, then the 11-year luminosity of the Earth will anticorrelate with the 11-year luminosity of the Sun.
-3 Summary and Outlook
+III Summary and Outlook
 The main result is the following: three “free” parameters – the magnetic field in the tachocline of the order of $\sim 10^{7}~G$ (see Fig.

### v3

@@
-An alternative theory of magnetic flux tubes in strong fields via axion origin photons
+A phenomenological model of magnetic flux tubes in strong fields induced by axion-origin photons
 Abstract
-In our alternative theory, built around the coincidence of experimental and theoretical data, three ”free” parameters – the magnetic field in the tachocline of the order of $\sim 10^{7}~G$ (see Fig.
-(A.1) and Eq.
-(A17) in V.
-D.
-Rusov et al.
-(2021)), the axion mass $m_{a}\sim 3.2\cdot 10^{-2}~eV$ (see Eq.
-(11) in V.
-D.
-Rusov et al.
-(2021)), and the asymmetric dark matter (ADM) in the Universe with $m_{ADM}\sim 5~GeV$ ((see V.
-D.
-Rusov et al.
-(2021); A.
-C.
-Vincent et al.
-(2016)) – give a complete solution to the problem of the theory of magnetic flux tubes in strong fields with 11-year variations of axion-origin photons, which are caused by and anticorrelated to the 11-year variations in density of ADM, gravitationally captured on the Sun.
-[…]
-show]siiis@te.net.ua
+We propose a phenomenological model in which empty flux tubes in strong magnetic fields are associated with both the conversion of axions to photons of axion origin to generate coronal heating, and the simultaneous preservation of the Parker-Biermann cooling effect by converting high-energy photons from the radiative zone to axions from the convective zone through the O-loop tachocline, is the source of the rise of the magnetic flux tube from the tachocline to the photosphere in the form of sunspots.
+The model is characterized by two free parameters, justified by existing theoretical and observational constraints, the axion mass $m_{a}\sim 3.2\cdot 10^{-2}~eV$, and asymmetric dark matter (ADM) with a particle mass $m_{ADM}\sim 5~GeV$ gravitationally captured by the Sun.
+In this scenario, temporal modulations of the axion density are caused by anticorrelated 11-year modulations of ADM gravitationally captured by the Sun.
+We also obtained theoretical estimates of the velocities and times in empty magnetic tubes in strong fields rising from the tachocline to the surface of the photosphere in the form of sunspots.
+This ensured complete agreement with known experimentally measured velocities and times using local helioseismological time-distance inversions along the ray trajectories of acoustic waves.
+This allows us to study not only the subsurface structure and dynamics of active regions, but also the depths of the empty magnetic flux tube, extending up to 200,000 km from the tachocline to the photosphere.
+The proposed model links the phenomenology of axions with the magnetic structures of the Sun and provides testable consequences for astrophysical and solar observations.
 I Introduction
-It is known that the unsolved problem of energy transport by magnetic flux tubes at the same time represents another unsolved problem related to the sunspot darkness (see 2.2 in M.
-Rempel & R.
-Schlichenmaier (2011)).
+It is well known that helioseismological methods are well-suited for studying the Sun’s rotation and structure from helioseismic reversals, but they have difficulty distinguishing the effects of very strong magnetic fields from other structural changes, and therefore cannot definitively confirm the presence of a magnetic field in the solar tachocline exceeding $10^{5}~G$.
+In other words, it is difficult to distinguish the effects of very strong magnetic fields in the tachocline, but this does not mean they should be discarded due to their virtually undetectable nature using direct seismic methods.
+This raises the question of why the dynamo action is only associated with a magnetic field of about $10^{5}~G$, but is definitely not associated with a stronger magnetic field of about $10^{7}~G$?
+It is known that “…even if complete knowledge of the solar dynamo existed – and it does not…”, as has been pointed out throughout the review [1] – a number of crucial questions require answers, including, at least: How do sunspots with structure, rotation velocity, and 11-year luminosity form and disappear?
+Why has spot formation not been captured even in local models, let alone global models?
+How do turbulence coefficients – turbulent pumping, turbulent diffusion – change with rotation velocity, luminosity, and internal structure?
+Why the inability to model accurately enough the surface layers where the density drops vertiginously is, according to [2], the common characteristic of all existing dynamo models?
+At that, using the so-called new era of 3D modeling and observations [1], they (solar physicists), surprisingly, believe that the study of the solar dynamo will be definitely solved within the next 10-20 years.
+But the most important thing is that any dynamo action that involves, for example, the 11-year oscillations of sunspots or a sunspot butterfly diagram, must contain a small number of “free parameters” – usually three or less (see Dahlquist’s key theorem [3]) – because of the mathematical structure of the problem, where the independence of the equations, the stability of the system, and the dimensionality restrictions reduce the number of “degrees of freedom” needed to determine the solution.
+We could raise a few more questions, but it is enough for us that “…not enough for a scientific understanding of the solar dynamo.
+For there is an unfortunate revelation: with four free parameters you can provide a satisfactory fit to the horizon of New York or Beijing” – as was said by the great Eugene N.
+Parker, who in 2009 in his last article [4] first began to strongly doubt the existence of the solar dynamo.
+Now, in contrast to the no more than of three “free” parameters of the solar dynamo, consider our next two “free” parameters: the dark matter axions in the Sun’s core and the dark matter asymmetry (ADM) gravitationally captured by the Sun.
+These variables in the corresponding models cannot be predicted by the theory itself and must be determined from experimental data, as they are related to fundamental physical constants.
+Let us first consider the first “free parameter” – the dark matter axions in the Sun’s core, which are necessary to solve the well-known problem of energy transfer by magnetic flux tubes from the tachocline to the Sun’s photosphere.
+It is known that the unsolved problem of energy transport by magnetic flux tubes at the same time represents another unsolved problem related to the sunspot darkness (see 2.2 in [5]).
 Of all the known concepts playing a noticeable role in understanding the connection between the energy transport and sunspot darkness, let us consider the most significant theory, in our view.
-It is based on the Parker-Biermann cooling effect (E.
-N.
-Parker, 1955a; L.
-Biermann, 1941; E.
-N.
-Parker, 1979a) and originates from the early works of L.
-Biermann (1941) and H.
-Alfvén (1942).
-As you know, the Parker-Biermann cooling effect (E.
-N.
-Parker, 1955a; L.
-Biermann, 1941; E.
-N.
-Parker, 1979a), which plays a role in our current understanding, originates from L.
-Biermann (1941) and H.
-Alfvén (1942): in a highly ionized plasma, the electrical conductivity can be so large that the magnetic fields are frozen into the plasma.
+It is based on the Parker-Biermann cooling effect [6, 7, 8] and originates from the early works of Biermann [7] and Alfvén [9].
+As you know, the Parker-Biermann cooling effect [6, 7, 8], which plays a role in our current understanding, originates from Biermann [7] and Alfvén [9]: in a highly ionized plasma, the electrical conductivity can be so large that the magnetic fields are frozen into the plasma.
 Biermann realized that the magnetic field in the spots themselves can be the cause of their coolness – it is colder because the magnetic field suppresses the convective heat transport.
 Hence, the darkness of the spot is due to a decrease in surface brightness.
-Parker (E.
-N.
-Parker, 1955a, 1974a, 1974b, 1974, 1979a) has pointed out that the magnetic field can be compressed to the enormous intensity only by reducing the gas pressure within the flux tube relative to the pressure outside, so that the external pressure compresses the field.
+Parker [6, 10, 11, 12, 8] pointed out that the magnetic field can be compressed to the enormous intensity only by reducing the gas pressure within the flux tube relative to the pressure outside, so that the external pressure compresses the field.
 The only known mechanism for reducing the internal pressure sufficiently is a reduction of the internal temperature over several scale heights so that the gravitational field of the Sun pulls the gas down out of the tube (as described by the known barometric law $dp/dz=-\rho g$).
-Hence it appears that the intense magnetic field of the sunspot is a direct consequence of the observed reduced temperature (E.
-N.
-Parker, 1955a).
-On the other hand, Parker (E.
-N.
-Parker, 1974, 1977) has also pointed out that the magnetic inhibition of convective heat transport beneath the sunspot, with the associated heat accumulation below, raises the temperature in the lower part of the field.
+Hence it appears that the intense magnetic field of the sunspot is a direct consequence of the observed reduced temperature [6].
+On the other hand, Parker [12, 13] has also pointed out that the magnetic inhibition of convective heat transport beneath the sunspot, with the associated heat accumulation below, raises the temperature in the lower part of the field.
 The barometric equilibrium leads to enhanced gas pressure upward along the magnetic field, causing the field to disperse rather than intensify.
-Consequently, E.
-N.
-Parker (1974) argued that the temperature of the gas must be influenced by something more than the inhibition of heat transport!
-Our unique alternative idea is that the explanation of sunspots is based not only on the suppression of convective heat transfer by a strong magnetic field (of the order of $\sim 10^{7}~G$ (V.
-D.
-Rusov et al., 2021)) through the enhanced cooling of the Parker-Biermann effect (E.
-N.
-Parker, 1974a), but also on the appearance of the axions of photonic origin (Fig 3, Fig.
-B.1 in V.
-D.
-Rusov et al.
-(2021)) from the tachocline to the photosphere, which is confirmed by the “disappearance” of the heat, and consequently, the temperature in the lower part of the magnetic tube (E.
-N.
-Parker, 1974, 1977) due to the axions of photonic origin from the photon-axion oscillations in the O-loop near the tachocline (see Fig.
-3).
-This means that the appearance of axions of photonic origin, which remove the problem of the temperature rise in the lower part of the magnetic tube, and the photons of axionic origin, which have the free path (Rosseland length; see Fig.
-B.3 in (V.
-D.
-Rusov et al., 2021)) from the tachocline to the photosphere, are the explanation of sunspots based not only on the suppression of convective heat transport by a strong magnetic field (V.
-D.
-Rusov et al., 2021), but also on the indispensable existence of the Parker-Biermann cooling effect.
-At the same time, we clearly understand that stronger fields would seriously suppress the action of the dynamo (see (E.
-E.
-Deluca & P.
-A.
-Gilman, 1986)).
-On the other hand, we understand that the existence of the Parker-Biermann cooling effect is associated with the so-called thermomagnetic Ettingshausen-Nernst effect (see Apendix A in V.
-D.
-Rusov et al.
-(2021)).
-Due to the large temperature gradient in the tachocline, the thermomagnetic EN effect (A.
-V.
-Ettingshausen & W.
-Nernst, 1886; E.
-H.
-Sondheimer, 1948; L.
-J.
-Spitzer, 1956; Y.
-B.
-Kim & M.
-J.
-Stephen, 1969) creates electric currents that are inversely proportional to the strong magnetic field of the tachocline.
-As we showed earlier (V.
-D.
-Rusov et al., 2021), the toroidal magnetic field of tachocline by means of the thermomagnetic Ettingshausen-Nernst effect
+Consequently, [12] argued that the temperature of the gas must be influenced by something more than the inhibition of heat transport!
+Our idea is that the explanation of sunspots is based not only on the suppression of convective heat transfer by a strong magnetic field (of the order of $\sim 10^{7}~G$ [14]) through the enhanced cooling of the Parker-Biermann effect [10], but also on the appearance of the axions of photonic origin (Fig 2, Fig.
+B.1 in [14]) from the tachocline to the photosphere, which is confirmed by the “disappearance” of the heat, and consequently, the temperature in the lower part of the magnetic tube [12, 13] due to the axions of photonic origin from the photon-axion oscillations in the O-loop near the tachocline (see Fig.
+2).
+This means that the appearance of axions of photonic origin, with which the problem of temperature increase in the lower part of a strong magnetic tube disappears, and the photons of axionic origin, which have the free path (Rosseland length; see Fig.
+B.3 in [14]) from the tachocline to the photosphere, are the explanation of sunspots based not only on the suppression of convective heat transport by a strong magnetic field [14], but also on the indispensable existence of the Parker-Biermann cooling effect.
+At the same time, we clearly understand that stronger fields (e.g.
+of the order of $10^{7}~G$) would seriously suppress the action of the dynamo (see [15]).
+On the other hand, we understand that the existence of the Parker-Biermann cooling effect is associated with the so-called thermomagnetic Ettingshausen-Nernst effect (see Apendix A in [14]).
+Due to the large temperature gradient in the tachocline, the thermomagnetic EN effect [16, 17, 18, 19] creates electric currents that are inversely proportional to the strong magnetic field of the tachocline.
+It means that, using the thermomagnetic EN effect, a simple estimate of the magnetic pressure of an ideal gas in the tachocline of e.g.
+the Sun,
 $\frac{B_{tacho}^{2}}{8\pi}=p_{ext}\approx 6.5\cdot 10^{13}\frac{erg}{cm^{3}}~~at~~0.7R_{Sun},$ (1)
@@
 $B_{tacho}^{Sun}=4.1\cdot 10^{7}~G,$ (2)
-Using the strong MFT field, we are interested in the existence of dark matter axions identical to solar axions, which strongly affects the the conversion of axion origin photons in magnetic O-loops near the tachocline, and are thus related to the so-called thermomagnetic Ettingshausen-Nernst effect through the Parker-Biermann cooling effect (see Apendix A in V.
-D.
-Rusov et al.
-(2021)).
-Since the variations of axion origin photons are the result of the changes in Sun luminosity, then unlike a self-excited dynamo, an unexpected yet simple question arises: is there a dark matter chronometer hidden deep in the Sun?
-And consequently, the next question arises: who controls the Sun or who measures the precise “clock” inside the Sun’s core?
-Figure 1: An illustration of the main possible processes of a magnetically active star of the Sun type.
-(a) $\alpha$-effect, $\Omega$-effect and BL mechanism as components of the solar dynamo model.
-The $\Omega$-effect (blue) depicts the transformation of the primary poloidal field into a toroidal field by differential rotation.
-Regeneration of the poloidal field is then performed either by the $\alpha$-effect (top) or by the BL mechanism (yellow in the middle).
-In case of $\alpha$-effect, the toroidal field at the base of the convection zone is subject to cyclonic turbulence.
-In the BL mechanism, the main process of regeneration of the poloidal field (based on the $\Omega$-effect (blue)) is the formation of sunspots on the surface of the Sun from the rise of buoyant toroidal flux tubes from the base of the convection zone.
-The magnetic fields of these sunspots closest to the equator in each hemisphere diffuse and join, and the field due to the spots closer to the poles has a polarity opposite to the current that initiates rotation of the polarity.
-The newly formed polar magnetic flux is transported by the meridional flow to deeper layers of the convection zone, thereby creating a new large-scale poloidal field.
-Derived from S.
-Sanchez et al.
-(2014).
-(b) 11-year variations of sunspot magnetic fields as a component of the solar anti-dynamo.
-The formation of the thermomagnetic EN effect (see L.
-J.
-Spitzer (1962, 2006); V.
-Rusov et al.
-(2015)), emphasizes that this process is associated with the continuous transformation of toroidal field into poloidal one ($T\rightarrow P$ transformation), but not vice versa ($P\rightarrow T$).
-For this purpose, we note (seeV.
-D.
-Rusov et al.
-(2021)) that the modulations of the ADM halo in the Galactic Center (GC), which is an indicator of the periods of S-stars, e.g.
-S-102 with a period of about 11 years, lead (via vertical density waves from the black hole to the solar neighborhood) to modulations of the ADM density, gravitationally trapped in the interior of the Sun.
-This means that anticorrelation modulations of the ADM density control the solar luminosity or sunspot cycles, in which the sunspot variations are identical to the 11-year variations of the magnetic flux tubes (see Section II): maximum (red lines) and minimum (green-red lines).
+Using the strong MFT field, we are interested in the existence of dark matter axions identical to solar axions, which strongly affects the conversion of axion origin photons in magnetic O-loops near the tachocline, and are thus related to the so-called thermomagnetic Ettingshausen-Nernst effect through the Parker-Biermann cooling effect (see Apendix A in [14]).
 In order to answer this question, let us first consider all the unexpected and intriguing implications of the 11-year modulations of the ADM density in the solar interior and around the BH (see Sect.
-3 in V.
-D.
-Rusov et al.
-(2021)).
-A unique result of our model is the fact (see Sect.
-3 in V.
-D.
-Rusov et al.
-(2021)), that the periods, velocities and modulations of the S-stars are the essential indicator of the modulation of the ADM halo density in the fundamental plane of the Galaxy center, which closely correlates with the density modulation of the baryon matter near the supermassive BH.
-If the modulations of the ADM halo at the GC lead to modulations of the ADM density on the surface of the Sun (through vertical density waves from the disk to the solar neighborhood), then there is an ”experimental” anticorrelation identity between such indicators as the ADM density modulation in the solar interior and the number of sunspots.
+3 in [14]).
+It is well known that virtually all physicists, and astrophysicists in particular, know that asymmetric dark matter (ADM) exists in the Universe, and of course, in our galaxy, with the highest density of ADM located precisely at the galactic center, i.e.
+near the black hole.
+At the same time, we know that the periods, velocities, and modulations of the S-stars are the essential indicator of the modulation of the ADM halo density in the fundamental plane of the Galaxy’s center.
+Surprisingly, between the black hole and the disk, there are also so-called amazing S-stars  [^1] Let us remind from where the fairly young S-stars (Fig.
+8a in [14]) appear in our galaxy between the black hole and the disk.
+Since we know that in our galaxy, in addition to the supermassive black hole (SMBH), there is the so-called intermediate black hole (IMBH; see evidence [136]).
+If we take into account the evolution of binary SMBH-IMBH, then when one body (SMBH) exchanges partners of binary stars orbiting the IMBH (Fig.
+8b in [14]), and through the extreme gravitational-tidal field, one star is captured (SMBH) and loses energy, while the other escapes, receives all this energy and, due to the hypervelocity of B-type stars – the IMBH sling (Fig.
+8b in [14]), simply flies out of the galaxy.
+From here it becomes clear how in other galaxies a young S-star can end up between a black hole and a disk., of which there are just over 30.
+Most intriguingly, the closest S-star to the black hole is called S-102, whose orbital period around the black hole is 11 years (see Fig.
+6 and Fig.
+8 in [14]).
+From this we understand that when the S-102 passes very close to the black hole, the high speed of this star is such that dark matter is practically not captured by this star.
+But when the star S-102 orbits farther from the black hole, the star gravitationally captures these ADMs.
+If the modulations of the ADM halo at the GC lead to modulations of the ADM density on the surface of the Sun (through vertical density waves from the disk to the solar neighborhood), then there is an ”experimental” anticorrelation identity between such indicators as the ADM density modulation in the solar interior and the sunspot cycles.
 Or equivalently, between the modulation of solar axions (or photons of axion origin) and the sunspot cycles!
@@
 Meanwhile, these experiments together with the astrophysical and cosmological limitations leave a rather narrow band for the permissible parameters of invisible axion (e.g.
-$10^{-6}eV\leqslant m_{a}\leqslant 10^{-2}eV$ (G.
-G.
-Raffelt, 2004, 2008a)).
-The PQ mechanism, solving the strong CP problem in a very elegant way (R.
-D.
-Peccei & H.
-R.
-Quinn, 1977a, b; F.
-Wilczek, 1978; S.
-Weinberg, 1978), is especially attractive here, since the axion is also a candidate for dark matter (J.
-Preskill et al., 1983; L.
-Abbott & P.
-Sikivie, 1983; M.
-Dine & W.
-Fischler, 1983; M.
-Kawasaki & K.
-Nakayama, 2013; D.
-J.
-Marsh, 2016; L.
-Di Luzio et al., 2020; P.
-Sikivie, 2021).
-At present, we have shown that axions born in the core of the Sun, and photons of axion origin (in the solar corona with a temperature of one million degrees) yield a tight constraint on the axion-photon coupling constant $g_{a\gamma}\sim 4.4\cdot 10^{-11}~GeV^{-1}$ and mass $m_{a}\sim 3.2\cdot 10^{-2}~eV$ (see Fig.
-2a; also Fig.1 in V.
-D.
-Rusov et al.
-(2021)).
-Most importantly, hadronic axions produced in the solar core are controlled by anticorrelated 11-year variations in the density of asymmetric dark matter (ADM) gravitationally trapped in the solar interior (via vertical density waves from the black hole disk to the solar neighborhood), with the hard part of the solar photon spectrum being of axion origin, which is defined as the product of the fraction of axions in the solar core and the fraction of the sunspot area.
+$10^{-6}eV\leqslant m_{a}\leqslant 10^{-2}eV$ (see [21, 22]; Fig.
+13 in [23], Fig.
+24 in [24])).
+The PQ mechanism, solving the strong CP problem in a very elegant way [25, 26, 27, 28], is especially attractive here, since the axion is also a candidate for dark matter [29, 30, 31].
+At present, we have shown that axions born in the core of the Sun, and photons of axion origin (in the million-degree solar corona) yield a tight constraint on the axion-photon coupling constant $g_{a\gamma}\sim 4.4\cdot 10^{-11}~GeV^{-1}$ and mass $m_{a}\sim 3.2\cdot 10^{-2}~eV$ (see Fig.
+1a; also Fig.1 in [14]).
+Most importantly, hadronic axions produced in the solar core are controlled by anticorrelated 11-year variations in the density of asymmetric dark matter (ADM) gravitationally trapped in the solar interior (via vertical density waves from the black hole disk to the solar neighborhood), with the hard part of the solar photon spectrum being of axion origin, which is defined as the product of the axions fraction in the solar core and the fraction of the sunspot area.
 The latter is the result of solar luminosity variations (see Eqs.
-(27)-(28) in V.
-D.
-Rusov et al.
-(2021)).
+(27)-(28) in [14]).
 This means that solar luminosity variations and other solar cycles are the result of 11-year variations in ADM density, which are anticorrelated with the density of hadronic axions in the solar core.
+Interestingly, in contrast to our Fig.
+1a, in Fig.
+1b, we can see a well-known strong constraint on the axion-photon coupling $g_{a\gamma}\sim 4.4\cdot 10^{-11}~GeV^{-1}$, but in a wide mass range (see Fig.
+1b and Eq.
+(2) in [32]) from an analysis of a sample of 39 Galactic Globular Clusters.
+However, in our case (Fig.
+1a) the wide range of masses [32] is, oddly enough, replaced by a single mass $m_{a}\sim 3.2\cdot 10^{-2}~eV$, which fits not only the axions of galactic globular clusters [32], but also all axion stars in our galaxy, including neutron stars [33], supernovae [34, 35], and the Sun [14].
 Figure 1: (a) Summary of astrophysical, cosmological and laboratory constraints on axions.
-Comprehensive axion parameter space, highlighting two main front lines of direct detection experiments: helioscopes (CAST (CAST Collaboration, 2017)) and haloscopes (ADMX (S.
-J.
-Asztalos et al., 2010), and RBF (W.
-U.
-Wuensch et al., 1989) and microwave resonators (B.
-M.
-Brubaker et al., 2017; L.
-Zhong et al., 2018; T.
-Braine et al., 2020; S.
-Lee et al., 2020; K.
-M.
-Backes et al., 2021)).
-The astrophysical bounds from horizontal branch and massive stars are labeled “HB” (G.
-G.
-Raffelt, 2008b) and “Cepheids” (G.
-Carosi et al., 2013), respectively, and there are also astrophysical hints (WD cooling hints, and HB hint).
-The QCD motivated models (KSVZ (J.
-E.
-Kim, 1979; M.
-Shifman et al., 1980)) for axions lay in the yellow diagonal band.
+Comprehensive axion parameter space, highlighting two main front lines of direct detection experiments: helioscopes (CAST [36]) and haloscopes (ADMX [37], and RBF [38] and microwave resonators [39, 40, 41, 42, 43]).
+The astrophysical bounds from horizontal branch and massive stars are labeled “HB” [44] and “Cepheids” [45], respectively, and there are also astrophysical hints (WD cooling hints, and HB hint).
+The QCD motivated models (KSVZ [46, 47]) for axions lay in the yellow diagonal band.
 A plot of $g_{a\gamma}$ versus $m_{a}$ with the most stringent results (solid lines) and sensitivity perspectives (dashed lines) of observations and experiments directly comparable to the different phases of IAXO are shown, BabyIAXO, IAXO, and an upgraded version of IAXO, IAXO+.
 The yellow band denotes the region of the parameter space favoured by QCD axion models.
-The red star marks the values of the axion mass $m_{a}\sim 3.2\cdot 10^{-2}~eV$ and the axion-photon coupling constant $g_{a\gamma}\sim 4.4\cdot 10^{-11}~GeV^{-1}$, which were first obtained experimentally (see V.
-D.
-Rusov et al.
-(2021)).
-(b) $R$ parameter constraints, which compares the numbers of stars in the the horizontal branch (HB) and in the upper portion of the red giant branch (RGB), to helium mass fraction $Y$ and axion coupling $g_{a\gamma}$ (adopted from (A.
-Ayala et al., 2014)).
-The resulting bound on the axion ($g_{10}=g_{a\gamma}/(10^{-10}~GeV^{-1})$ is somewhere between rather conservative $0.5<g_{10}<0.8$ and most aggressive $0.35<g_{10}<0.5$ (A.
-Friedland et al., 2013).
+The red star marks the values of the axion mass $m_{a}\sim 3.2\cdot 10^{-2}~eV$ and the axion-photon coupling constant $g_{a\gamma}\sim 4.4\cdot 10^{-11}~GeV^{-1}$, which were first obtained experimentally (see [14]).
+(b) $R$ parameter constraints, which compares the numbers of stars in the the horizontal branch (HB) and in the upper portion of the red giant branch (RGB), to helium mass fraction $Y$ and axion coupling $g_{a\gamma}$ (adopted from [32]).
+The resulting bound on the axion ($g_{10}=g_{a\gamma}/(10^{-10}~GeV^{-1})$ is somewhere between rather conservative $0.5<g_{10}<0.8$ and most aggressive $0.35<g_{10}<0.5$ [48].
 The red line marks the value of the axion–photon coupling constant $g_{a\gamma}\sim 4.5\cdot 10^{-11}~GeV^{-1}$ adopted from Eq.
-(2) in (A.
-Ayala et al., 2014).
+(2) in [32].
 The blue shaded area represents the bounds from Cepheids observation.
 The yellow star corresponds to $Y=0.254$ and the bounds from HB lifetime (yellow dashed line).
-Interestingly, in contrast to our Fig.
-2a, in Fig.
-2b, we can see a well-known strong constraint on the axion-photon coupling $g_{a\gamma}\sim 4.5\cdot 10^{-11}~GeV^{-1}$, but in a wide mass range (see Fig.
-2b and Eq.
-(2) in (A.
-Ayala et al., 2014)) from an analysis of a sample of 39 Galactic Globular Clusters.
-However, in our case (Fig.
-2a) the wide range of masses in (A.
-Ayala et al., 2014) iso, ddly enough, replaced by a single mass $m_{a}\sim 3.2\cdot 10^{-2}~eV$, which fits not only the axions of galactic globular clusters (A.
-Ayala et al., 2014), but also all axion stars in our galaxy, including neutron stars (M.
-Buschmann et al., 2021), supernovae (P.
-Carenza et al., 2019, 2021), and the Sun (V.
-D.
-Rusov et al., 2021).
 Taking into account the above, we structure our paper as follows.
@@
 Finally, in Section III, we provide a summary and outlook for this paper.
-II Physics of the theory of practically empty magnetic flux tubes at strong fields, $\sim 10^{7}~G$, at which there is a free path for photons of axion origin from the tachocline to the photosphere.
-This raises the fundamental question of the self-exciting action of the dynamo, the remarkable physics of which was discovered some 70 years ago (see E.
-N.
-Parker (1955b, 1957)), beginning with the advent of Babcock’s solar magnetograph (H.
-W.
-Babcock & H.
-D.
-Babcock, 1955), and still today, but which, as no surprisingly, has not received a complete and final solution.
-Avoiding an alternative theory, most physicists believe that it is necessary to study more deeply the known physical problems of various aspects of the theory of solar dynamo (see B.
-B.
-Karak et al.
-(2014)) before passing a verdict on it.
-This continues to this day (see B.
-B.
-Karak (2024)).
-In order for our alternative theory (for example, magnetic flux tubes in strong fields using photons of axion origin) to solve all its problems (see Sections II.1-II.2), it is necessary to solve the most important problem: In what way come sunspots, generated by a strong magnetic field in a flux tube from the base of the convection zone to the surface of the Sun, have 11-year variations without any dynamo?
-The answer is very simple!
-Almost all physicists, and in particular dynamo physicists, know that in the Universe, and, of course, in our galaxy, there are ADM, and the highest density of ADM is located exactly at the center of the Galaxy, i.e.
-near the black hole.
-At the same time, we know that between the black hole and the disk there are also so-called amazing S-stars [^1] Let us remind from where the fairly young S-stars (Fig.
-8a in (V.
-D.
-Rusov et al., 2021)) appear in our galaxy between the black hole and the disk.
-Since we know that in our galaxy, in addition to the supermassive black hole (SMBH), there is the so-called intermediate black hole (IMBH; see evidence (S.
-Takekawa et al., 2019)).
-If we take into account the evolution of binary SMBH-IMBH, then when one body (SMBH) exchanges partners of binary stars orbiting the IMBH (Fig.
-8b in V.
-D.
-Rusov et al.
-(2021)), and through the extreme gravitational-tidal field, one star is captured (SMBH) and loses energy, while the other escapes, receives all this energy and, due to the hypervelocity of B-type stars – the IMBH sling (Fig.
-8b in V.
-D.
-Rusov et al.
-(2021)), simply flies out of the galaxy.
-From here it becomes clear how in other galaxies a young S-star can end up between a black hole and a disk., which have a little more than 30 stars.
-The most interesting thing is that the closest S-star to the black hole is called S-102, which has an 11-yer period of around the black hole (see Fig.
-6 and Fig.
-8 in V.
-D.
-Rusov et al.
-(2021)).
-From here, we understand that when the period of S-102 is very close to the black hole, then the high speed of this star is such that the dark matter is practically not captured by this star.
-But when the elliptical star S-102 rotates farther from the black hole, then the star gravitationally captures these ADMs.
-This means that the 11-year variations in the density of dark matter via vertical density waves from the black hole disk to the solar neighborhood V.
-D.
-Rusov et al.
-(2021) are also captured in the solar interior!
-Hence, the main result is that the 11-year variations in e.g.
-sunspots are never associated with any dynamo action.
-We could raise a few more questions, but it is enough for us that “…not enough for a scientific understanding of the solar dynamo.
-For there is an unfortunate revelation: with four free parameters you can provide a satisfactory fit to the horizon of New York or Beijing” – as was said by the great E.
-N.
-Parker (2009).
-In this regard, let us return to the introduction of our section.
+II Physics of the theory of practically empty flux tubes in strong fields of the order $\sim 10^{7}~G$
 Of all the known concepts that play a significant role in understanding the connection between the energy transfer of a magnetic flux tube and the darkness of sunspots, we will consider what we believe to be the most significant.
-It is based on the Parker-Biermann cooling effect (E.
-N.
-Parker, 1955a; L.
-Biermann, 1941; E.
-N.
-Parker, 1979a) in strong magnetic fields, which explains how the result of the suppression of Parker’s convective heat transfer manifests itself in the lower part of the magnetic flux tube (see Fig.
-3).
-In order to understand the physics of the Parker’s suppression of the convective heat transport in strong magnetic fields, we need to tun to the dark matter axions, born in the core of the Sun.
-With solar axions and the existence of a magnetic O-loop inside the MFT near the tachocline, the answer becomes very simple.
-When a magnetic O-loop is formed inside the MFT near the tachocline through the Kolmogorov turbulent cascade (see Fig.
-4), the high-energy photons from the radiation zone experience the axion-photon oscillations in this O-loop, and the so-called axions of photonic origin appear under the sunspot.
-This means that the cooling effect of Parker-Biermann exists due to the disappearance of barometric equilibrium (E.
-N.
-Parker, 1974a) and, as a consequence, the manifestation of the photon free path (Rosseland length; see Fig.
-B.3 in V.
-D.
-Rusov et al.
-(2021)) from the tachocline to the photosphere, which is confirmed by axions of photonic origin after photon-axion oscillations in the O-loop (see Fig.
-3).
+It is based on the Parker-Biermann cooling effect [6, 7, 8] in strong magnetic fields, which explains how the result of the suppression of Parker’s convective heat transfer manifests itself in the lower part of the magnetic flux tube (see Fig.
+2).
+To understand the physics of the Parker’s suppression of the convective heat transfer in strong magnetic fields, we need to tun to the dark matter axions originating from the Sun’s core.
+Taking into account solar axions and the existence of a magnetic O-loop within the magnetic field near the tachocline (see Fig.
+3d), the answer becomes very simple.
+When a magnetic O-loop forms within the magnetic field near the tachocline via a turbulent Kolmogorov cascade (see Fig.
+3), dark matter axions and high-energy photons from the radiation zone to the tachocline undergo both axion-photon and photon-axion oscillations in this O-loop during the magnetic field in the tachocline, during which so-called of axion origin photons and photon origin axions emerge under a sunspot.
+This means that the Parker-Birman cooling effect exists due to the disappearance of barometric equilibrium [10, 11] and, as a consequence, the manifestation of the photon mean free path (Rosseland length; see Fig.
+B.3 in [14]) from the tachocline to the photosphere, which is confirmed by the existence of both axion-photon oscillations during the formation of axion origin photons, and simultaneously by the existence of photon-axion oscillations during the formation of photon origin axions during MTF in the tachocline (see Fig.
+2).
 Figure 2: Topological effects of magnetic reconnection inside the magnetic tubes with the “magnetic steps” (Fig.
-B.1 in V.
-D.
-Rusov et al.
-(2021)).
-The left panel shows the temperature and pressure change along the radius of the Sun from the tachocline to the photosphere (J.
-Bahcall & M.
-Pinsonneault, 1992), $L_{MS}$ is the height of the magnetic shear steps.
+B.1 in [14]).
+The left panel shows the temperature and pressure change along the radius of the Sun from the tachocline to the photosphere [49], $L_{MS}$ is the height of the magnetic shear steps.
 At $R\sim 0.72~R_{Sun}$ the vertical magnetic field, which is developed from the horizontal part of the magnetic field (step) with the participation of the O-loop through the well-known Kolmogorov turbulent cascade (see Fig.
-4), reaches $B_{z}\sim 3600$ T, and the magnetic pressure $p_{ext}=B^{2}/8\pi\simeq 5.21\cdot 10^{13}~erg/cm^{3}$ (J.
-Bahcall & M.
-Pinsonneault, 1992).
+3), reaches $B_{z}\sim 3600$ T, and the magnetic pressure $p_{ext}=B^{2}/8\pi\simeq 5.21\cdot 10^{13}~erg/cm^{3}$ [49].
 The very cool regions along the entire convective zone caused by the Parker-Biermann cooling effect have the virtually zero internal gas pressure, i.e.
@@
 between the tachocline and the “ring” (or more precisely between the magnetic wall and the O-loop of the tube; see Figs.
-3 and 4), and which is applied without the Parker-Biermann cooling effect, allows to determine the convective heating $(dQ/dt)_{2}$ (see Sect.
+2 and 3), and which is applied without the Parker-Biermann cooling effect, allows to determine the convective heating $(dQ/dt)_{2}$ (see Sect.
 II.1).
-Figure 4: Kolmogorov turbulent cascade (A.
-Kolmogorov, 1941; A.
-N.
-Kolmogorov, 1968, 1991) and primary magnetic reconnection (which differs sharply from the secondary (see Fig.
-5b) in the lower layers inside the unipolar magnetic tube (Fig.
-B.2 in V.
-D.
-Rusov et al.
-(2021)).
-Common to these various turbulent systems is the presence of the inertial region of Kolmogorov, through which the energy is cascaded from large to small scales.
+Figure 3: Kolmogorov turbulent cascade [50, 51] and primary magnetic reconnection (which differs sharply from the secondary (see Fig.
+4) in the lower layers inside the unipolar magnetic tube (Fig.
+B.2 in [14]).
+Common to these various turbulent systems is the presence of the Kolmogorov’s inertial region, through which the energy is cascaded from large to small scales.
 In this case, dissipative mechanisms (as a consequence of the “primary” magnetic reconnection) overcome the turbulent energy during plasma heating.
@@
 That’s why due to the primary magnetic reconnection in the lower layers of the flux tubes (see Fig.
-4d) this is not the final stage of the simulation.
+3d) this is not the final stage of the simulation.
 The essence of a practically empty tube in strong fields of $\sim 10^{7}~G$, which is born for the first time without a dynamo of any type (see Fig.
-5a), is associated with the physics of turbulent reconnection of magnetic bipolar structures (see Fig.
-5b,c).
+4a), is associated with the physics of turbulent reconnection of magnetic bipolar structures (see Fig.
+4b,c).
 It is very important that this is due to the topological effects of secondary magnetic reconnection in the lower layers of the magnetic tube.
 From here, the bipolar part of the $\Omega$-loop rearranges itself at its base, compressing the $\Omega$-loop (blue lines; see Fig.
-5b,c) and, as a consequence, organize a free O-loop (blue lines; Fig.
-5d).
+4b,c) and, as a consequence, organize a free O-loop (blue lines; Fig.
+4d).
 Figure 4: Sketch of primary or secondary magnetic reconnection near the tachocline.
@@
 The blue arrows show the motion of matter leading to the connection of the “legs” of the loop and their “flying away” from the surface of the Sun.
-As described by E.
-N.
-Parker (2009, 1994); H.
-C.
-Spruit et al.
-(1987); P.
-R.
-Wilson et al.
-(1990), the upward convection flow around the rising $\Omega$-loop brings its “legs” together in such a way that the magnetic field reconnection occurs across this loop.
+As described by [52, 53, 54], the upward convection flow around the rising $\Omega$-loop brings its “legs” together in such a way that the magnetic field reconnection occurs across this loop.
 This cuts off the magnetic loop from the azimuthal magnetic field, turning it into an O-loop (see Fig.
-4 in (E.
-N.
-Parker, 1994) and Fig.
-3 in (H.
-C.
-Spruit et al., 1987)).
+4 in [52] and Fig.
+3 in [53]).
 After that the azimuthal magnetic field restores its initial configuration and becomes ready for another process with $\Omega$-loop.
@@
 It is also necessary to explain the physical interpretation of the overshoot process near the tachocline and estimate the velocity $v_{rise}$ and time $\tau_{d}$ of the magnetic tube rise from the overshoot boundary layer – starting with the azimuthal magnetic flux strength of $B_{tacho}\sim 4\cdot 10^{7}~G$ (see Eq.
-A.1 in (V.
-D.
-Rusov et al., 2021)).
+A.1 in [14]).
 One of the final goals of this section is to determine the general regularities of the theory of magnetic flux tubes in strong fields of the order of $\sim 10^{7}~G$ using photons of axion origin – the so-called universal Ballegooijen-Fan-Fisher model (Sec.
 II.1), which are generated by the magnetic buoyancy of practically empty tubes, and as a consequence, their rise from the tachocline to the solar surface (Figs.
-3-5).
+2-4).
 Another goal is the essence of the physics of the primary and secondary reconnection of magnetic flux tubes in the lower layers of the tachocline (Sect.
@@
 II.1-II.2) are caused by the existence (of dark matter?) – solar axions generated in the core of the Sun.
-II.1 Radiative $(dQ/dt)_{1}$ sec-reconnectionand convective $(dQ/dt)_{2}$ heatings and the buoyant rise of magnetic flux tubes and the phenomenon of solar axions
+II.1 Radiative $(dQ/dt)_{1}$ and convective $(dQ/dt)_{2}$ heatings with the concern of the Parker-Biermann cooling effect and axions of photon origin
 The first problem is devoted to the study of the effect of virtually empty magnetic tubes and the phenomenon of solar axions.
 The assumption that the virtually empty magnetic tubes (Fig.
-3) are neutrally buoyant ($\rho_{int}=\rho_{ext}$ (E.
-N.
-Parker, 1994)) implies that the temperature inside these tubes is lower than that of the ambient medium (Fig.
+2) are neutrally buoyant ($\rho_{int}=\rho_{ext}$ [52]) implies that the temperature inside these tubes is lower than that of the ambient medium (Fig.
 2 and Fig.
-7a).
-This leads to the heat inflow, and consequently, the flux tube rises up (see E.
-N.
-Parker (1975) or Sect.
-8.8 in (E.
-N.
-Parker, 1979b)).
-For a horizontal tube with a cross-section of radius $a$ the rise velocity follows from the Parker’s analysis (E.
-N.
-Parker (1975, 1979b), also Eq.
-(60) in (A.
-A.
-van Ballegooijen, 1982)):
+4a).
+This leads to the heat inflow, and consequently, the flux tube rises up (see [55] or Sect.
+8.8 in [56]).
+For a horizontal tube with a cross-section of radius $a$ the rise velocity follows from the Parker’s analysis ([55, 56], also Eq.
+(60) in [57]):
 $v_{rise}=2\frac{H_{p}}{\tau_{d}}\frac{B^{2}}{8\pi p_{ext}}\left(-\delta+0.12\frac{B^{2}}{8\pi p_{ext}}\right)^{-1},$ (3)
-where $H_{p}=\Re T_{ext}/g=p_{ext}/g\rho_{ext}=0.08R_{Sun}$ (E.
-Böhm-Vitense, 1958; H.
-C.
-Spruit, 1977; A.
-S.
-Brun et al., 2011) is the pressure scale height at the tachocline, $T_{ext}$ and $p_{ext}$ are the external gas temperature and pressure, $\delta\equiv Y=\nabla_{e}-\nabla_{ad}=-c_{p}^{-1}dS/d\xi=-c_{p}^{-1}H_{p}dS/dz$ is the dimensionless entropy gradient (see A.
-A.
-van Ballegooijen (1982); R.
-Smolec & P.
-Moskalik (2008, 2010)), $\nabla_{e}\equiv d\ln T_{e}/d\ln p_{e}$ and $\nabla_{ad}\equiv(\partial\ln T_{e}/\partial\ln p_{e})_{s}$ are the local and adiabatic temperature gradients in external and internal plasma (H.
-Spruit, 1974; A.
-A.
-van Ballegooijen, 1982; J.
-Christensen-Dalsgaard et al., 1995), $s$ is the specific entropy, $c_{p}$ is the heat capacity at constant pressure, and $\tau_{d}$ is the rise time of the radiative heating of the magnetic flux tube:
+where $H_{p}=\Re T_{ext}/g=p_{ext}/g\rho_{ext}=0.08R_{Sun}$ [58, 59, 60] is the pressure scale height at the tachocline, $T_{ext}$ and $p_{ext}$ are the external gas temperature and pressure, $\delta\equiv Y=\nabla_{e}-\nabla_{ad}=-c_{p}^{-1}dS/d\xi=-c_{p}^{-1}H_{p}dS/dz$ is the dimensionless entropy gradient (see [57]), $\nabla_{e}\equiv d\ln T_{e}/d\ln p_{e}$ and $\nabla_{ad}\equiv(\partial\ln T_{e}/\partial\ln p_{e})_{s}$ are the local and adiabatic temperature gradients in external and internal plasma [61, 62, 63], $s$ is the specific entropy, $c_{p}$ is the heat capacity at constant pressure, and $\tau_{d}$ is the rise time of the radiative heating of the magnetic flux tube:
 $\tau_{d}=\frac{c_{p}\rho a^{2}}{k_{e}}\simeq c_{p}\rho a^{2}\left[\frac{c_{p}F_{tot}}{g}\left(1+\frac{2\ell_{ov}}{5H_{p}}\right)^{\nu}\right]^{-1}.$ (4)
-where for the fully ionized gas $c_{p}=2.5\Re=2.5p_{ext}/\rho_{ext}T_{ext}$ ($\Re$ is the the equation of state $p_{ext}=\rho_{ext}\Re T_{ext}=nk_{B}T_{ext}$ (E.
-Böhm-Vitense, 1958)), $T(z)$ and $\rho(z)$ are the mean temperature and density; $k_{e}$ is the radiative heat conductivity (see Eq.
-(35) in A.
-A.
-van Ballegooijen (1982)); $\ell_{ov}\approx 0.37H_{p}$ (A.
-A.
-van Ballegooijen, 1982; J.
-Christensen-Dalsgaard et al., 2011) is the thickness of the overshoot layer; the total radiative energy flux $F_{tot}=L/(4\pi r^{2})$ depends on the Sun luminosity $L$; $g$ is the gravitational acceleration.
-Next we apply the condition of hydrostatic equilibrium, $dp/dz=\rho g$, when the adiabatic temperature gradient $(dT/dz)_{ad}=g/c_{p}$ may be used, and the neutral buoyancy of the flux tube in the overshoot zone $(|\delta T|/T_{ext})^{-1}\sim\beta\equiv 8\pi p_{ext}/B^{2}$.
+where for the fully ionized gas $c_{p}=2.5\Re=2.5p_{ext}/\rho_{ext}T_{ext}$ ($\Re$ is the the equation of state $p_{ext}=\rho_{ext}\Re T_{ext}=nk_{B}T_{ext}$ [58]), $T(z)$ and $\rho(z)$ are the mean temperature and density; $k_{e}$ is the radiative heat conductivity (see Eq.
+(35) in [57]); $\ell_{ov}\approx 0.37H_{p}$ [62, 63] is the thickness of the overshoot layer; the total radiative energy flux $F_{tot}=L/(4\pi r^{2})$ depends on the Sun luminosity $L$; $g$ is the gravitational acceleration.
+Next we apply the condition of hydrostatic equilibrium, $dp/dz=-\rho g$, when the adiabatic temperature gradient $(dT/dz)_{ad}=g/c_{p}$ may be used, and the neutral buoyancy of the flux tube in the overshoot zone $(|\delta T|/T_{ext})^{-1}\sim\beta\equiv 8\pi p_{ext}/B^{2}$.
 This way we are able to estimate the time of the radiative and/or convective diffusion $\tau_{d}$ (see Eq.
@@
 $F_{tot}=\frac{L}{4\pi R_{tacho}^{2}}=H_{p}\frac{\nabla_{rad}}{\nabla_{e}}\left(\frac{dQ}{dt}\right)_{1}.$ (7)
-Here $\nabla_{rad}=(\partial\ln T_{ext}/\partial\ln p_{ext})_{rad}$ is the radiative equilibrium temperature gradient; $(dQ/dt)_{1}$ is the rate of radiative heating, which only depends on the thermodynamic parameters $k_{e}$ and $T_{ext}$ of the ambient plasma, depending only on the radial distance from the Sun center (A.
-A.
-van Ballegooijen, 1982; J.
-Christensen-Dalsgaard et al., 1995).
-As a result, it is not difficult to show that the van Ballegooijen model combining equations (3)-(7) gives the final expression for the rise time by radiation heating from the boundary layer of the overshoot to the solar surface,
+Here $\nabla_{rad}=(\partial\ln T_{ext}/\partial\ln p_{ext})_{rad}$ is the radiative equilibrium temperature gradient; $(dQ/dt)_{1}$ is the rate of radiative heating, which only depends on the thermodynamic parameters $k_{e}$ and $T_{ext}$ of the ambient plasma, depending only on the radial distance from the Sun center [57, 62].
+As a result, it is not difficult to show that the van Ballegooijen model combining equations (3)-(7) gives the final expression for the rise time by radiation heating $(dQ/dt)_{1}$ from the boundary layer of the overshoot to the solar surface,
 $\tau_{d}\approx\frac{2}{\beta}T_{ext}\left[\frac{1}{c_{p}\rho_{e}}\left(\frac{dQ}{dt}\right)_{1}\right]^{-1},$ (8)
 at which we have a very important (see section II.2.1) value of
-$\tau_{d}\approx\left.5\frac{1}{\beta}p_{ext}\middle/\left(\frac{dQ}{dt}\right)_{1}\right.\approx 0.84\cdot 10^{8}~sec\approx 2.5~years,$ (9)
+$\tau_{d}\approx\left.5\frac{1}{\beta}p_{ext}\middle/\left(\frac{dQ}{dt}\right)_{1}\right.\approx 0.81\cdot 10^{8}~sec\approx 2~years,$ (9)
 by means of $1/\beta=B^{2}/8\pi\rho_{ext}\approx 2.9\cdot 10^{-5}$, which is identical to the magnetic field $B\approx B_{eq}\sim 2\cdot 10^{5}~G$ (see $B^{2}/8\pi=p_{ext}=nk_{B}T_{ext}\approx 1.665\cdot 10^{13}~erg\cdot cm^{-3}$ (at $0.8R_{Sun}$ (see e.g.
 Fig.
-3)) and $(dQ/dt)_{1}\approx 29.7~erg\cdot cm^{-3}\cdot s^{-1}$ (see Eq.
-(18) in Y.
-Fan & G.
-Fisher (1996)).
+2)) and $(dQ/dt)_{1}\approx 29.7~erg\cdot cm^{-3}\cdot s^{-1}$ (see Eq.
+(18) in [64]).
 At the same time, the rate of the MFT rise from the tachocline to the solar surface
-$v_{rise}=H_{p}\nabla_{ad}\frac{1}{p_{ext}}\left(\frac{dQ}{dt}\right)_{1}\left(|\delta|+0.12\frac{B^{2}}{8\pi p_{ext}}\right)^{-1},~~~\nabla_{ad}=\nabla_{e}\simeq 0.4,$ (10)
+$v_{rise}=H_{p}\nabla_{ad}\frac{1}{p_{ext}}\left(\frac{dQ}{dt}\right)_{1}\left(|\delta|+0.12\frac{B^{2}}{8\pi p_{ext}}\right)^{-1},$
+$\nabla_{ad}=\nabla_{e}\simeq 0.4,$ (10)
 at which we have a very important (see Sections II.1-II.2) value of
@@
 by means of $H_{p}\approx 5.6\cdot 10^{4}~km$ and $(|\delta|+0.12B^{2}/8\pi p_{ext})=(3/5)\times B^{2}/8\pi p_{ext}+0.12\times B^{2}/8\pi p_{ext}\approx 0.72\times 2.9\cdot 10^{-5}$ (see Eqs.
-(58-60) in (A.
-A.
-van Ballegooijen, 1982)), are almost identical to the equations (29)-(30) of Y.
-Fan & G.
-Fisher (1996).
+(58-60) in [57]), are almost identical to the equations (29)-(30) of [64].
 At the same time, it is easy to show that these values
@@
 provide a simple relationship between the rise time and the rate of ascent of the magnetic flux tube from the overshoot boundary layer to the surface of the Sun.
-Surprisingly, these equations (Eqs.
-(8) and (10)) are almost identical to equations (29) and (30) from Y.
-Fan & G.
-Fisher (1996), which arise from completely different equations (60)-(61) of J.
-Bahcall & M.
-Pinsonneault (1992).
+Surprisingly, these equations (Eqs. (8) and (10)) are almost identical to equations (29) and (30) from [64], which arise from completely different equations (60)-(61) of [49].
 Unlike the special van Ballegooijen model, we adopt the universal model of MFTs with
@@
 $\tau_{d}=\frac{c_{p}\rho a^{2}}{k_{e}}\left\{\left(\frac{dQ}{dt}\right)_{1}\left[1+\left(\frac{dQ}{dt}\right)_{2}\middle/\left(\frac{dQ}{dt}\right)_{1}\right]\right\}^{-1},$
-Here the second term $(dQ/dt)_{2}$ represents a convective diffusion across the flux tube that is due to the temperature difference ($\delta T\equiv T-T_{e}$) between the flux tube and the external plasma (see (Y.
-Fan & G.
-Fisher, 1996)).
+Here the second term $(dQ/dt)_{2}$ represents a convective diffusion across the flux tube that is due to the temperature difference ($\delta T\equiv T-T_{e}$) between the flux tube and the external plasma (see [64]).
 Using simple calculations of equations (3) and (II.1) for MFTs, it is easy to show that with the help of the
@@
 and
-$v_{rise}=H_{p}\nabla_{ad}\frac{1}{p_{ext}}\frac{dQ}{dt}\left(-\delta+0.12\frac{B^{2}}{8\pi p_{ext}}\right)^{-1},~~~\nabla_{ad}=\nabla_{e}\simeq 0.4,$ (16)
-is the general case of the so-called universal model of the van Ballegooijen-Fan-Fisher, which is completely identical to equation (27) of the model by Y. Fan & G. Fisher (1996):
-$\frac{d\Delta\rho}{dt}=\rho_{e}\frac{v_{rise}}{H_{p}}\left[\delta+\left(\frac{1}{\gamma}-\frac{2}{\gamma^{2}}\right)\frac{1}{\beta}-\frac{1}{\gamma}\frac{\Delta\rho}{\rho_{e}}\right]+\frac{2}{\gamma\beta}\rho_{e}\left[\frac{\partial(v\cdot I)}{\partial s}-\vec{v}\cdot\vec{k}\right]+\frac{\rho_{e}}{p_{e}}\nabla_{ad}\left[\left(\frac{dQ}{dt}\right)_{1}+\left(\frac{dQ}{dt}\right)_{2}\right],$ (17)
-where in case of quasi-static rising, it is necessary apply ($d\Delta\rho/dt$) and $\Delta\rho$ to zero (see equation (27) of Y.
-Fan & G.
-Fisher (1996)).
-Furthermore, according to Y.
-Fan & G.
-Fisher (1996), for simplicity, the lifting of a uniform horizontal magnetic flux tube in the region of a plane-parallel overshoot region is considered, under which the conditions $\partial(v\cdot I)/\partial s=0$ and $\vec{v}\cdot\vec{k}=0$ can be used in equation (27) of Y.
-Fan & G.
-Fisher (1996): $I\equiv\partial r/\partial s$ is the unit vector tangential to the flux tube, and $k\equiv\partial^{2}r/\partial s^{2}$ is the tube’s curvature vector.
+$v_{rise}=H_{p}\nabla_{ad}\frac{1}{p_{ext}}\frac{dQ}{dt}\left(-\delta+0.12\frac{B^{2}}{8\pi p_{ext}}\right)^{-1},$
+$\nabla_{ad}=\nabla_{e}\simeq 0.4,$ (16)
+is the general case of the so-called universal model of the van Ballegooijen-Fan-Fisher, which is completely identical to equation (27) of the model [64]:
+$\frac{d\Delta\rho}{dt}$ $=\rho_{e}\frac{v_{rise}}{H_{p}}\left[\delta+\left(\frac{1}{\gamma}-\frac{2}{\gamma^{2}}\right)\frac{1}{\beta}-\frac{1}{\gamma}\frac{\Delta\rho}{\rho_{e}}\right]+$ (17)
+$\frac{2}{\gamma\beta}\rho_{e}\left[\frac{\partial(v\cdot I)}{\partial s}-\vec{v}\cdot\vec{k}\right]+\frac{\rho_{e}}{p_{e}}\nabla_{ad}\left[\left(\frac{dQ}{dt}\right)_{1}+\left(\frac{dQ}{dt}\right)_{2}\right],$
+where in case of quasi-static rising, it is necessary apply ($d\Delta\rho/dt$) and $\Delta\rho$ to zero (see equation (27) of [64]).
+Furthermore, according to [64], for simplicity, the lifting of a uniform horizontal magnetic flux tube in the region of a plane-parallel overshoot region is considered, under which the conditions $\partial(v\cdot I)/\partial s=0$ and $\vec{v}\cdot\vec{k}=0$ can be used in equation (27) of [64]: $I\equiv\partial r/\partial s$ is the unit vector tangential to the flux tube, and $k\equiv\partial^{2}r/\partial s^{2}$ is the tube’s curvature vector.
 It should also be recalled that, taking into account the presence of subadiabaticity $\delta<0$ (see Eq.
-(58) in (A.
-A.
-van Ballegooijen, 1982), Eq.
-(27) in (Y.
-Fan & G.
-Fisher, 1996), and also Fig.
-6) of the overshoot tachocline and using, according to A.
-A.
-van Ballegooijen (1982), non-local mixing length theory (H.
-C.
-Spruit & A.
-A.
-van Ballegooijen, 1982; E.
-Böhm-Vitense, 1958; S.
-Z.
-Ali & S.
-Dey, 2020; M.
-A.
-Weber & Y.
-Fan, 2015), it can be shown that thin, neutrally buoyant flow tubes, stable in a stratified medium, provided that its field strength $B$ is smaller than a critical value $B_{c}$ (A.
-A.
-van Ballegooijen, 1982), which is approximately given by
+(58) in [57], Eq.
+(27) in [64], and also Fig.
+5) of the overshoot tachocline and using, according to [57], non-local mixing length theory [57, 65, 58, 66, 67], it can be shown that thin, neutrally buoyant flow tubes, stable in a stratified medium, provided that its field strength $B$ is smaller than a critical value $B_{c}$ [57], which is approximately given by
 $\frac{B_{c}^{2}}{8\pi p_{ext}}=-\gamma\delta=-\frac{5}{3}\delta.$ (18)
@@
 where we used the MFT heating rate which consists of the radiative heating rate $(dQ/dt)_{1}$ and convective heating rate $(dQ/dt)_{2}$ (see Eqs.
-(II.1)–(19), Eqs.
-(13) and (19) in (Y.
-Fan & G.
-Fisher, 1996) and Eqs.
-(8)–(9) in (M.
-A.
-Weber & Y.
-Fan, 2015)); the approximate ratio $|\delta T|/T_{ext}\sim 1/\beta$; the parameter $\alpha_{1}^{2}\approx 5.76$ (Y.
-Fan & G.
-Fisher, 1996; M.
-A.
-Weber & Y.
-Fan, 2015); $F_{rad}$ is the radiative energy flux (see Eq.
-(15) in (Y.
-Fan & G.
-Fisher, 1996), and also Eq.
+(19)–(21), Eqs.
+(13) and (19) in [64] and Eqs.
+(8)–(9) in [67]); the approximate ratio $|\delta T|/T_{ext}\sim 1/\beta$; the parameter $\alpha_{1}^{2}\approx 5.76$ [64, 67]; $F_{rad}$ is the radiative energy flux (see Eq.
+(15) in [64], and also Eq.
 (6) and Fig.
-1 in (M.
-A.
-Weber & M.
-K.
-Browning, 2016)); $\nabla_{e}\sim 1.287\nabla_{rad}$ (see Table 2 in (H.
-C.
-Spruit, 1977)); $H_{p}/a$ is the factor for the lower convection zone (see (R.
-M.
-Roettenbacher et al., 2016)), where it was previously believed that $a=(\Phi/\pi B_{tacho})^{1/2}\leqslant 0.1H_{p}$ (see 3.2.1 in (R.
-M.
-Roettenbacher et al., 2016)) with an average value of typical magnetic flux of $\Phi\sim 10^{20}-10^{22}~Mx$ (Y.
-Fan et al., 1993).
+1 in [68]); $\nabla_{e}\sim 1.287\nabla_{rad}$ (see Table 2 in [56]); $H_{p}/a$ is the factor for the lower convection zone (see [69]), where it was previously believed that $a=(\Phi/\pi B_{tacho})^{1/2}\leqslant 0.1H_{p}$ (see 3.2.1 in [57], 5.6.1 in [69]) with an average value of typical magnetic flux of $\Phi\sim 10^{20}-10^{22}~Mx$ [70, 71, 72].
 From this, we understand that it is with the strong fields that a sharp increase in convective heating $(dQ/dt)_{2}$ is necessary in the tube, which simultaneously leads to a sharp decrease in the outer width of the ring $a\equiv a_{conv}$ (between the O-loop and the tube walls in Fig.
-6) in a practically empty magnetic flux tube (see Fig.
+4a) in a practically empty magnetic flux tube (see Fig.
 2 and Fig.
-6).
+4a).
 In this case, the area of the “ring” of the magnetic tube is approximately equal to $2\pi r\times a_{conv}$, where $r$ is the radius of the magnetic tube, and $a_{conv}$ is the width of the “ring”.
 What is the physics behind the appearance of the “ring” between the O-loop and the walls of the magnetic tube (see Fig.
-6; also Fig.
-3)?
+5; also Fig.
+2)?
 In simple words, one can say the following.
@@
 The keV photons (see Fig.
-2 in J.
-E.
-Bailey et al.
-(2009)) coming from the radiation zone are turned into axions in the horizontal magnetic field of the O-loop (see Fig.
-3).
+2 in [73]) coming from the radiation zone are turned into axions in the horizontal magnetic field of the O-loop (see Fig.
+2).
 Therefore, the radiative heating almost vanishes in the virtually empty magnetic tube.
@@
 Meanwhile, the external plasma turns from sub-adiabatic to super-adiabatic ($\delta=\nabla_{e}-\nabla_{ad}>0$).
-At the same time, we remember that anticorrelated 11-year ADM density modulations in the Sun interior (see Fig.
-1b) and the ADM around BH (see Sect.
-3 in (V.
-D.
-Rusov et al., 2021)) are not only interconnected, but simultaneously, as consequence, are driving the 11-year modulations of the solar axion density (see Eqs.
-(27)-(28) in (V.
-D.
-Rusov et al., 2021)).
-As a result, assuming 11-year variations in convective heating $(dQ/dt)_{2}$ in the form of a maximum and minimum of the “ring” width of the $a\equiv a_{conv}$ (between the O-loop and the tube walls in Fig.
-6)
-$(a_{conv})_{max}\sim 3.7\cdot 10^{-4}H_{p}\approx 22~km,$ (22)
-$(a_{conv})_{min}\sim 3.7\cdot 10^{-3}H_{p}\approx 220~km,$ (23)
-we apply a new analysis of the universal vanBFF model (see Eqs.
-(32.1)-(18) analogous to (A.
-A.
-van Ballegooijen, 1982) or (19)-(21) (analogous to the last part of Eq.
-(13) and (19) in (Y.
-Fan & G.
-Fisher, 1996))), where the calculated values such as the magnetic flux $\Phi$ and the rise speed $(v_{rise})_{conv}$ of the MFT to the surface of the Sun, do not contradict the known observational data:
+At the same time, we remember that anticorrelated 11-year ADM density modulations in the Sun interior and the ADM around BH (see Sect.
+3 in [14]) are not only interconnected, but simultaneously, as consequence, are driving the 11-year modulations of the solar axion density (see Eqs.
+(27)-(28) in [14]).
+As a result, assuming 11-year variations in convective heating $(dQ/dt)_{2}$ in the form of a maximum and minimum in the width of the “ring” $a\equiv a_{conv}$ (between the O-loop and the tube walls in Fig.
+5), we apply a new analysis of the universal vanBFF model (see Eqs.
+(19)-(21) similar to the last part of Eqs.
+(13) and (19) in [64])), where the values such as the magnetic flux $\Phi$ (see [71, 72]) and the magnetic fields at the tachocline ($3.6\cdot 10^{7}~G$ in Fig.
+2) are consistent with known observational data:
 $\bullet$ values of the magnetic tube (see Fig.
-5a), based on the physics of convective heating $(dQ/dt)_{2}$, are related to the radius of the “ring” between the O-shaped contour and the walls of the magnetic tube (Fig.
-6), and which are also the result of maximum magnetic flux
+4a), based on the physics of convective heating $(dQ/dt)_{2}$, are related to the radius of the “ring” between the O-shaped loop and the walls of the magnetic tube (Fig.
+5), and correspond exactly to the “observed” values of the ratio between the maximum and minimum magnetic flux and the magnetic field at the tachocline [71, 72, 69].
+Therefore, we are interested in the observed magnetic fluxes
 $\Phi_{max}=2\pi r(a_{conv})_{min}(B_{tacho})_{max}\approx 10^{22}~Mx,$ (22)
-where $r\sim 220~km$, $(a_{conv})_{min}\sim 22~km$ and $(B_{tacho})_{max}\sim 3.6\cdot 10^{7}~G$, and minimum magnetic flux
-$\Phi_{min}=\pi(a_{conv})_{max}^{2}(B_{tacho})_{min}\approx 5\cdot 10^{21}~Mx,$ (25)
-where $\pi(a_{conv})_{max}^{2}\approx\pi(220^{2}-20^{2})~km^{2}=1.5\cdot 10^{15}~cm^{2}$ and $(B_{tacho})_{min}\sim 3.6\cdot 10^{6}~G$, and which is in good agreement with the observational data of C.
-Zwaan (1987) (see also Section 2.1 in (Y.
-Fan, 2021)) since the total magnetic flux ranges are $1-5\cdot 10^{22}~Mx$ (near the maximum) and $2-5\cdot 10^{21}~Mx$ (near the minimum) (see (K.
-L.
-Harvey et al., 1982)).
+where $\Phi_{max}\approx 10^{22}~Mx$ is at the maximum of the “measured” magnetic fluxes [71, 72] and the maximum of the “measured” magnetic field MFT $(B_{tacho})_{max}\sim 3.7\cdot 10^{7}~G$ (see Fig.
+2), which is associated with the so-called Ettingshausen-Nernst thermomagnetic effect, caused by the Parker-Biermann cooling effect.
+Then a minimum of magnetic flux appears
+$\Phi_{min}=2\pi r(a_{conv})_{max}(B_{tacho})_{min}\approx 5\cdot 10^{21}~Mx,$ (23)
+where $\Phi_{min}\approx 5\cdot 10^{21}~Mx$ is at the minimum of the “measured” magnetic fluxes [71, 72] and the minimum of the “measured” magnetic field MFT $(B_{tacho})_{min}\sim 3.7\cdot 10^{6}~G$.
 It is very important here that, on the one hand, the anchored magnetic field in the tachocline rises from the bottom of the convective zone to the solar photosphere in the form of thin isolated filaments known as flux tubes (see e.g.
-(E.
-N.
-Parker, 1979a)), and on the other hand, the well-known nonlinear dynamics of a thin magnetic tube was modeled in the form that has been used by H.
-C.
-Spruit (1981).
+[56]), and on the other hand, the well-known nonlinear dynamics of a thin magnetic tube was modeled in the form that has been used by [59].
 In this case, the magnetic tube is assumed to be thin in the sense that its cross-sectional radius $a_{conv}$ is negligibly small with respect to both the atmospheric scale height (i.e.
 $a_{conv}\ll H_{p}$ (see Fig.
-6)) and any variation scales along the tube.
+5)) and any variation scales along the tube.
 This means that each magnetic field line is roughly parallel to the axis of the tube, and the Spruit’s model is applicable only to untwisted magnetic tubes.
-This raises a rather complicated but simple question: In what way, without twisting, can a real magnetic tube, which is not integral and will not behave as a single object for a very long time (E.
-N.
-Parker, 1979a; H.
-C.
-Spruit, 1981; D.
-W.
-Longcope & I.
-Klapper, 1997), be obtained without destruction by hydrodynamic forces?
+This raises a rather complicated but simple question: In what way, without twisting, can a real magnetic tube, which is not integral and will not behave as a single object for a very long time [56, 59], be obtained without destruction by hydrodynamic forces?
 The answer, oddly enough, is very simple!
 For an untwisted thin magnetic tube, the pressure balance is always maintained not only along the entire radius of the “ring” between the O-shaped contour and the walls of the magnetic tube (Fig.
-6), which is solved without destruction by hydrodynamic forces, but at the same time leads to the suppression of convection along the inner radius of the O-shaped contour due to a sharp decrease in temperature and density (see $\rho_{int}^{0}\gg\rho_{int}$ in Fig.
-6).
+5), which is solved without destruction by hydrodynamic forces, but at the same time leads to the suppression of convection along the inner radius of the O-shaped contour due to a sharp decrease in temperature and density (see $\rho_{int}^{0}\gg\rho_{int}$ in Fig.
+5).
 The latter means that in very strong magnetic fields (e.g.
-$5\cdot 10^{5}~G<B<10^{8}~G$) an untwisted thin magnetic tube will, surprisingly, always be perpendicular to the azimuthal direction from the tachocline to the surface of the photosphere!
+$5\cdot 10^{5}~G<B<10^{7}~G$) an untwisted thin magnetic tube will, surprisingly, always be perpendicular to the azimuthal direction from the tachocline to the surface of the photosphere!
 The physics of this process is briefly as follows.
@@
 As a result, it opens a free path (in the radial direction!) for photons of axion origin (Rosseland length; see Fig.
-B.3 in (V.
-D.
-Rusov et al., 2021)) from the tachocline to the photosphere (see Fig.
-5a)!
-$\bullet\bullet$ the time of rise of convective heating of the flux tube from the tachocline to the photosphere (see Eq. (15)), which corresponds to the equations
+B.3 in [14]) from the tachocline to the photosphere (see Fig.
+4a)!
+From here, using Eq. (22) and Eq. (23), we obtain the cross-section values $a_{conv}$ (between the O-loop and the tube walls in Fig. 4a) at the minimum
+$(a_{conv})_{min}\sim 22~km\approx 3.7\cdot 10^{-4}H_{p}$ (24)
+and at the maximum
+$(a_{conv})_{max}\sim 220~km\approx 3.7\cdot 10^{-3}H_{p},$ (25)
+at which the rise time of convective heating of the flux tube from the tachocline to the photosphere (see Eq. (15)) corresponds to the equations
 $(\tau_{d})_{conv}=\frac{2}{\beta}T_{ext}\left\{\frac{1}{c_{p}\rho_{ext}}\left(\frac{dQ}{dt}\right)_{1}\left[1+\frac{\alpha_{1}^{2}}{\nabla_{e}}\frac{1}{\beta}\left(\frac{H_{p}}{a_{conv}}\right)^{2}\right]\right\}^{-1}$ (26)
@@
 $\cong\frac{5}{\beta}\rho_{ext}\mu^{-1}m_{p}^{-1}k_{B}T_{ext}\left\{\left(\frac{dQ}{dt}\right)_{1}\left[1+\frac{\alpha_{1}^{2}}{\nabla_{e}}\frac{1}{\beta}\left(\frac{H_{p}}{a_{conv}}\right)^{2}\right]\right\}^{-1}$ (29)
-at the maximum value of the 11-year variations of the magnetic flux tube in the photosphere for sunspots (see (27)), in the form
+where at the maximum the values of $(a_{conv})_{min}\approx 3.7\cdot 10^{-4}H_{p}$ and $1/\beta=B^{2}/8\pi p_{ext}\approx 1$ we represent 11-year variations of the magnetic flux tube in the photosphere for sunspots (see (28) or (27), in the form
 $(\tau_{d})_{conv}^{max}\approx 5\times(6.5\cdot 10^{13}~erg\cdot cm^{-3})/3\cdot 10^{9}~erg\cdot cm^{-3}\cdot s^{-1}$
@@
 $(\tau_{d})_{conv}^{max}\approx 1.1\cdot 10^{5}~s\sim 1.3~day,$ (31)
-where $p_{ext}=6.53\cdot 10^{13}~erg/cm^{3}$ (see A.16 in (V.
-D.
-Rusov et al., 2021), and also (J.
-Bahcall & M.
-Pinsonneault, 1992)), $T_{ext}=2.3\cdot 10^{6}~K$ and $\rho_{ext}=0.2~g/cm^{3}$ (J.
-Bahcall & M.
-Pinsonneault, 1992); $c_{p}=2.5\Re=2.5p_{ext}/\rho_{ext}T_{ext}$ ($\Re$ is the gas constant in the equation of state $p_{ext}=\rho_{ext}\Re T_{ext}=nk_{B}T_{ext}$ (E.
-Böhm-Vitense, 1958)); $\mu=0.58$ is mean molecular weight of hydrogen in overshoot tachocline, $m_{p}\cong 1.67\cdot 10^{-24}~g$ is the proton mass: ”2018 CODATA Value: proton mass”, The NIST Reference on Constants, Units, and Uncertainty, 20 May 2019; $k_{B}=1.38\cdot 10^{-23}~J/K$; $(dQ/dt)_{1}\approx 29.7~erg\cdot cm^{-3}\cdot s^{-1}$ (see Eq.
-(18) in (Y.
-Fan & G.
-Fisher, 1996)) is the rate of radiative heating; $\alpha_{1}^{2}\approx 5.76$ (Y.
-Fan & G.
-Fisher, 1996; M.
-A.
-Weber & Y.
-Fan, 2015); $\nabla_{e}\approx 0.4$ (see Table 2 in (H.
-Spruit, 1974)); $H_{p}=\Re T_{ext}/g=p_{ext}/g\rho_{ext}=0.08R_{Sun}$ (E.
-Böhm-Vitense, 1958; H.
-C.
-Spruit, 1977; A.
-S.
-Brun et al., 2011); $a_{conv}^{max}\sim 3.7\cdot 10^{-4}H_{p}$ (see Eq.
+where $p_{ext}=6.53\cdot 10^{13}~erg/cm^{3}$ (see [49]), $T_{ext}=2.3\cdot 10^{6}~K$ and $\rho_{ext}=0.2~g/cm^{3}$ [49]; $c_{p}=2.5\Re=2.5p_{ext}/\rho_{ext}T_{ext}$ ($\Re$ is the gas constant in the equation of state $p_{ext}=\rho_{ext}\Re T_{ext}=nk_{B}T_{ext}$ [58]); $\mu=0.58$ is mean molecular weight of hydrogen in overshoot tachocline, $m_{p}\cong 1.67\cdot 10^{-24}~g$ is the proton mass: ”2018 CODATA Value: proton mass”, The NIST Reference on Constants, Units, and Uncertainty, 20 May 2019; $k_{B}=1.38\cdot 10^{-23}~J/K$; $(dQ/dt)_{1}\approx 29.7~erg\cdot cm^{-3}\cdot s^{-1}$ (see Eq.
+(18) in [64]) is the rate of radiative heating; $\alpha_{1}^{2}\approx 5.76$ [64, 67]; $\nabla_{e}\approx 0.4$ (see Table 2 in [61]); $H_{p}=\Re T_{ext}/g=p_{ext}/g\rho_{ext}=0.08R_{Sun}$ [58, 59, 60]; $a_{conv}^{min}\sim 3.7\cdot 10^{-4}H_{p}$ (see Eq.
 (24)).
-In order to obtain the minimum values of the 11-year variations of the magnetic flux tube in the photosphere for sunspots, we have to calculate a new value for $p_{ext}=(nk_{B}T)_{min}$, which together with $a_{conv}\sim 3.7\cdot 10^{-3}~H_{p}$ (Eq.
-(23)) is related to Eqs.
-(32)-(33).
-Using the value of fully ionized hydrogen plasma with $Z=1$ in the tachocline, which is related to the equation $T^{1/4}n=const$ (see Eq.
-(A.9) from Appendix A in (V.
-D.
-Rusov et al., 2021)), we obtain the relation
-$(T^{1/4}\rho_{ext})_{min}=(T^{1/4}\rho_{ext})_{max}=((2.3\cdot 10^{6}~K)^{1/4}\times 0.2~g/cm^{3})_{max}.$ (32)
-When we apply the temperature “minimum” value $(T\sim 10^{6}~K)_{min}$ (see Eq. (28) from (V. D. Rusov et al., 2021)), then we get the following relation
-$\mu\times m_{p}\times(T^{1/4}n)_{min}=((10^{6}~K)^{1/4}\rho_{ext})_{min}=\mu\times m_{p}\times(T^{1/4}n)=((2.3\cdot 10^{6}~K)^{1/4}\times 0.2~g/cm^{3})_{max},$ (33)
-from which we obtain the complete solution of the “minimum” rise time of the convective heating of the flux tube from the tachocline to the photosphere (see Eq. (15), which corresponds to the equation
-$(\tau_{d})_{conv}^{min}\approx 5.7\cdot 10^{5}~s\sim 7~days,$ (34)
-where $p_{ext}=(nk_{B}T)_{min}=3.56\cdot 10^{13}~erg/cm^{3}$, $T_{ext}=10^{6}~K$, $\rho_{ext}\approx 0.25~g/cm^{3}$ (see Eq. 33).
 And, as a consequence, the magnitude of the MFT rise speed (see Eq. (16)) will be equal to
@@
 $(v_{rise})_{conv}^{max}\sim 1.40~km/s,$ (33)
-which are almost identical to the observational data of the known works (see about 1.4 m/s at 62-75 and 42-55 Mm for AR11726, April 2013 in A.
-G.
-Kosovichev et al.
-(2016, 2018); see about 1.2 km/s at 16-32 Mm for AR NOAA 6432, South Pole in 1991 January 4-5, in A.
-Kosovichev (1996)), and at the minimum of the solar cycle
-$(v_{rise})_{conv}^{min}\sim 0.25~km/s,$ (37)
-which are also almost identical to the observational data, e.g.
-of the known works by S.
-Ilonidis et al.
-(2011, 2012, 2013).
-This is due to the fact that these works, on the one hand, showed the measured velocities of the magnetic flux escape from the deep interior of the Sun into the atmosphere in the form of sunspots (see $\sim 1~km/s$ for 2003 October 25 (Fig.
-5.14 in S.
-Ilonidis et al.
-(2012); Fig.
-8 in S.
-Ilonidis et al.
-(2013))), and on the other hand, using deep-focus time-distance helioseismology T.
-L.
-Duvall et al.
-(1993, 1996); A.
-G.
-Kosovichev et al.
-(2016) found several strong emerging flux events (for AR 10488, October 26, 2003, + ARs 8164 and 8171, 23-27 February 1998 + AR 7978, 06 July 1996) at depths of 42-75 Mm up to 6 h (using only previous data) before entering the photosphere and 1-2 days before until the appearance of the detected magnetic structures on the surface with an average speed of 0.3 to 0.6 km/s (S.
-Ilonidis et al., 2011).
-This means that, unlike the works of S.
-Ilonidis et al.
-(2011, 2012, 2013), we use the velocities of emerging flows on the surface of sunspots between the maximum ((31) and (36)) and minimum ((34) and (37)) of the solar cycle (for AR 10488, October 26, 2003, + ARs 8164 and 8171, 23-27 February 1998 + AR 7978, 06 July 1996), which correspond close to the maximum $\sim 1~km/s$ (for 2003 October), lower (by assumption!) than middle of the maximum and minimum 0.3 km/s +0.3 km/s (for ARs 8164 and 8171, 23-27 February 1998) and low $\sim 0.25~km/s$ (for AR 7978, 06 July 1996).
-From here, we get the average value of the rate of rise of the magnetic tube from the maximum (for 2003 October) to the minimum (for July 1996)
-$\frac{1.0(2003)+0.30(1998)+0.30(1998)+0.25(1996)}{4}km/s\sim 0.45~km/s$ (38)
-which due to our mean velocity uncertainty of $45_{-0.15}^{+0.15}~km/s$ almost exactly corresponds to the known uncertainty in the average velocity from 0.3 to 0.6 km/s (S. Ilonidis et al., 2011)!
-In other words, this means that our (maximum (31) and (36)) and minimum ((34) and (37)) are the remarkable existence of a maximum and minimum of 11-year variations of the photosphere magnetic flux tube for sunspots, which, unlike the solar dynamo, are predetermined as flux tube magnetic field sources (anchored in the overshoot tachocline), and are coupled by the holographic mechanism of quantum gravity (see Fig.
-1b), and anticorrelated 11-year ADM density variations gravitationally trapped in the interior of the Sun (see Section 3 in (V.
-D.
-Rusov et al., 2021)).
-In order to show signatures of appearing areas of sunspots inside the Sun before they appear on the surface, we used, according to A.
-G.
-Kosovichev et al.
-(2016), the measurements of plasma flows in the upper convection zone, provided by the Time-Distance Helioseismology pipeline developed for the analysis of solar oscillation data obtained by the Helioseismic and Magnetic Imager (HMI) aboard the Solar Dynamics Observatory (SDO), to investigate the subsurface dynamics of the emerging active region (AR) NOAA 11726.
-Figure 7c shows the distribution of the effective phase shift (“helioseismic index”) for the depth range 62-75 Mm measured on April 19, 2013, 03:00 UT, when there was no significant magnetic flux on the surface.
-The figure 7d shows the variations of the helioseismic perturbation associated with the emerging AR at two different depths: 62-75 Mm and 42-55 Mm depending on time.
-One of the greatest advantages of the remote helioseismology is that the full-disk observations of solar variations allow to build a special scheme of measurement and selection of acoustic waves signals (Fig.
-7b), passing though certain sub-surface areas.
-For example, S.
-Ilonidis et al.
-(2013) developed a special procedure of deep focusing which is able to detect large emerging active areas more than a day before their appearance on the surface (Fig.
-7).
-The rise velocity determined by tracking the magnetic flux declination with depth, is about 1.4 km/s, which is very close to the rise velocity in the deep layers A.
-G.
-Kosovichev et al.
-(2016), e.g.
-near tachocline!
-Interestingly, the results of several attempts to detect the emerging magnetic flux before its appearance in the photosphere may be compared to our theoretical estimates (maximum (31) and (36)) and minimum ((34) and (37)) and demonstrate a surprising agreement in terms of the magnetic tube rising velocity and time.
-Figure 7: Detection of emerging active region AR 11726: a) magnetograms of helioseismic and magnetic imager (HMI) on board of the Solar Dynamics Observatory (SDO) of April 19, 2013, 3:00 UT, prior the emergence; the rings show the area where the acoustic oscillation signal was measured to detect the subsurface signal at the central point; b) a vertical cut through the computational grid used in the time-distance inversions, and a sample of ray paths of acoustic waves, which were used for measuring travel times (A.
-G.
-Kosovichev & J.
-Duvall, 2008; T.
-L.
-Duvall et al., 1993, 1996); c) variations of a travel-time index showing the AR perturbation located the depth of 62-75 Mm on April 19, 2013, 3:00 UT; d) variations of the helioseismic perturbation associated with the emerging AR at two different depths: 42-55 Mm and 62-75 Mm, as a function of time.
-The characteristic rise time estimated using the delays in two layers at depths 42-55 Mm and 62-75 Mm is about 1.4 km/s.
-Derived from A.
-G.
-Kosovichev et al.
-(2016).
-Below we briefly and simply show how our fundamental physics of the rise of strong magnetic flux tubes with $B\gg 10^{5}~G$, but without dynamo action, differs from the classical physics of an active region, in which the rise of magnetic tubes from a tachocline to the photosphere is determined by the result of the convective action of the dynamo in the main part of the solar convection zone (with 0.15-0.25 km/s; see e.g.
-Y.
-Fan (2021); P.
-Charbonneau (2020); P.
-Charbonneau & D.
-Sokoloff (2023) and Fig.
-21 of H.
-Hotta & H.
-Iijima (2020).
-Let us first draw some attention to what the remarkable physicist Yuhong Fan says about the essence of magnetic fields in the solar convection zone: “The question of whether, or to what extent, a strong toroidal magnetic field stored in the overshoot region at the base of the convection zone, generated by a deep seated solar dynamo process, is responsible for the formation of solar active regions remains to be more rigorously investigated.(Y.
-Fan, 2021)
-Recalling the words of Yuhong Fan that “it remains to be more thoroughly investigated” to what extent is strong toroidal magnetic field, control by means of the self-generating dynamo in the overshoot tachocline region, one must also remember the existence of the amazing 11-year variation of dark matter density (see e.g.
-V.
-D.
-Rusov et al.
-(2021)), and, for example, the holographic mechanism of quantum gravity (see S.
-W.
-Hawking (2015a, b); S.
-W.
-Hawking et al.
-(2016); V.
-D.
-Rusov et al.
-(2021)), which strongly influence the Sun, the stars, the Universe?!
-This means that, unlike the holographic mechanism of quantum gravity on the Sun, where the magnetic flux tubes in strong fields ($\sim 4.1\cdot 10^{7}~G$ in the tachocline are identical to the magnetic core of the Sun), such physics of formation, emergence and “decay” of sunspots, which is considered an integral part of the convective dynamo (see e.g.
-Y.
-Fan & F.
-Fang (2014) and Fig.
-21 in H.
-Hotta & H.
-Iijima (2020)), and, as a consequence, the formation of active regions from the uppermost convection zone to the corona (see e.g.
-F.
-Chen et al.
-(2017, 2022)), is still an unsolved problem of the 70-year evolution and the current state of solar and stellar dynamo theory (see e.g.
-P.
-Charbonneau & D.
-Sokoloff (2023) and P.
-Charbonneau (2020)), which, although good, cannot be on the Sun, and certainly not in our Universe!
-In contrast to the classical physics of rise of magnetic flux tubes from the tachocline to the photosphere, which are driven by a self-generating dynamo in the main part of the solar convective zone, we already know that,
-  • first, a strong toroidal magnetic field stored in the overshoot tachocline region is determined by the thermomagnetic Ettingshausen-Nernst effect (see Apendix A in V.
-D.
-Rusov et al.
-(2021)), which exactly (in contrast to dynamo!) coincides with $B_{tacho}\sim 10^{7}~G$.
-In this case, we indirectly showed that, using the fundamental holographic principle of quantum gravity (see Apendix C in V.
-D.
-Rusov et al.
-(2021)), the repelling toroidal magnetic field of the tachocline exactly “neutralizes” the magnetic field in the Sun’s core (see Eq.
-1 and W.
-A.
-Fowler et al.
-(1955); S.
-Couvidat et al.
-(2003)), in which the projections of the magnetic fields of the tachocline and the core have the equal value but the opposite directions.
-  • second, we underline that this process is associated with continuous transformation of toroidal magnetic field into poloidal magnetic field ($T\rightarrow P$ transformation), but not vice versa ($P\rightarrow T$).
-It means that the holographic mechanism is the main process of regeneration of the primary toroidal field in the tachocline, and the formation of the processes of these fields shears at the base of the convective zone leads to the emergence of buoyancy of toroidal magnetic tubes rising to the surface of the Sun.
-  • third, our physics of the magnetic flux tubes rise is related to the fact that if the axions, which are born in the Sun’s core, are directly converted into X-rays near the tachocline, then the axion-photon oscillations predetermine the appearance of magnetic sunspot cycles.
-This is due to the fact that the formation of sunspots and their cycles is a consequence, according to V.
-D.
-Rusov et al.
-(2021), of the anticorrelation 11-year cycles of ADM density modulation inside the Sun, which, not surprisingly, are a consequence of the 11-year ADM halo density modulation in the fundamental plane of the galactic center, which closely correlates with the density modulation of baryonic matter near the supermassive black hole.
-From here, it is easy to show (see V.
-D.
-Rusov et al.
-(2021)) in what way the anticorrelation identity between indicators of the ADM density modulation inside the Sun and the number of sunspots (or the correlation identity between indicators of modulation of solar axions (or photons of axion origin) and sunspot cycles) is realized.
-Finally, let us show that if the rise velocity of the buoyant magnetic tube, determined by the magnetic field values maximum ($B_{max}\sim 3.6\cdot 10^{7}~G$) and minimum ($B_{max}\sim 3.6\cdot 10^{6}~G$) of the magnetic field and the width of the “ring” $(a_{conv})_{max}\sim 3.7\cdot 10^{-4}~H_{p}$ or $(a_{conv})_{min}\sim 3.7\cdot 10^{-3}~H_{p}$, causes the MFT to appear on the surface in the form of sunspots, then the parameters of the universal vanBFF model will be associated with the magnetic cycles that are directly related to the observed features of the cycle of both double sunspot maxima (see Sect.
-II.2.2) and the tilt angle of Joy’s law (see Sect.
-II.2.1).
+which are almost identical to the observational data of the known works (see about 1.4 km/s in [74, 75, 76, 77]), and at the minimum of the solar cycle
+$(v_{rise})_{conv}^{min}\sim 0.14~km/s,$ (34)
+which are almost identical to the observational data of the known works (see about $\sim 0.15~km/s$ in [78]), from which we obtain the “minimum” time of increase of convective heating $(dQ/dt)_{2}$ of the flux tube from the tachocline to the photosphere (see Eq.
+(15)), which corresponds to the equation
+$(\tau_{d})_{conv}^{min}\approx 10^{6}~s\sim 10~days,$ (35)
+where $p_{ext}=(nk_{B}T)_{min}=3.56\cdot 10^{13}~erg/cm^{3}$, $T_{ext}=10^{6}~K$, $\rho_{ext}\approx 0.25~g/cm^{3}$ (see Eq. (27)).
+Ultimately, these Eq.
+(32) are the source of the first (fast in time) maximum of solar activity (e.g.
+sunspots) from the observed double maxima of 11-year cycles, while the second (two-year in time) maximum of sunspots is associated with observational data on the slope of Joy’s law (see Fig.
+8b2,c on the right).
+This raises an intriguing point regarding helioseismology.
+One of the first solutions in our section is the buoyancy of a virtually empty magnetic flux tube in strong magnetic fields of $\sim 10^{7}~G$ near the tachocline, determined by the thermomagnetic Ettingshausen-Nernst effect (see Apendix A in [14]), manifesting itself as “visible” sunspots on the solar surface.
+This is due to the fact that helioseismological inversions, which are well suited not only for studying the rotation and structure of the Sun, can sometimes be well suited due to the mediated effects of very strong magnetic fields at empty magnetic flux tubes.
+Let us repeat once again that it is the existence of empty magnetic flux tubes that allows photons of axion origin to “rush” past the tachocline to the photosphere both at high speed (see $\sim 1.4~km/s$ in [74, 75, 76, 77]) and at short times (see $\sim 1.3~days$ for Eqs.
+(31)-(32))!
+Thus, they can unequivocally confirm the presence of a magnetic field in the solar tachocline exceeding $10^{5}~G$ to, for example, $10^{7}~G$.
+On the other hand, it is the existence of “visible” sunspots in such strong fields that causes the dynamo effect to completely disappear.
+The answer, surprisingly, is quite simple.
+It means that the “measured” velocity (maximum about $\sim 1.4~km/s$ [74, 75, 76, 77] and minimum about $\sim 0.15~km/s$ in [78]) of magnetic flux tubes, which are associated with structural changes, for example, in the form of cross-sections of these tubes at the maximum ($\sim 10^{22}~Mx$) or minimum ($5\cdot 10^{21}~Mx$) of the “measured” magnetic fluxes [71, 72], can be easily determined at the maximum $\sim 10^{7}~G$ or minimum $\sim 10^{6}~G$ values of the magnetic fields (see Eqs.
+(22)-(23)).
+The modernity of this considered helio-seismological inversion remains to this day!
+In other words, the modernity of this well-thought-out helioseismological inversion remains to this day!
 II.2 Physics of primary and secondary reconnection of magnetic flux tubes in the lower layers of the tachocline
@@
 4d), and thus participates in the formation of axion origin photons (see Fig.
-3, left) and axions of photon origin (see Fig.
-3, right) from the O-loop to the photosphere, is the source of sunspots, which (due to the cooling of the Parker-Biermann effect inside the tube) rise from the tachocline to the surface of the Sun.
+2, left) and axions of photon origin (see Fig.
+2, right) from the O-loop to the photosphere, is the source of sunspots, which (due to the cooling of the Parker-Biermann effect inside the tube) rise from the tachocline to the surface of the Sun.
+Figure 6: Topological effects of secondary magnetic reconnection in the lower (left) or upper (right) layers of the magnetic tube.
+Here, the unipolar part of the $\Omega$-loop is rearranged at its base, compressing the $\Omega$-loop (blue lines) and, as a consequence, organizing a free O-loop (green lines) (a,e) to form (blue-yellow lines) the first part of the secondary magnetic reconnection (b,f).
+The yellow lines show the movement of matter leading to the loop connection “leg” (see analogous Fig.
+4 in [52]).
+If the O-loop (primary magnetic reconnection, green lines) can randomly have different directions of magnetic fields, then the secondary magnetic reconnection can generate loop “legs” in different layers, for example, in lower layers (c and d as an analog of Fig.
+4b) and upper layers (e and f as an analog of Fig.
+4 in [79]).
 On the other hand, we are interested in the essence of the very important secondary magnetic reconnection (see Fig.
-5b), which is the source of the observed double maxima of one sunspot cycle – it shifts by 2.5 years between the primary and secondary solar maxima when the magnetic tube rises, and is associated with both the observational data on the tilt angle of Joy’s law and the “disappearance” of sunspots on the solar surface.
+6b), which is the source of the observed double maxima of one sunspot cycle – it shifts by 2.5 years between the primary and secondary solar maxima when the magnetic tube rises, and is associated with both the observational data on the tilt angle of Joy’s law and the “disappearance” of sunspots on the solar surface.
 In Fig.
 6 one can see the conditions for the secondary reconnection between the O-loop (green lines) and the unipolar part of the $\Omega$-loop (blue lines) that can organize them in the lower layer (Fig.
-8, left) or in the upper layer (Fig.
-8, right), thereby showing the appearance of bipolar magnetic tubes in various versions (Fig.
-8c,d and g,h).
-Figure 8: Topological effects of magnetic reconnection in the lower (left) or upper (right) layers of the magnetic tube.
-Here, the unipolar part of the $\Omega$-loop is rebuilt on its base, compressing the $\Omega$-loop (blue lines) to form a free O-loop (green lines) (Fig.
-5a).
-The yellow lines show the movement of the substance leading to the connection of the “leg” loop (see analogous Fig.
-4 in E.
-N.
-Parker (1994)).
-If the O-loop (primary magnetic reconnection, green lines) can randomly have different directions of magnetic fields, then the secondary magnetic reconnection can generate loop “legs” in different layers, for example, in lower layers (Fig.
-8c and also Fig.
-8d as an analog of Fig.
-5b) and upper layers (Fig.
-8e and also Fig.
-8f as an analog of Fig.
-4 in S.
-Jabbari et al.
-(2016)).
-From here, since the sunspot cycle is identical to the maxima and minima of the magnetic tube number on the surface of the Sun, we are interested in the physics of the maxima and minima both during the primary reconnection for the time $(\tau_{d})_{conv}$ of the rise of the almost empty magnetic tube from the tachocline to the solar photosphere, which are associated with the features of the tilt angle of Joy’s law (see Fig.
-10c), and during the secondary reconnection for the time $(\tau_{n})_{rec}$ of the departure of the same rise of the almost empty magnetic tube from the surface of the Sun and, as a consequence, their “disappearance” on the Sun.
-II.2.1 On the ascending magnetic tube and the formation of sunspots both near the tacho-cline with $\sim 10^{7}~G$ in the absence of Joy’s law and in the middle region with $\sim 10^{5}~G$ in the presence of Joy’s law
+6, left) or in the upper layer (Fig.
+6, right), thereby showing the appearance of bipolar magnetic tubes in various versions (Fig.
+6c,d and g,h).
+From here, since the sunspot cycle is identical to the maxima and minima of the magnetic tube number on the surface of the Sun, we are interested in the physics of the maxima and minima both during the primary reconnection for the time $(\tau_{d})_{conv}$ of the rise of the almost empty magnetic tube from the tachocline to the solar photosphere (see about $\sim 1-2$ days at max in Eq.
+(31) and about $\sim 10$ days at min in Eq.
+35), and during the secondary reconnection $(\tau_{n})_{rec}$, which are associated with the features of the tilt angle of Joy’s law (see Fig.
+8c), during the rise of the almost empty magnetic tube from the surface of the Sun (see about $\sim 2$ years in Fig.
+8b2 and c) and, as a consequence, their “disappearance” on the Sun (see about <0.5<0.5 years in Fig.
+8c).
+II.2.1 Secondary reconnection of the MFT without axions, but with concern of turbulent diffusion and the tilt angle of the Joy law
 The first part of the scenario consists in discussing the physics of magnetic flux tube buoyancy at strong fields and with the participation of axion origin photons (see Fig.
-3), associating with the formation of an O-loop (see the existence of primary magnetic reconnection in Fig.
-4) inside the magnetic flux tube near the tachocline.
-For the neutral buoyancy condition ($\rho_{int}=\rho_{ext}\approx 0.2~g/cm^{3}$; see Figs.
-3 and 6) and the strong toroidal field of the magnetic tube, $B_{Sun}=4.1\cdot 10^{7}~G$, as well as the average width of the “thin” ring $a_{conv}\sim 3.7\cdot 10^{-4}~H_{p}$ (see Eq.
-(22) and Fig.
-6) between the O-loop and the walls of the magnetic tube (see Fig.
-6), it is easy to show the estimate of the time $(\tau_{d})_{conv}$ (see Eqs.
-(27)-(34)) and the velocity $(v_{rise})_{conv}$ (see Eq.
-(16) and also Eqs.
-(36)-(37) of the convective rise from the tachocline to the photosphere, which for such large magnetic fields has a significant amount of rise of the magnetic tubes at all latitudes (see Fig.
-5a and the red lines in Fig.
-9a).
-Figure 9: Scheme of turbulent reconstruction of the toroidal magnetic field in the convective zone (CZ): (a) meridional circulation (closed blue poloidal field lines), which is the source of the toroidal field in the tachocline (thin black lines) based on the holographic mechanism of quantum gravity (see Fig.
-1b), and the magnetic buoyancy of the flux tubes (red arrows) with a magnetic field of $\sim 10^{7}~G$; (b) In the CZ, through secondary turbulent reconnection of the magnetic tube near the tachocline (see Fig.
-10b2,c), effects of both macroscopic turbulent plasma diamagnetism are observed, facilitating the formation of two layers on the surface (see analogous (V.
-N.
-Krivodubskij, 2005, 2021; V.
-N.
-Krivodubskii, 1992)) with different vertical directions of horizontal field transfer (see “black” arrows), and the conditions of magnetic field transfer along $\nabla\rho$ due to small-scale magnetic pulsations (the so-called magnetic pumping, see “white” arrows), acting against buoyancy.
+1), associating with the formation of an O-loop (see the existence of primary magnetic reconnection in Fig.
+2) inside the magnetic flux tube near the tachocline.
+For the condition of neutral buoyancy ($\rho_{int}=\rho_{ext}\approx 0.2~g/cm^{3}$ [49]; see A.16 in [14], Figs.
+2 and 5) and the strong toroidal field of the magnetic tube, $B_{Sun}=4.1\cdot 10^{7}~G$, as well as the condition of the average width of the “thin” ring $a_{conv}^{min}\sim 3.7\cdot 10^{-4}~H_{p}$ (see Eq.
+(24)) between the O-shaped loop and the walls of the magnetic tube (see Fig.
+5), it is easy to recall the estimate of the time $(\tau_{d})_{conv}^{max}\sim 1.3~days$ (see Eq.
+(31)) and velocity $(v_{rise})_{conv}^{max}\sim 1.40~km/s$ (see Eq.
+(33), as well as the time $(\tau_{d})_{conv}^{min}\sim 10~days$ (see Eq.
+(35)) and velocity $(v_{rise})_{conv}^{min}\sim 0.14~km/s$ (see Eq.
+(34) of the convective rise from the tachocline to the photosphere, which for such large magnetic fields has a significant amount of rise of the magnetic tubes at all latitudes (see Fig.
+4a and the red arrows of the magnetic flux tubes in Fig.
+7a).
+Figure 7: Scheme of turbulent reconstruction of the toroidal magnetic field in the convective zone (CZ): (a) meridional circulation (closed blue poloidal field lines), which is the source of the toroidal field in the tachocline (thin black lines), and the magnetic buoyancy of the flux tubes (red arrows) with a magnetic field of $\sim 10^{7}~G$; (b) In the CZ, through secondary turbulent reconnection of the magnetic tube near the tachocline (see Fig.
+8b2,c), effects of both macroscopic turbulent plasma diamagnetism are observed, facilitating the formation of two layers on the surface (see analogous [80, 81, 82]) with different vertical directions of horizontal field transfer (see “black” arrows), and the conditions of magnetic field transfer along $\nabla\rho$ due to small-scale magnetic pulsations, acting against buoyancy.
 However, in the deep layers in the near-equatorial region, the $\nabla\rho$ effect provides upward advection, which, on the contrary, directs magnetic buoyancy, and thus, can facilitate the penetration of strong fields to the surface.
@@
 Therefore, the primary magnetic reconnection itself in the lower layers of the magnetic tubes (see Fig.
-5a), is not the final stage of the modeling.
+4a), is not the final stage of the modeling.
 The essence of a practically empty tube in strong fields of $\sim 10^{7}~G$ is related (through the primary magnetic reconnection) to the physics of the secondary turbulent reconnection of magnetic bipolar structures (see Fig.
-5b, and also Fig.
-4 in (A.
-S.
-Brun et al., 2011)).
+4b, and also Fig.
+4 in [60]).
 Here, on its basis, the bipolar part of the $\Omega$-loop is reconstructed, compressing the $\Omega$-loop (blue lines: see Fig.
-5b,c) near the tachocline and simultaneously lifting them (from 2 to 3 years) to the surface of the Sun, and, as a consequence, after some time (from several days to several months (see S.
-K.
-Solanki (2003); K.
-Petrovay & L.
-van Driel-Gesztelyi (1997)) a free O-loop appears (blue lines in Fig.
-5d), which eventually “disintegrates” or, more precisely, the O-loop (without spots) flies away from the surface of the Sun (see Fig.
-5d).
+4b,c) near the tachocline and simultaneously lifting them (from 2 to 3 years) to the surface of the Sun, and, as a consequence, after some time (from several days to several months (see [83, 84]) a free O-loop appears (blue lines in Fig.
+4d), which eventually “disintegrates” or, more precisely, the O-loop (without spots) flies away from the surface of the Sun (see Fig.
+4d).
 This raises a rather unexpected but important question: How is the secondary magnetic reconnection of flux tubes with $B_{tacho}^{Sun}=B_{z}=4.1\cdot 10^{7}~G$, leading to a real sharp decrease in the magnetic field to $B_{eq}\leqslant 2\cdot 10^{5}~G$ at $\sim 0.80R_{Sun}$ (see e.g.
 Fig.
-10b,c), related to the existence of downward turbulent diamagnetic pumping and rotational $\nabla\rho$-pumping at relatively low magnetic fields, and as a consequence, at low escape velocities of flux tubes from the solar surface?
+8b,c), related to the existence of downward turbulent diamagnetic pumping and rotational $\nabla\rho$-pumping at relatively low magnetic fields, and as a consequence, at low escape velocities of flux tubes from the solar surface?
 In this regard, we obtain an answer to this important question.
-For this purpose, we consider the second part of the scenario, based on the remarkable idea of L.
-L.
-Kichatinov (1991), which includes the generation of the magnetic field of $\leqslant 2\cdot 10^{5}~G$ near the bottom of the convective zone (see Fig.
-C.1b in V.
-D.
-Rusov et al.
-(2021) for the holographic mechanism) and transfer of the toroidal field from the deep layers at different latitudes.
-It is very important that the efficiency of the magnetic buoyancy transfer is associated with the participation of two processes: macroscopic turbulent diamagnetism (Y.
-B.
-Zel’dovich, 1957, 1956; L.
-J.
-Spitzer, 1957) and rotational $\nabla\rho$-pumping (L.
-L.
-Kichatinov, 1991).
+For this purpose, we consider the second part of the scenario, based on the remarkable idea of L.L.
+Kitchatinov [84], which includes the generation of the magnetic field of $\leqslant 2\cdot 10^{5}~G$ near the bottom of the convective zone and transfer of the toroidal field from the deep layers at different latitudes.
+It is very important that the efficiency of the magnetic buoyancy transfer is associated with the participation of two processes: macroscopic turbulent diamagnetism [85, 86, 87] and rotational $\nabla\rho$-pumping [84].
 Let us note some important properties of macroscopic turbulent diamagnetism.
-It is known that Y.
-B.
-Zel’dovich (1956, 1957) and L.
-J.
-Spitzer (1956) discovered the diamagnetism of inhomogeneously turbulent conducting liquids, in which the inhomogeneous magnetic field moves as a single whole.
+It is known that Zel’dovich [86, 85] and Spitzer [87] in 1957 discovered the diamagnetism of inhomogeneously turbulent conducting liquids, in which the inhomogeneous magnetic field moves as a single whole.
 In this case the turbulent fluid, for example, with nonuniform effective diffusivity $\eta_{T}\approx(1/3)vl$ (see Fig.
-1 in  L.
-L.
-Kitchatinov & G.
-Rüdiger (2008), and also (Y.
-B.
-Zel’dovich, 1957, 1956; L.
-J.
-Spitzer, 1957)) behaves like a diamagnetic one and carries the magnetic field with the effective velocity (H.
-Köhler, 1973)
+1 in [88], and also [85, 86, 87]) behaves like a diamagnetic one and carries the magnetic field with the effective velocity [89]
 $\vec{v}_{dia}=-\frac{1}{2}\nabla\eta_{T},$ (36)
@@
 (36) shows the meaning of turbulent magnetism: it is not paramagnetic magnetism, so magnetic fields repel from regions with relatively high turbulent intensity.
-In other words, macroscopic turbulent plasma diamagnetism and, as a consequence, the so-called macroscopic diamagnetic effect (see K.-H.
-Rädler (1968a, b)) in the physical sense is the displacement of the averaged magnetic field $B$ from regions with increased intensity of turbulent pulsations to regions with less developed turbulence (F.
-Krause & K.-H.
-Raedler, 1980; L.
-L.
-Kichatinov & G.
-Ruediger, 1992).
+In other words, macroscopic turbulent plasma diamagnetism and, as a consequence, the so-called macroscopic diamagnetic effect (see [90, 91]) in the physical sense is the displacement of the averaged magnetic field $B$ from regions with increased intensity of turbulent pulsations to regions with less developed turbulence [92, 93].
 However, there is an interesting problem of the diamagnetic process caused by inhomogeneous turbulent intensity with allowance for the total nonlinearities in the magnetic field.
@@
 For example, at strong magnetic fields of flux tubes, at $B>10^{5}~G$, e.g.
-$\geqslant 10^{7}~G$, the diamagnetic effect becomes almost negligible, in particular, strong magnetic damping of diamagnetism $\sim B^{-3}$ is obtained for super-equipartitions of fields (F.
-Chen et al., 2022) when turbulence is close to two-dimensional (Y.
-B.
-Zel’dovich, 1957).
-On the other hand, for very weak fields the diamagnetic pumping, which is predetermined by the intensity of turbulence at $B<10^{5}~G$, is a very effective process (F.
-Chen et al., 2022).
-Among the known limiting cases of weak and strong magnetic fields from the tachocline to the surface, we are interested in the addition to the portions of strong fields (via the secondary Lazaryan-Vishniac reconnection (see A.
+$\geqslant 10^{7}~G$, the diamagnetic effect becomes almost negligible, in particular, strong magnetic damping of diamagnetism $\sim B^{-3}$ is obtained for super-equipartitions of fields [94] when turbulence is close to two-dimensional [85].
+On the other hand, for very weak fields the diamagnetic pumping, which is predetermined by the intensity of turbulence at $B<10^{5}~G$, is a very effective process [94].
+Among the known limiting cases of weak and strong magnetic fields from the tachocline to the surface, we are interested, in the addition to the portions of strong fields (via the secondary Lazaryan-Vishniac reconnection (see  [95, 96, 97], and also §III and §IX in [98]) near the tachocline, which have a very slow Sweet-Parker velocity $(V_{rec})_{2}$ (see [99, 100]) for Fig.
+8b2), the part of the weak magnetic field from $0.8~R_{Sun}$ to near $0.85~R_{Sun}$ (see Fig.
+8b2,c), which constitutes the super-equipartitions of field $B\geqslant B_{eq}$ and is the source of two processes: diamagnetic pumping, and rotational $\nabla\rho$-pumping.
+This is due to the fact that the secondary magnetic reconnection of the flux tubes (see Figs.
+6b,d,f,g) leads to the real increase in the upward magnetic field to $B\approx 2\cdot 10^{5}~G$ at $\sim 0.8~R_{Sun}$, which is identical to the $1/\beta=B^{2}/8\pi p_{ext}\approx 2.9\cdot 10^{-5}$ (see e.g.
+Eq.
+(9), and Fig.
+8b2,c).
+This means that the real toroidal magnetic flux tube is a consequence of the formation of the secondary reconnection, which generates both the “anti-buoyant” effects of downward turbulent diamagnetic transport and the rotational effect of the magnetic $\nabla\rho$-pumping at polar latitudes, and the “buoyant” effects of magnetic flux tubes at equatorial latitudes as a result of the upward rotational effect of magnetic $\nabla\rho$-pumping.
+Figure 8: An alternative theory of magnetic flux tubes in strong fields by means of axion origin photons associated with primary and secondary reconnections.
+[Ref: S.K.
+Solanki (2003) [101], K.
+Petrovay & L.
+van Driel-Gesztelyi (1997) [83].]
+Figure 8: (cont.) An alternative theory of magnetic flux tubes in strong fields by means of axion origin photons associated with primary and secondary reconnections.
+[Ref: A.
 Lazarian & E.
 T.
-Vishniac (1999); A.
+Vishniac (1999) [95]; A.
 Lazarian et al.
-(2004); G.
+(2004) [96], (2020) [98]; G.
 Kowal et al.
-(2009), and also §III and §IX in (A.
-Lazarian et al., 2020)) near the tachocline, which have a very slow Sweet-Parker velocity $(V_{rec})_{2}$ (see E.
+(2009) [97]; P.
+A.
+Sweet (1958) [100]; E.
 N.
-Parker (1957); P.
-A.
-Sweet (1958)) for Fig.
-10b2), the part of the weak magnetic field from $0.8~R_{Sun}$ to near $0.85~R_{Sun}$ (see Fig.
-10b2,c), which constitutes the super-equipartitions of fields $B\geqslant B_{eq}$ and is the source of two processes: diamagnetic pumping, and rotational $\nabla\rho$-pumping.
-This is due to the fact that the secondary magnetic reconnection of the flux tubes (see Figs.
-8b,d,f,g) leads to the real increase in the upward magnetic field to $B\approx 2\cdot 10^{5}~G$ at $\sim 0.8~R_{Sun}$, which is identical to the $1/\beta=B^{2}/8\pi p_{ext}\approx 2.9\cdot 10^{-5}$ (see e.g.
-Eq.
-(9), and Fig.
-10b2,c).
-This means that the real toroidal magnetic flux tube is a consequence of the formation of the secondary reconnection, which generates both the “anti-buoyant” effects of downwardly directed turbulent diamagnetic transport and the rotational effect of the magnetic $\nabla\rho$-pumping at polar latitudes, and the “buoyant” effects of magnetic flux tubes at equatorial latitudes as a result of the upward rotational effect of magnetic $\nabla\rho$-pumping.
-Figure 10: An alternative theory of magnetic flux tubes in strong fields by means of axion origin photons associated with primary and secondary reconnections.
-Figure 11: (cont.) An alternative theory of magnetic flux tubes in strong fields by means of axion origin photons associated with primary and secondary reconnections.
-In this regard, we consider turbulence with quasi-isotropic spectral tensor (L.
-L.
-Kichatinov, 1987, 1991), which is certainly the simplest representation for inhomogeneous turbulence (see Eq.
-(2.12) in L.
-L.
-Kichatinov & G.
-Ruediger (1992)).
+Parker (1955a) [6], (1957) [99], (1974) [11], (1974a) [12], (1974b) [10], (1977) [13]; V.
+D.
+Rusov et al.
+(2021) [14]]
+In this regard, we consider turbulence with quasi-isotropic spectral tensor [102, 84], which is certainly the simplest representation for inhomogeneous turbulence (see Eq.
+(2.12) in [93]).
 As a consequence, the information on the spectral properties of turbulence (given by Eqs.
-(2.12)-(2.16) from L.
-L.
-Kichatinov & G.
-Ruediger (1992)) is sufficient to reduce the expression for the average electromotive force $\varepsilon$ (see Eq.
-(2.1) L.
-L.
-Kichatinov & G.
-Ruediger (1992)) to its traditional form, where only integrations over the wave number $k$ and frequency $\omega$ remain.
+(2.12)-(2.16) from [93]) is sufficient to reduce the expression for the average electromotive force $\varepsilon$ (see Eq.
+(2.1) [93]) to its traditional form, where only integrations over the wave number $k$ and frequency $\omega$ remain.
 After such shortening it is possible to get (see Eq.
-3.1 in L.
-L.
-Kichatinov & G.
-Ruediger (1992))
+3.1 in [93])
 $\vec{F}=(\vec{v}_{dia}+\vec{v}_{dens})\times\vec{B}$ (37)
@@
 We are interested in the problem of reconstructing a strong toroidal field of flux tubes $\sim 10^{7}~G$ (see Figs.
-9a,b with the magnetic tube profile, and Figs.
-10a,b,c with the magnetic tube in full face), transform regions of the average magnetic field $B\sim 2\cdot 10^{5}~G$ in the convective zone (see Figs.
-10b,c), and thereby allow the organization of the amazing balance between the magnetic buoyancy, turbulent diamagnetism, and the rotationally modified $\nabla\rho$-effect.
+7a,b for the magnetic tube profile, and Figs.
+8a,b,c for the magnetic tube in full face), transform the regions of the average magnetic field $B\sim 2\cdot 10^{5}~G$ in the convective zone (see Figs.
+8b,c), and thereby allow the organization of the amazing balance between the magnetic buoyancy, turbulent diamagnetism, and the rotationally modified $\nabla\rho$-effect.
 In this regard, we first consider the physics of linear magnetic diffusivity, which is known to govern two processes: macroscopic turbulent diamagnetism and rotational $\nabla\rho$-pumping.
@@
 Here we assumed that, according to Figs.
-10b2,c, the radial profile of turbulent diffusion is a smooth-convex function with a maximum, $\eta_{T}\approx 10^{11}~cm/s$, approximately in the middle of the CZ at a depth of $z=1.30\cdot 10^{5}~km$ at $\sim 0.8R_{Sun}$, which is described by the distance of $2.3\cdot 10^{4}~km$ between $0.8R_{Sun}$ and near $0.85R_{Sun}$ (see Figs.
-10b2,c) and with the depth of the bottom of the convective zone $z_{0}\approx 1.84\cdot 10^{5}~km$ (see M.
-Stix (2002); R.
-Kippenhahn et al.
-(2012); J.
-Schumacher & K.
-R.
-Sreenivasan (2020)).
+8b2,c, the radial profile of turbulent diffusion is a smooth-convex function with a maximum, $\eta_{T}\approx 10^{11}~cm/s$, approximately in the middle of the CZ at a depth of $z=1.30\cdot 10^{5}~km$ at $\sim 0.8R_{Sun}$, which is described by the distance of $2.3\cdot 10^{4}~km$ between $0.8R_{Sun}$ and near $0.85R_{Sun}$ (see Figs.
+8b2,c) and with the depth of the bottom of the convective zone $z_{0}\approx 1.84\cdot 10^{5}~km$ (see [103, 104, 105]).
 In order to consider the balance of the three velocities $v_{B}$, $v_{dia}$, and $v_{dens}$, it is necessary to apply the widely used approximation of the mixing length (see e.g.
-E.
-Böhm-Vitense (1958); P.
-Bradshaw (1974); D.
-O.
-Gough (1977); D.
-Gough (1977); A.
-J.
-Barker et al.
-(2014); A.
-Brandenburg (2016)), which, according to L.
-L.
-Kichatinov (1991), fully satisfies this goal.
+[58, 106, 107, 108]), which, according to [84], fully satisfies this goal.
 This approximation will be understood as the replacement of nonlinear terms along with time derivatives in the equations for fluctuating fields by means of $\tau$-relaxation terms, i.e.
-instead of equations (3.1) and (3.9) from L.
-L.
-Kichatinov (1991), we now have the equation of the radial speed of the toroidal field in the convection zone
+instead of equations (3.1) and (3.9) from [84], we now have the equation of the radial speed of the toroidal field in the convection zone
 […]
@@
 $v_{dens}^{light}\sim 3\times\eta_{T}\times g_{0.8R_{Sun}}/(\gamma-1)c_{p}T_{0.8R_{Sun}}\approx 4.1\cdot 10^{1}~cm/s,$ (43)
-where $\tau\approx l/(\langle u^{2}\rangle^{\circ})^{1/2}$ is a typical lifetime of a convective eddy; $l$ is the mixing length; $\langle u^{2}\rangle^{\circ}$ is the mean intensity of fluctuating velocities for original turbulence; $\theta$ is the latitude (scalar); $e_{r}\nabla\rho/\rho=-e_{r}g/[(\gamma-1)c_{p}T]$ , where $e_{r}$ is the radial unit vector; $T\cong 1.352\cdot 10^{6}~K$ is the temperature at $0.8R_{Sun}$ (J.
-Bahcall & M.
-Pinsonneault, 1992); $g=g_{0}(R_{Sun}/r)$ is the gravity, where $g_{0}=2.74\cdot 10^{4}~cm/s^{2}$ is the surface value of solar gravity, and $g_{0.8R_{Sun}}=4.22\cdot 10^{4}cm/s^{2}$ at $0.8R_{Sun}$; $c_{p}=3.4\cdot 10^{8}~cm^{2}s^{-2}K^{-1}$ (fully ionized hydrogen) is the specific heat at constant pressure; $\gamma=5/3$ is the ratio of specific heats $c_{p}/c_{V}$; $3\kappa=\tau\langle u^{2}\rangle^{\circ}$, where $\kappa\equiv\eta_{T}\approx 10^{11}~cm^{2}/s$ is turbulent diffusivity coefficient, and the mixing length relation $\langle u^{2}\rangle^{\circ}=-\nabla\Delta Tl^{2}g/(4T)$, where the $\nabla\Delta T$ is superadiabatic temperature gradient; $v_{p}=(1/6)\tau\langle u^{2}\rangle^{\circ}(\nabla\rho/\rho)$ is the velocity of the magnetic field transfer caused by the density gradient (see Eq.
-(36) in S.
-I.
-Vainshtein & L.
-L.
-Kichatinov (1983)); $\hat{\Omega}=Co=2\tau\Omega$ is the Coriolis number (reciprocal of the Rossby number), where $\Omega$ is the rotation speed, $\tau$ is the turnover time, and the functions
+where $\tau\approx l/(\langle u^{2}\rangle^{\circ})^{1/2}$ is a typical lifetime of a convective eddy; $l$ is the mixing length; $\langle u^{2}\rangle^{\circ}$ is the mean intensity of fluctuating velocities for original turbulence; $\theta$ is the latitude (scalar); $e_{r}\nabla\rho/\rho=-e_{r}g/[(\gamma-1)c_{p}T]$ , where $e_{r}$ is the radial unit vector; $T\cong 1.352\cdot 10^{6}~K$ is the temperature at $0.8R_{Sun}$ [49]; $g=g_{0}(R_{Sun}/r)$ is the gravity, where $g_{0}=2.74\cdot 10^{4}~cm/s^{2}$ is the surface value of solar gravity, and $g_{0.8R_{Sun}}=4.22\cdot 10^{4}cm/s^{2}$ at $0.8R_{Sun}$; $c_{p}=3.4\cdot 10^{8}~cm^{2}s^{-2}K^{-1}$ (fully ionized hydrogen) is the specific heat at constant pressure; $\gamma=5/3$ is the ratio of specific heats $c_{p}/c_{V}$; $3\kappa=\tau\langle u^{2}\rangle^{\circ}$, where $\kappa\equiv\eta_{T}\approx 10^{11}~cm^{2}/s$ is turbulent diffusivity coefficient, and the mixing length relation $\langle u^{2}\rangle^{\circ}=-\nabla\Delta Tl^{2}g/(4T)$, where the $\nabla\Delta T$ is superadiabatic temperature gradient; $v_{p}=(1/6)\tau\langle u^{2}\rangle^{\circ}(\nabla\rho/\rho)$ is the velocity of the magnetic field transfer caused by the density gradient (see Eq.
+(36) in [109]); $\hat{\Omega}=Co=2\tau\Omega$ is the Coriolis number (reciprocal of the Rossby number), where $\Omega$ is the rotation speed, $\tau$ is the turnover time, and the functions
 $\phi_{n}(\hat{\Omega})=(1/8)I_{n}(\Omega,k,\omega)$ (44)
-(see $I_{1}$ and $I_{2}$ in Eqs. (3.12) and (3.21) in L. L. Kichatinov (1991)) are
+(see $I_{1}$ and $I_{2}$ in Eqs. (3.12) and (3.21) in [84]) are
 $\phi_{1}(\hat{\Omega})=\frac{1}{4\hat{\Omega}^{2}}\left[-3+\frac{\hat{\Omega}^{2}+3}{\hat{\Omega}}\arctan\hat{\Omega}\right$], (45)
 $\phi_{2}(\hat{\Omega})=\frac{1}{8\hat{\Omega}^{2}}\left[1+\frac{\hat{\Omega}^{2}-1}{\hat{\Omega}}\arctan\hat{\Omega}\right$] (46)
-(see also analogous Eq. (19) and Fig. 2 in L. L. Kitchatinov & A. A. Nepomnyashchikh (2016)), which describe the rotational effect on turbulent convection.
+(see also analogous Eq. (19) and Fig. 2 in [110]), which describe the rotational effect on turbulent convection.
+From here, we obtain an the estimate of the Coriolis number for solar convection in the deep layer, which should be $\hat{\Omega}\approx 20$.
+By virtue of the equations (44)-(46) and the function $\phi_{n}$ (see Eq.
+(44)) for the value $\hat{\Omega}\approx 20$ (see from $0.8R_{Sun}$ in Fig.
+8b2,c), the following values are adopted:
+$\phi_{1}\cong 0.0171,~~~\phi_{2}\cong 0.0098,$ (47)
+at which the radial velocity $v_{dens}^{light}$ in Eq.
+(40) of toroidal field transport changes the sign at the latitude $\theta^{*}=\arccos\sqrt{\varphi_{2}/\varphi_{1}}\cong 41^{\circ}$, being negative (downward) for $\theta>\theta^{*}$ and positive (upward) for $\theta<\theta^{*}$.
+Using Eq.
+(42), we find that the value of the radial velocity of toroidal field transport in the convective zone $v_{dens}^{light}$ (see Eqs.
+(40)-(42)) near low latitudes (e.g.
+$\theta^{*}=\arccos(0.985)\cong 10^{\circ}$; see also Fig.
+7b) is almost identical to the value of the velocity (43), which was previously calculated using the toroidal field $B\approx 2\cdot 10^{5}~G$ at $0.8R_{Sun}$ (see Eq.
+(9) and Fig.
+8b,c) in the $\Re_{dens}(\beta,\varphi)$ (see Eq.
+(3.13) in [93]).
 Secondly, we get diamagnetic pumping rate
@@
 Thirdly, in addition to the radial profile of turbulent diffusion $\eta_{T}\approx 10^{11}~cm^{2}/s$, which is related to the diamagnetic pumping velocity $v_{dia}\approx 1.7\cdot 10^{2}~cm/s$ and the distance $2.3\cdot 10^{4}~km$ between $0.8R_{Sun}$ and near $0.85R_{Sun}$ (see the “green” line of the turbulent-toroidal magnetic field in Fig.
-10b,c), we need to calculate the magnetic buoyancy velocity $v_{B}$, which is associated with a super-uniformly distributed field strength $B\sim 2\cdot 10^{5}~G$ (see Fig.
-9b and Fig.
-10b,c).
-It should be remembered that for buoyant flux tubes with super-equipartition field strength $B\sim 2\cdot 10^{5}~G\geqslant(H_{p}/a)^{1/2}B_{eq}\approx 3.16B_{eq}$ with $a\approx 0.1H_{p}$, where, as Yuhong Fan wrote (see Section 7 in (Y.
-Fan, 2021)), $B_{eq}\approx 6.1\cdot 10^{4}~G$ is in equipartition with the kinetic energy density of the convective downflows, the magnetic buoyancy of the tubes dominates the hydrodynamic force from the convective downflows and the flux tubes can rise unaffected by convection”, but significantly affects the turbulent diamagnetic pumping!
+7b and Fig.
+8b,c), we need to calculate the magnetic buoyancy velocity $v_{B}$, which is associated with a super-uniformly distributed field strength $B\sim 2\cdot 10^{5}~G$ (see Fig.
+7b and Fig.
+8b,c).
+It should be remembered that for buoyant flux tubes with super-equipartition field strength $B\sim 2\cdot 10^{5}~G\geqslant(H_{p}/a)^{1/2}B_{eq}\approx 3.16B_{eq}$ with $a\approx 0.1H_{p}$, where, as Yuhong Fan wrote (see Section 7 in [70]), $B_{eq}\approx 6.1\cdot 10^{4}~G$ is in equipartition with the kinetic energy density of the convective downflows, the magnetic buoyancy of the tubes dominates the hydrodynamic force from the convective downflows and the flux tubes can rise unaffected by convection”, but significantly affects the turbulent diamagnetic pumping!
 For this, using our so-called universal van Ballegooijen-Fan-Fisher model with equations (8)-(12) with the values of the magnetic flux tube with $\sim 2\cdot 10^{5}~G$, the pressure of $1.665\cdot 10^{13}~erg/cm^{3}$ at $0.8R_{Sun}$ and the rate of radiative heating $(dQ/dt)_{1}\approx 29.7~erg\cdot cm^{-3}\cdot s^{-1}$, we obtained the magnetic buoyancy velocity $v_{B}\approx 1.7\cdot 10^{2}~cm/s$, which, surprisingly, exactly coincides with the diamagnetic pumping velocity $v_{dia}\approx 1.7\cdot 10^{2}~cm/s$.
@@
 One of the main tasks was related to the solution of the maximum near-equatorial time $(\tau_{n})_{max}$ of the rise of the magnetic flux tube in the region between $0.8R_{Sun}$ and near $0.85R_{Sun}$ (see Fig.
-9b and Fig.
-10b,c), which is associated with the distance of turbulent diffusion $l_{diffus}=2.3\cdot 10^{4}~km$ and the speed of rotational $\nabla\rho$-pumping $v_{dens}\approx 4.1\cdot 10~cm/s$:
+7b and Fig.
+8b,c), which is associated with the distance of turbulent diffusion $l_{diffus}=2.3\cdot 10^{4}~km$ and the speed of rotational $\nabla\rho$-pumping $v_{dens}\approx 4.1\cdot 10~cm/s$:
 $(\tau_{n})_{max}$ $=\frac{l_{diffus}}{v_{dens}}\approx 0.56\cdot 10^{8}~s\sim 2~years,$
@@
 When the toroidal fields of magnetic tubes rising from the tachocline to the solar surface, but without the participation of the turbulent diffusion distance (see Eq.
-39), have strong fields of the order of $2\cdot 10^{5}~G$ (see Fig.
-9a and Fig.
-10a), then they
+36), have strong fields of the order of $2\cdot 10^{5}~G$ (see Fig.
+7a and Fig.
+8a), then they
 $v_{B}(>2\cdot 10^{5}~G)\gg v_{dia}+v_{dens},$ (53)
@@
 31 and $(\tau_{d})_{conv}^{min}$ in Eq.
-34).
+35).
 The most interesting thing is that, in addition to the sunspots after the rise of the magnetic tube with $B\sim 10^{7}~G$ at all latitudes (see Fig.
-9a and Fig.
-10a) and the sunspots after the rise of the magnetic tube with $B\leqslant 2\cdot 10^{5}~G$ only at low latitudes (see Fig.
-9b and Fig.
-10c), which are the source of the double maxima of the 11-year solar cycles (see the full experimental evidence in Section II.2.2), we are interested in the problem or, more precisely, the physics of the tilt angle of Joy’s law (V.
-G.
-Ivanov, 2012; M.
-Dasi-Espuig et al., 2010) and how it is related to the magnetic $\nabla\rho$-pumping?
-First, about the Coriolis force.
-According to Käpylä, P.
-J.
-et al.
-(2014), the rotational effect on the flow can be measured with the local Coriolis number $\hat{\Omega}\equiv Co=2\tau\Omega$, especially if $\tau$ is estimated on the basis of the theory of mixing length, which predicts values of $\hat{\Omega}$ reaching more than 10 in the deep layers (see e.g.
-M.
-Ossendrijver (2003); A.
-Brandenburg & K.
-Subramanian (2005); P.
-Käpylä (2011); P.
-Käpylä et al.
-(2011)).
-However, on the other hand, according to Käpylä, P.
-J.
-et al.
-(2014), the question arises of whether or not there are differential rotation bistabilities in which antisolar (stellar-like) and/or solar-like rotation depends on the initial state or launch history, or, in addition, the question of whether there are such deep layers of solar and stellar convection zones or not is still open.
-In contrast to the work of Käpylä, P.
-J.
-et al.
-(2014), already in 2021, astrophysicist Krivodubskii (V.
-N.
-Krivodubskij, 2005, 2021) recalled (see page 72 in (V.
-N.
-Krivodubskij, 2005)) that exactly “…Modern helioseismologocal inversions (J.
-Schou et al., 1998; R.
-Howe et al., 2000; H.
-M.
-Antia, 2002; S.
-Basu & H.
-M.
-Antia, 2003; S.
-Basu & H.
-Antia, 2008) show that the differential rotation rate, $\Omega(r,\theta)$, in the SCZ can be divided into domains of fast (equatorial) and slow (polar) rotation.
-And, what is more important, the radial angular-velocity gradient $\partial\Omega/\partial r$ has opposite signs in these domains.
-The angular velocity increases along the radius, $\partial\Omega/\partial r>0$, at the near-equator region, and, in contrast, it decreases along the radius, $\partial\Omega/\partial r<0$, at the polar regions.” The modernity of this considered helio-seismological inversion remains to this day!
-At the same time, we will show a little later that this rotating magnetic $\nabla\rho$-pumping due to the second part (slow) secondary reconnection of the $\Omega$-loop (see $(V_{rec})_{2}$ in Figs.
-10b2,c) (through the primary part (fast) secondary reconnection of the $\Omega$-loop (see $(V_{rec})_{1}$ in Fig.
-10b1, also in Fig.5a-d, Fig.
-8b,d, and also Fig.
-8c)) both in the polar region of slow rotation, where the angular velocity increases inward of the Sun ($\partial\Omega/\partial r<0$), in the equatorial region of fast rotation, where the angular velocity mainly decreases with depth ($\partial\Omega/\partial r>0$).
-Moreover, the direction of the magnetic $\nabla\rho$-pumping (up or down, see Fig 10b2,c) is susceptible to the sign of the factor in Eqs.
-(52)-(53), which depends on the polar angle (colatitude) and the behavior of the Coriolis number functions in the convection zone.
-Below we show that the estimate of the Coriolis number for solar convection in the deep layer should be $\hat{\Omega}\approx 20$.
-Due to Eqs.
-(47)-(49) and the function $\phi_{n}$ for the quantity $\hat{\Omega}\approx 20$, the following values are adopted:
-$\phi_{1}\cong 0.0171,~~~\phi_{2}\cong 0.0098,$ (56)
-at which the radial velocity $v_{dens}^{light}$ in Eq.
-(43) of toroidal field transport changes the sign at the latitude $\theta^{*}=\arccos\sqrt{\varphi_{2}/\varphi_{1}}\cong 41^{\circ}$, being negative (downward) for $\theta>\theta^{*}$ and positive (upward) for $\theta<\theta^{*}$.
-Using (45), we find that the value of the radial velocity in the convective zone $v_{dens}^{light}$ (see Eqs.
-(43)-(45)) near low latitudes (e.g.
-$\theta^{*}=\arccos(0.985)\cong 10^{\circ}$; see also Fig.
-9b) almost completely coincides with the value of the speed (46), which was previously calculated using the toroidal field $B\approx 2\cdot 10^{5}~G$ at $0.8R_{Sun}$ (see Eq.
-(9) and Fig.
-10b,c) in the $\Re_{dens}(\beta,\varphi)$ (see Eq.
-(3.13) in (L.
-L.
-Kichatinov & G.
-Ruediger, 1992)).
+7a and Fig.
+8a) and the sunspots after the rise of the magnetic tube with $B\leqslant 2\cdot 10^{5}~G$ only at low latitudes (see Fig.
+7b and Fig.
+8c), which are the source of the double maxima of the 11-year solar cycles (see the full experimental evidence in Section II.2.2), we are interested in the problem or, more precisely, the physics of the tilt angle of Joy’s law [111, 112] and how it is related to the magnetic $\nabla\rho$-pumping?
 From here, this raises an unexpected question: In what way, in addition to the visible magnetic tube at strong fields at all latitudes (see Fig.
-9a and Fig.
-10a), the “invisible” magnetic tube at high latitudes (see Fig.
-10b2) and the “visible” magnetic tube at low latitudes (see Fig.
-10c), depending on the time $(\tau_{rec})_{2}\sim 2.5~years$ of reconnection of the horizontal field in the tachocline (which is equal to the sum of $(\tau_{n})_{max}\approx 2~years$ and the lifetime of sunspots ($\leqslant 0.5~years$) from several days to several months (see S.
-D’Silva & A.
-R.
-Choudhuri (1993); Y.
-Fan et al.
-(1993, 1994)) surprisingly, practically does not depend on the direction in the radial plane down or up?
+7a and Fig.
+8a), the “invisible” magnetic tube at high latitudes (see Fig.
+8b2) and the “visible” magnetic tube at low latitudes (see Fig.
+8c), depending on the time $(\tau_{rec})_{2}\sim 2.5~years$ of reconnection of the horizontal field in the tachocline (which is equal to the sum of $(\tau_{n})_{max}\approx 2~years$ and the lifetime of sunspots ($\leqslant 0.5~years$) from several days to several months (see [113, 69, 114]) surprisingly, practically does not depend on the direction in the radial plane down or up?
 The answer, surprisingly, is not very simple.
 We know that the vertical compression (see Fig.
-9a and Fig.
-10b1) of the descending nearly empty anchored magnetic tube in strong fields of $\sim 10^{7}~G$ (and via convective heating $(dQ/dt)_{2}$ is associated with a horizontal “stretching” or expansion of this magnetic tube near the tachocline, where the counter-movement of magnetic fields leads to the emergence of a collisionless interaction with fast (opposite) velocities $(V_{rec})_{1}$ (and times $(\tau_{rec})_{1}$ (<<10 days)) of the first part of the fast secondary (turbulent) reconnection (see Fig.
-10b1).
+7a and Fig.
+8b1) of the descending nearly empty anchored magnetic tube in strong fields of $\sim 10^{7}~G$ (and via convective heating $(dQ/dt)_{2}$ is associated with a horizontal “stretching” or expansion of this magnetic tube near the tachocline, where the counter-movement of magnetic fields leads to the emergence of a collisionless interaction with fast (opposite) velocities $(V_{rec})_{1}$ (and times $(\tau_{rec})_{1}$ (<<10 days)) of the first part of the fast secondary (turbulent) reconnection (see Fig.
+8b1).
 When the vertical compression leads to horizontal stronger “stretching” of the magnetic tube in weak fields of $\geqslant 2\cdot 10^{5}~G$ (and through radiative heating $(dQ/dt)_{1}$, then the counter motion of magnetic fields (see $(V_{rec})_{1}$ for Fig.
-10b1) leads to the emergence of the second part of the secondary (turbulent) reconnection with a very slow Sweet-Parker speed $(V_{rec})_{2}$ (see Fig.
-10b2; also A.
-Lazarian & E.
-T.
-Vishniac (1999); A.
-Lazarian et al.
-(2004); G.
-Kowal et al.
-(2009); A.
-Lazarian et al.
-(2020); E.
-N.
-Parker (1957); P.
-A.
-Sweet (1958)).
+8b1) leads to the emergence of the second part of the secondary (turbulent) reconnection with a very slow Sweet-Parker speed $(V_{rec})_{2}$ (see Fig.
+8b2; also [95, 96, 97, 98, 99, 100]).
 This means that the emerging weak fields of $\geqslant 2\cdot 10^{5}~G$ at distances from $0.80R_{Sun}$ to near $0.85R_{Sun}$ are a remarkable source of rotating magnetic $\nabla\rho$-pumping with the rise time of the magnetic tube $(\tau_{n})_{max}\sim 2.0~years$ from the tachocline to the solar surface (see Fig.
-10c), and the lifetime of sunspots ($\leqslant 0.5~years$), and consequently, the time of the magnetic tube evaporation $(\tau_{rec})_{2}\leqslant 2.5~years$ and their disappearance on the solar surface.
+8c), and the lifetime of sunspots ($\leqslant 0.5~years$), and consequently, the time of the magnetic tube evaporation $(\tau_{rec})_{2}\leqslant 2.5~years$ and their disappearance on the solar surface.
 Since the study of the tendency of the tilt angle of Joy’s law is very important for understanding the evolution of the solar magnetic field, then unlike the heavy calculations of theoretical estimates of the averaged angle tilt of sunspots with increasing latitude, we can summarize our results (see Eqs.
-(5)-(9) in S.
-D’Silva & A.
-R.
-Choudhuri (1993)), which can simultaneously be expressed in the form of simple and understandable physics of Joy’s law.
-It is known that, according to S.
-D’Silva & A.
-R.
-Choudhuri (1993) and Y.
-Fan et al.
-(1993), the Coriolis force is proportional to the magnetic buoyancy velocity $v_{B}$, which can be estimated taking into account the balance between the buoyancy force (the term on the left) and the drag force (left term) and the drag force (right term):
+(5)-(9) in [113]), which can simultaneously be expressed in the form of simple and understandable physics of Joy’s law.
+It is known that, according to [113] and [69], the Coriolis force is proportional to the magnetic buoyancy velocity $v_{B}$, which can be estimated taking into account the balance between the buoyancy force (the term on the left) and the drag force (left term) and the drag force (right term):
 $\frac{B^{2}}{8\pi H_{p}}=C_{D}\frac{\rho_{ext}v^{2}_{B}}{\pi a},$ (54)
@@
 Taking into account Eqs.
-(18) and (19) by Y.
-Fan et al.
-(1994) and using the theoretical values of the tilt angle $\alpha$, the latitude angle $\theta$ and the radial parameter $\xi=r/R_{Sun}\sim 0.80$, one can easily obtain a simplified equation in the following form:
+(18) and (19) by [114] and using the theoretical values of the tilt angle $\alpha$, the latitude angle $\theta$ and the radial parameter $\xi=r/R_{Sun}\sim 0.80$, one can easily obtain a simplified equation in the following form:
 $\sin{(tilt)}\propto\frac{\sqrt{a(\xi,\theta)}}{B}\sin{(latitude)}.$ (55)
 This means that looking at equation (55), we understand that when magnetic tubes contain relatively large radii $a$ (due to radiative heat transfer $(dQ/dt)_{1}$ (see Fig.
-10b2,c) and low magnetic fields ($\leqslant 10^{5}~G$), then the tilt angle of Joy’s law are born from here, and when magnetic tubes have practically zero radii $a$ (due to convective heat transfer $(dQ/dt)_{2}$ (see Fig.
-10b2,c) and strong high magnetic fields ($10^{5}~G$), then the tilt angle of Joy’s law disappears!
+8b2,c) and low magnetic fields ($\leqslant 10^{5}~G$), then the tilt angle of Joy’s law are born from here, and when magnetic tubes have practically zero radii $a$ (due to convective heat transfer $(dQ/dt)_{2}$ (see Fig.
+8b2,c) and strong high magnetic fields ($10^{5}~G$), then the tilt angle of Joy’s law disappears!
 It can be said in a completely different way.
 If radiative heat transfer $(dQ/dt)_{1}$ (see Fig.
-10b2,c) or the presence of magnetic $\nabla\rho$-pumping is used in magnetic tubes, then, as usual, the tilt angles of Joy’s law appear in sunspots, when convective heat transfer $(dQ/dt)_{2}$ or the disappearance of magnetic $\nabla\rho$-pumping (see Fig.
-10b2,c) is used in magnetic tubes, then, strangely enough, the tilt angles of Joy’s law disappear in sunspots!
-In this regard, we are interested in an intriguing question: Where and how do sunspots appear, in which, according to M.
-N.
-Gnevyshev (1963), the maxima of the double peak in the sunspot cycle appear?
+8b2,c) or the presence of magnetic $\nabla\rho$-pumping is used in magnetic tubes, then, as usual, the tilt angles of Joy’s law appear in sunspots, when convective heat transfer $(dQ/dt)_{2}$ or the disappearance of magnetic $\nabla\rho$-pumping (see Fig.
+8b2,c) is used in magnetic tubes, then, strangely enough, the tilt angles of Joy’s law disappear in sunspots!
+In this regard, we are interested in an intriguing question: Where and how do sunspots appear, in which, according to [115], the maxima of the double peak in the sunspot cycle appear?
 And how does the first maximum – the maximum number of spots – physically differ from the second – the maximum of large spots?
@@
 Secondly, according to equation (54), when the flux tube has a large radius and a low magnetic field, then this magnetic tube, surprisingly, is the source of the formation of the tilt angle (via the Coriolis force, and not necessarily on the surface of the photosphere (see Fig.
-10c)) and a large area of the spot on the surface of the Sun.
-Conversely, when the flux tube has a small radius and a high magnetic field, then the magnetic tube, on the one hand, is not connected to the source of the formation of the tilt angle, and on the other hand, surprisingly, has a small spot area on the surface of the Sun.
-II.2.2 New physics of observed double maxima of sunspot cycles and solutions of time interval or Gnevyshev gap
-The phenomenon of a double peak in the sunspot cycle was discovered by M.
-N.
-Gnevyshev (1963) during a study of coronal emission and areas of active formations in the 11-year cycle 19 (M.
-N.
-Gnevyshev, 1963).
+8c)) and a large area of the spot on the surface of the Sun.
+Conversely, when the flux tube has a small radius and a high magnetic field, then the magnetic tube, on the one hand, is not connected to the source of the formation of the tilt angle of Joy’s law (through the Coriolis force), and on the other hand, surprisingly, has a small spot area on the surface of the Sun.
+This raises a second intriguing point.
+One of the first solutions in our section is secondary reconnection of the MFT, which involves a turbulent diffusivity of $10^{11}~cm/s$ and a tilt angle of Joy’s law between magnetic fields ranging from $2\cdot 10^{5}~G$ (at $0.8R_{Sun}$) to $6\cdot 10^{4}~G$ (at $0.85R_{Sun}$).
+Moreover, due to the relationship between turbulent diffusivity and the tilt angle of Joy’s law, the direction of magnetic $\nabla\rho$-pumping (upward or downward (see Fig.
+8b2,c)) is sensitive to the sign of the multiplier in equations (45)-(45), which depends on the polar angle (co-latitude) and the behavior of the Coriolis number functions in the convective zone.
+It is crucial that the existence of secondary MFT reconnection generates a secondary maximum in the 11-year sunspot cycle.
+This is because when the primary reconnection, in which the buoyancy of a practically empty magnetic flux tube at $\sim 10^{7}~G$ rises from the tachocline to the photosphere for $\sim 2$ to $10$ days, transitions to secondary reconnection, the latter is associated with a buoyancy of practically MFT from $\sim 6\cdot 10^{4}~G$ from $0.85R_{Sun}$ (see Fig.
+8c)) to the photosphere.
+From this we understand that since the lifetime of $\sim 2$ years for the rotating magnetic $\nabla\rho$-pumping from $2\cdot 10^{5}~G$ at $0.8R_{Sun}$ (see Fig.
+8c) to $6\cdot 10^{4}~G$ is related to the lifetime of the order of $\sim 10$ days for the buoyancy of almost MFT from $\sim 6\cdot 10^{4}~G$ to the photosphere, this means that the secondary reconnection of MFT from the tachocline to the photosphere is also equal to $\sim 2$ years.
+In other words, the connection between primary and secondary reconnections (or first and second maxima) is the connection of MFTs between lifetimes of the order of $\sim 2$ to $10$ days and lifetimes of the order of $\sim 2$ years, which is preserved by 11-year sunspot cycles without any dynamo.
+II.2.2 Double maxima of 11-year sunspot cycles and the Gnevyshev gap
+The phenomenon of a double peak in the sunspot cycle was discovered by Mstislav Gnevyshev in 1963 [115] during a study of coronal emission and areas of active formations in the 11-year cycle 19 [115].
 First, Gnevyshev analyzed the evolution of the annual average intensity of the green coronal spectral line $\lambda 530.3$ (which serves as a direct indicator of magnetic activity and coronal heating) in five-degree heliolatitude intervals during the cycle 19 and discovered two maxima in the intensity of the corona.
-According to data in (A.
-Antalova & M.
-N.
-Gnevyshev, 1965; M.
-N.
-Gnevyshev, 1967, 1977) the first maximum of the area of sunspot groups coincides with the main maximum of the 11-year cycle for Wolf numbers, while the second maximum is associated with an increase in the number of large sunspots (the so-called maximum of sunspot intensities) (Y.
-I.
-Vitinskij et al., 1986; M.
-Kopecký & G.
-V.
-Kuklin, 1969).
-In other words, the first maximum is the maximum in the number of sunspots, and the second is the maximum of large sunspots (M.
-Kopecký & G.
-V.
-Kuklin, 1969).
+According to data in [116, 117, 118], the first maximum of the area of sunspot groups coincides with the main maximum of the 11-year cycle for Wolf numbers, while the second maximum is associated with an increase in the number of large sunspots (the so-called maximum of sunspot intensities) [119, 120].
+In other words, the first maximum is the maximum in the number of sunspots, and the second is the maximum of large sunspots [120].
 The more large spots in the cycle, the more distinct the two peaks.
 These features were later confirmed and for the cycle 20 (see Fig.
-12), including for the solar northern and southern hemispheres separately (M.
-N.
-Gnevyshev, 1963).
+9), including for the solar northern and southern hemispheres separately [115].
+In contrast to the “new solar dynamo era” (see e.g.
+[80, 70, 121, 122, 1, 123, 2, 124]), we used our alternative theory of magnetic flux tubes at strong fields via photons of axion origin.
+The essence of double maximum 11-year sunspot cycles, the source of which is the generation of a magnetic field near the bottom of the convective zone and the subsequent rise of the field from the deep layers to the surface in the photosphere, manifests itself at strong fields only (see Fig.
+8a, Fig.
+8b2,c).
+This means that the two maxima are related as follows: the first maximum (the primary reconnection of the empty MFT with $10^{7}~G$ at $0.72R_{Sun}$; see Figs.
+7a, 8a), which generates a very “fast” velocity to the solar surface (on the order of several days; see Eq.
+(31) $\sim 1-2~days$) is transformed into a secondary maximum (the secondary reconnection of the empty MFT with $10^{6}~G$ at $0.85R_{Sun}$; see Figs.
+7a, 8a), which then generates a “very slow” velocity to the solar surface (on the order of two years; see Eq.
+(52) of $(\tau_{n})_{max}\sim 2.0~years$ for Fig.
+8b2,c).
+In other words, the “very slow” second maximum depends only on two interrelated causes: the very slow secondary reconnection (see Fig.
+8b2,c) of the magnetic flux tube, which depends both at a strong magnetic field of $2\cdot 10^{5}~G$, having two sections between $0.7R_{Sun}$ and near $0.80R_{Sun}$, and between $0.85R_{Sun}$ and the photosphere (see Fig.
+8b2,c), and, surprisingly, at a weak magnetic field of $2\cdot 10^{5}~G$, where the so-called effects of turbulent diamagnetic transport and rotating $\nabla\rho$-pumping appear for the first time, which provide both the process of blocking magnetic buoyancy at high latitudes (see Fig.
+8b2), and at near-equatorial latitudes predetermines the process of dominance of the Coriolis force with rotating $\nabla\rho$-pumping with $2\cdot 10^{5}~G$ at a distance $2.3\cdot 10^{4}~km$ between $0.80R_{Sun}$ and near $0.85R_{Sun}$ (see Fig.
+8c), which ultimately generates a “shift” of the arcuate ascending loop in the form of the tilt angle of Joy’s law on the solar surface (see Fig.
+8c).
 Figure 9: Variation in the number of sunspots for six cycles from 19 to 24 associated with double maxima of 11-year solar cycles: (a) solid red lines of sunspot cycles – smoothed values (https://sidc.be/SILSO) Royal Observatory of Belgium 2023, August 23.
 (b) theoretical assessment of double maxima of 11-year solar cycles: solid line – “smoothed” blue values of sunspot cycles, in which the sum of the curves with the first maximum in the form of a “red” value and the second maximum in the form of a “green” value almost exactly coincides with a solid “blue” line of the corresponding cycle.
-In contrast to the “new solar dynamo era” (see e.g.
-V.
-N.
-Krivodubskij (2005); Y.
-Fan (2021); Z.
-Zhang & J.
-Jiang (2022); Z.
-Zhang et al.
-(2023); P.
-Charbonneau & D.
-Sokoloff (2023); G.
-M.
-Vasil et al.
-(2024)), we used our alternative theory of magnetic flux tubes at strong fields via photons of axion origin.
-The essence of double maximum 11-year sunspot cycles, the source of which is the generation of a magnetic field near the bottom of the convective zone and the subsequent rise of the field from the deep layers to the surface in the photosphere, appear only at strong fields (see Fig.
-9a), which generate a very “fast” (on the order of several days; see Eq.
-(31) $\sim 1-2~days$) first maximum (see Fig.
-9a), and, as a consequence, the obligatory addition of a section at weak fields (see Fig.
-10b2,c), at which it is the first maximum that is transformed into a very “slow” (on the order of two years!
-See $(\tau_{n})_{max}\sim 2.0~years$ second maximum (see Fig.
-10b2,c)!
-In other words, the very “slow” second maximum depends only on two interrelated causes: the very slow secondary reconnection (see Fig.
-10b2,c) of the magnetic flux tube, which depends both at a strong magnetic field of $2\cdot 10^{5}~G$, having two sections between $0.70R_{Sun}$ and near $0.80R_{Sun}$ and between $0.85R_{Sun}$ and the photosphere (see Fig.
-10b2,c), and, surprisingly, at a weak magnetic field of $2\cdot 10^{5}~G$, where the so-called effects of turbulent diamagnetic transport and rotating $\nabla\rho$-pumping appear for the first time, which provide both the process of blocking magnetic buoyancy at high latitudes (see Fig.
-10b2), and at near-equatorial latitudes predetermines the process of dominance of the Coriolis force with rotating $\nabla\rho$-pumping with $\geqslant 2\cdot 10^{5}~G$ at a distance $2.3\cdot 10^{4}~km$ between $0.80R_{Sun}$ and near $0.85R_{Sun}$ (see Fig.
-10c), which ultimately generates a “shift” of the arcuate ascending loop in the form of the tilt angle of Joy’s law on the solar surface (see Fig.
-10c).
-In this case, the key role in the proposed mechanism of double maxima is played by two “waves” of toroidal fields, in which the first wave, without the participation of secondary reconnection, first appears at all latitudes at the lower base of the CZ, and then, as a consequence, instead of the first wave without the participation of secondary reconnection, a second wave appears, which is the same primary wave, but with the participation of secondary reconnection, only at the latitudes of the equatorial surface of the Sun.
+In this case, the key role in the proposed mechanism of double maxima is played by two “waves” of toroidal fields, in which the first wave, with a strong field of $\sim 10^{7}~G$ and without the participation of secondary reconnection, first arises at all latitudes, and then, as a consequence, instead of the first wave, a second wave appears with a weak magnetic field of $2\cdot 10^{5}~G$ between $0.80R_{Sun}$ and $0.85R_{Sun}$, which is obtained from the same primary wave, but with the participation of secondary reconnection, only at the latitudes of the equatorial surface of the Sun.
 That is to say, the first maximum is the maximum of the number of spots that are sources of the magnetic flux tube (anchored in the overshoot tachocline) with $B\sim 10^{7}~G$ and with the participation of primary reconnection (see Fig.
 3 and Fig.
-5a; also Fig.
-9a), and the second maximum is the maximum of large spots that are the result of the appearance of magnetic flux tubes with $B\sim 10^{7}~G$ and with the participation of secondary reconnection (see Fig.
-5b and Fig.
-10b2,c).
+4a; also Fig.
+7a), and the second maximum is the maximum of large spots that are the result of the appearance of magnetic flux tubes with $B\sim 10^{7}~G$ and with the participation of secondary reconnection (see Fig.
+4b and Fig.
+8b2,c).
 From here, we are interested, in fact, in the very important secondary magnetic reconnection (see Fig.
-9b and Fig.
-10b2,c) during the rise of the magnetic flux tube in the solar photosphere, which is one of the two sources of the observed double maxima of one sunspot cycle – it shifts by $\sim 2.5~years$ between the primary (see the “red” curves in Fig.
-12b) and secondary (see the “green” curves in Fig.
-12b) sunspot maxima (see “smoothed” blue curves in Fig.
-12b).
+7b and Fig.
+8b2,c) during the rise of the magnetic flux tube in the solar photosphere, which is one of the two sources of the observed double maxima of one sunspot cycle – it shifts by $\sim 2.5~years$ between the primary (see the “red” curves in Fig.
+9b) and secondary (see the “green” curves in Fig.
+9b) sunspot maxima (see “smoothed” blue curves in Fig.
+9b).
 Moreover, it is the secondary maximum of sunspots that is also associated with both observational data of the tilt angle of Joy’s law (see Fig.
-10c from the right) and with the disappearance of sunspots on the surface of the Sun (see Fig.
-5c,d).
+8c from the right) and with the disappearance of sunspots on the surface of the Sun (see Fig.
+4c,d).
 And finally, we realize that the observed double maxima, for example in cycles 19-24 (see Fig.
-12a), also have different ratios between the numbers of very “fast” primary maxima $(S_{n})_{max}^{primary}$ (see the red curves in Fig.
-12b; also Fig.
-9a and Fig.
-10a) and very “slow” secondary maxima (see the “green” curves in Fig.
-12b)
-$(S_{n})_{max}^{primary}/(S_{n})_{max}^{secondary}\approx\begin{cases}2&at\sim 300~sunspots\\ 1.5&at\sim 150\div 200~sunspots\\ 1&at\sim 100~sunspots\end{cases}$ (59)
-and at the same time have almost the same difference between the secondary reconnection time $(\tau_{rec})_{2}\leqslant 2.5~years$ of the horizontal field in the tachocline (which is equal to the sum of $(\tau_{n})_{max}\approx 2$ and the lifetime of sunspots ($\leqslant 0.5~years$) from several days to several months (see S.
-K.
-Solanki (2003); K.
-Petrovay & L.
-van Driel-Gesztelyi (1997)), which shift by $\sim 2.0~years$,
+9a), also have different ratios between the numbers of very “fast” primary maxima $(S_{n})_{max}^{primary}$ (see the red curves in Fig.
+9b; also Fig.
+7a and Fig.
+8a) and very “slow” secondary maxima (see the “green” curves in Fig.
+9b)
+$(S_{n})_{max}^{primary}/(S_{n})_{max}^{secondary}\approx\begin{cases}2&at\sim 300~sunspots\\ 1.5&at\sim 150\div 200\\ 1&at\sim 100~sunspots\end{cases}$ (56)
+and at the same time have almost the same difference between the secondary reconnection time $(\tau_{rec})_{2}\leqslant 2.5~years$ of the horizontal field in the tachocline (which is equal to the sum of $(\tau_{n})_{max}\approx 2$ and the lifetime of sunspots ($\leqslant 0.5~years$) from several days to several months (see [101, 83]), which shift by $\sim 2.0~years$,
 $(\tau_{n})_{max}=(\tau_{rec})_{max}^{secondary}-(\tau_{rec})_{max}^{primary}\sim 2.0~years$ (57)
 between the “experimental” data of double sunspot maxima – very “slow” second maxima $(\tau_{rec})_{max}^{secondary}$ and very “fast” primary maxima $(\tau_{rec})_{max}^{primary}$ (see Fig. 9b).
-This raises an unexpected but valid question: How does the unique source of the butterfly diagram is only in the form of 11-year fundamental sunspots or is only in the form of alternating 22-year fundamental solar activity cycles?
-Here we show that, as expected, the source of the butterfly diagram is found only in the form of 11-year fundamental sunspots (see M.
-N.
-Gnevyshev (1967)).
-Below we show our solution with the so-called latitude-time interval (or gap) between the nearest butterflies of the remarkable solar physicist M.
-N.
-Gnevyshev (1967).
+This raises a valid question: How is the unique source of the butterfly diagram found only in the form of 11-year variations in sunspot numbers or only in the form of alternating 22-year fundamental cycles of solar activity?
+Here we show that, as expected, the source of the butterfly diagram is found only in the 11-year variations of the sunspot numbers (see [117]).
+Below we present an understanding of the so-called latitude-time interval (or gap) between nearby butterflies by the remarkable solar physicist Gnevyshev [117].
 To understand the physics of the Gnevyshev gap, we need to briefly describe the solution of the following interrelated problems:
-  • First, these are 11-year variations of sunspots (see V.
-D.
-Rusov et al.
-(2021)).
+  • First, these are 11-year variations of sunspots (see [14]).
 A unique result of our model is the fact that the 11-year periods, velocities and modulations of the S102 star (see Fig.
-6b in V.
-D.
-Rusov et al.
-(2021)) near the black hole are a significant indicator of the density modulation of the ADM halo in the fundamental plane of the Galactic center, which closely correlates with the density modulation of baryonic matter near the SMBH.
+6b in [14]) near the black hole are a significant indicator of the density modulation of the ADM halo in the fundamental plane of the Galactic center, which closely correlates with the density modulation of baryonic matter near the SMBH.
 If the ADM halo modulations in the GC black hole lead to ADM density modulations on the solar surface (via vertical density waves between the disk and the black hole to the solar neighborhood), then there is an “experimental” anticorrelation identity between such indicators as the 11-year ADM density modulation in the solar interior and the modulations of solar axions (or photons of axion origin), and equivalently between the modulations of solar axions (or photons of axion origin) and the sunspot cycles!
-  • Secondly, the most surprising fact is that the magnetic flux tube model at strong fields of the order of $10^{7}~G$ is a consequence of the thermomagnetic EN effect in the tachocline (see Fig.
-C.1a in V.
-D.
-Rusov et al.
-(2021)), which is in our case a remarkable source of the process with a continuous transformation of the toroidal field into a poloidal one (transformation $T\rightarrow P$), but not vice versa ($P\rightarrow T$) (see Fig.
-1b and Eq.
-(2); also Fig.
-C.1a in V.
-D.
-Rusov et al.
-(2021)).
-  • Thirdly, we know why we never see sunspots at the solar poles.
+  • Secondly, we know why we never see sunspots at the solar poles.
 The answer is quite simple.
 We know that both the tilt angle between the geometric and magnetic meridians, which are tilted by 7.25 degrees relative to the ecliptic, and the magnetic field of the northern (or southern) part of the meridional circulation, which radially corresponds to a length of about $\sim 22^{\circ}$, add up to about $\sim 30^{\circ}$.
-It follows clearly from this that the magnetic field of the meridional circulation part (see e.g.
-Fig.1b), will always slow down the magnetic field of the flux tube, which will almost never rise to the surface of the photosphere.
+It follows clearly from this that the magnetic field of the meridional circulation part will always slow down the magnetic field of the flux tube, which will almost never rise to the surface of the photosphere.
 This means that the magnetic fields of the flux tubes will only appear below about $90^{\circ}-30^{\circ}\approx 60^{\circ}$ at high latitudes.
 And that is precisely why no one observes sunspots at the poles of the Sun above $60^{\circ}$!
-  • And fourthly, we can finally observe (see Fig.
-6a of V.
-D.
-Rusov et al.
-(2021) for Japanese x-ray telescope Yohkoh (1991–2001)) and obtain the physics of axion origin photons only in those sunspots which are practically independent of the differential rotation of the Sun in the volume butterfly diagram.
+  • And Thirdly, we can finally observe (see Fig.
+6a of [14] for Japanese x-ray telescope Yohkoh (1991–2001)) and obtain the physics of photons of axion origin, which are associated with primary reconnection, only in those sunspots which are practically independent of the differential rotation of the Sun (when the equator rotates faster than the poles) in the butterfly diagram.
 Clearly, in other sunspots which depend exactly on secondary reconnection of magnetic flux tubes and, as a consequence, on the tilt angle of Joy’s law on the surface of the Sun (see Fig.
-10c), there can never be occurrences of axion origin photons in sunspots due to the differential rotation of the Sun!!!
+8c), there can never be occurrences of axion origin photons in sunspots due to the differential rotation of the Sun.
+Figure 10: The time variations of total solar irradiance (TSI; orange curve; see Fig.
+9 in [125]), simulated solar magnetic flux (black curve; see Fig.
+7 in [126]; and also Fig.
+16 in [127]) (both curves were smoothed with 29-year weighted moving average); the geomagnetic field secular variations (the Y-component; red curve) measured at the Eskdalemuir observatory [128]; variations in the Earth’s rotation speed (green curve; joined data from Fig.
+3 in [129], Fig.
+3.9 in [130], and the EOP C01 IAU1980 (1846-now) dataset [131]); the variations of the average global ocean level (Pacific Decadal Oscillation (PDO) [132] + Atlantic Multidecadal Oscillation (AMO) [133]; blue curve) with a trend in mm/years near 60-year oscillation (see Fig.
+1 in [134]).
+The latter three curves were first smoothed using 5-year moving average, and then the 11-year moving average.
+All considered time variations between the Sun and the Earth always have a time interval of approximately $\sim 10\div 11$ years.
+The only serious disagreements between the reconstructions and observations occur during the Second World War, especially in the period 1944-1945 (see [135]).
 From here we now return to the physics of the Gnevyshev gap.
@@
 Conversely, when the ADM density in the solar interior is high (or the sunspot count is low), then the Gnevyshev intervals will be longer in time.
-As no surprisingly, it is the presence of ADM density modulations that shows that the appearance of “our indicated” Gnevychev intervals is the result of precisely 11-year variations of butterfly cycles with sunspots!
+As no surprisingly, it is the presence of ADM density modulations that shows that the appearance of “our indicated” Gnevychev gap is the result of precisely 11-year variations of butterfly cycles with sunspots.
 The most interesting thing is that it is the 11-year variations of sunspot cycles or photons of axion origin (see Figs.
-5a and 6b in V.
-D.
-Rusov et al.
-(2021)) that are anticorrelated not only with the 11-year modulations of ADM density on the Sun, but also with the 11-year modulations of ADM density in the solar neighborhood, and, in particular, on Earth (see e.g.
+5a and 6b in [14]) that are anticorrelated not only with the 11-year modulations of ADM density on the Sun, but also with the 11-year modulations of ADM density in the solar neighborhood, and, in particular, on Earth (see e.g.
 Fig.
-13).
-Figure 13: The time variations of total solar irradiance (TSI; orange curve; see Fig.
-9 in (J.
-Pelt & O.
-Kärner, 2012)), simulated solar magnetic flux (black curve; see Fig.
-7 in M.
-Dikpati et al.
-(2008); and also Fig.
-16 in V.
-D.
-Rusov et al.
-(2013)) (both curves were smoothed with 29-year weighted moving average); the geomagnetic field secular variations (the Y-component; red curve) measured at the Eskdalemuir observatory (World Data Centre of Geomagnetism (2025)World Data Centre of Geomagnetism (Edinburgh), Edinburgh); variations in the Earth’s rotation speed (green curve; joined data from Fig.
-3 in N.
-S.
-Sidorenkov (2016), Fig.
-3.9 in N.
-S.
-Sidorenkov (2009), and the EOP C01 IAU1980 (1846-now) dataset by International Earth Rotation and Reference Systems Service (2025)); the variations of the average global ocean level (Pacific Decadal Oscillation (PDO), NOAA Physical Sciences Laboratory (2025) + Atlantic Multidecadal Oscillation (AMO), NOAA Physical Sciences Laboratory (2023); blue curve) with a trend in mm/years near 60-year oscillation (see Fig.
-1 in D.
-P.
-Chambers et al.
-(2012)).
-The latter three curves were first smoothed using 5-year moving average, and then the 11-year moving average.
-All considered time variations between the Sun and the Earth always have a time interval of approximately $\sim 10\div 11$ years.
-The only serious disagreements between the reconstructions and observations occur during the Second World War, especially in the period 1944-1945 (see C.
-K.
-Folland et al.
-(2018)).
+10).
 First, we recall that the physics of experimental data related to the anticorrelation between 11-year variations in the luminosity of the Sun and the Earth (see Fig.
-13) currently cannot explain one of the most mystical phenomena of solar-terrestrial physics, which is contained in well-known experiments, for example, between the temporal variations of the magnetic flux of the solar tachocline (black curve; see Fig.
-7 in M.
-Dikpati et al.
-(2008); and also Fig.
-16 in V.
-D.
-Rusov et al.
-(2013)), and secular variations in the Earth’s magnetic field (Y-component from World Data Centre of Geomagnetism (2025)World Data Centre of Geomagnetism (Edinburgh) (Edinburgh)), or variations in the Earth’s rotation speed (see Fig.
-3 in N.
-S.
-Sidorenkov (2016), Fig.
-3.9 in N.
-S.
-Sidorenkov (2009), and EOP C01 IAU1980 (1846-now) dataset by International Earth Rotation and Reference Systems Service (2025)), as well as the variations in the mean global sea level (Pacific Decadal Oscillation (PDO) (NOAA Physical Sciences Laboratory, 2025) + Atlantic Multidecadal Oscillation (AMO) (NOAA Physical Sciences Laboratory, 2023)) with a trend in mm/years near 60-year oscillation (see Fig.
-1 in D.
-P.
-Chambers et al.
-(2012)).
-From here follows the main argument for solving the above problem (see Fig.
-13; and also Fig.
-6 in V.
-D.
-Rusov et al.
-(2021)).
-We understand that it is the presence of asymmetric dark matter (ADM) in the Earth’s core that can, on the one hand, be the reason for solving the density deficit in the inner core, and on the other hand, achieve the necessary consensus (see D.
-Antonangeli et al.
-(2018); E.
-Edmund et al.
-(2019)) not only regarding the absolute values of the solid iron velocities under the conditions of the inner core, but also regarding the iron melt velocities from pressure and temperature.
-If this is true, then we can prove that the inner core is not only a hidden reservoir of hydrogen on Earth (see e.g.
-T.
-Sakamaki et al.
-(2016), but also the cause of the existence of a heat flow source from the Earth’s core, which, based on ADM variability (see V.
-D.
-Rusov et al.
-(2021)), generates 11-year thermal, magnetic and temperature cycles in all parts of the Earth, including from the Earth’s surface to the troposphere.
-This means that when the maximum of magnetic fluxes on the Earth’s surface exceeds the energy of solar irradiation in the troposphere, and the minimum of these fluxes on the Earth’s surface becomes equal to solar irradiation in the troposphere, then the 11-year luminosity of the Earth will anticorrelate with the 11-year luminosity of the Sun.
+10) currently cannot explain one of the most mystical phenomena of solar-terrestrial physics, which is contained in well-known experiments, for example, between the temporal variations of the magnetic flux of the solar tachocline (black curve; see [126] and Fig.
+16 in [127]), and secular variations in the Earth’s magnetic field (Y-component from [128]), or variations in the Earth’s rotation speed (see [129, 130] ), as well as the variations in the mean global sea level (Pacific Decadal Oscillation (PDO) [132] + Atlantic Multidecadal Oscillation (AMO) [133]) with a trend in mm/years near 60-year oscillation (see Fig.
+1 in [134]), between which there is a time lag of about 10-11 years, respectively (see Fig.
+10; and also Fig.
+6 in [14].
 III Summary and Outlook
-The main result is the following: three “free” parameters – the magnetic field in the tachocline of the order of $\sim 10^{7}~G$ (see Fig.
-(A.1) and Eq.
-(A17) in V.
-D.
-Rusov et al.
-(2021)), axions with mass $m\sim 3.2\cdot 10^{-2}~eV$ and coupling constant $g_{a\gamma}\sim 4.4\cdot 10^{-11}~GeV^{-1}$ (see Eq.
-(11) in V.
-D.
-Rusov et al.
-(2021)) and the asymmetry dark matter (ADM) in the Universe with $m\sim 5~GeV$ (see V.
-D.
-Rusov et al.
-(2021); A.
-C.
-Vincent et al.
-(2016)) – are the source of a complete solution of the problem of an alternative theory of magnetic flux tubes in strong fields using 11-year variations of axion origin photons, which are caused by anticorrelations of 11-year variations of ADM density, gravitationally captured on the Sun.
-From here, we understand that the fraction of dark matter in the Universe, which is about $\sim$27%, should be the reason for the appearance of a very large amount of dark matter with $m\sim 5~GeV$, for example, in our Milky Way galaxy.
-This means that the dark matter ADM in our galaxy should be not only in stars, but also on planets, for example, the Sun and the Earth.
-As not surprisingly, a simple but fundamental conclusion emerges.
-The anticorrelation of the 11-year variations in the ADM density, gravitationally trapped both on the Sun and on the Earth, must necessarily participate in all the physical problems of dark matter, which are related, for example, not only to our solution of magnetic flux tubes in strong fields, describing the physics of sunspots, but also to the solution of the global climate of the Earth, describing the physics of global warming due to temperature and magnetic fields, but, crucially, without the existence of anthropogenic enhancement of the greenhouse effect.
-The latter suggests that, in an intriguing perspective, the experimental observation of anticorrelations between 11-year variations in the luminosity of the Sun and the Earth (see Fig.
-13) leads to the emergence of dark matter asymmetry in the Earth’s core, which, as not surprisingly, controls, among other things, the magnetic field from the core to the Earth’s surface.
-This means that when the maximum magnetic field at the Earth’s surface exceeds the solar irradiation energy in the troposphere, and the minimum of such energy flux at the Earth’s surface is equal to the solar irradiation energy in the troposphere, then the 11-year luminosity of the Earth will anticorrelate with the 11-year luminosity of the Sun.
-This opens up a powerful prospect for observing both the axions, which in this mass range may be the main component of dark matter, and the anticorrelation of 11-year variations in the density of ADMs gravitationally trapped on both the Sun and Earth.
+Most importantly, the existing void in the tube causes high velocities and short lifetimes of sunspots from the tachocline to the photosphere, which is consistent with experimental data.
+This means that the modernity of a well-developed helioseismological inversion, surprisingly, persists to this day.
+The perspective is to understand the physics of the anticorrelation between 11-year variations of total solar irradiance and 11-year variations of the magnetic field and other Earth parameters (see Fig.
+10).